POLARIS  Network Activity Log; Last Updated: Tue Jul 10 20:49:45 EDT 2007
No Date Time UT Location Activity Comments
1 2021-10-27 12:56 A3PN TDMA Halt Removed from TDMA (CYG253, Carina 115) until next visit. Station down due to antenna shift.
2 2021-09-09 ACKN New Station Start DC (Solar Panels).
3 2021-07-20 20:20 ACKN Temp Stop Stopped recording in ACKN ringbuffers to swap electronics.
4 2021-07-20 20:21 ACKN Hardware Swap Cygnus 160/HRD 211 replaced with Cygnus 220/Trident 141.
5 2021-07-20 21:00 ACKN SR Change Sampling rate from to 20 sps to 40 sps
6 2021-07-20 21:00? ACKN Restart Started recording in new ACKN ringbuffers.
7 2021-08-13 18:28 ACKN Firmware Upgrade Upgraded to Libra V5.20, TS V1.20, TR V1.50 EPLD V1.08
8 2021-09-14 15:30-16:30 ACKN SSPB Check Corrected SSPB temperature profile from flat 45 dB to manufacturer specs.
9 2021-09-21 12:23 ACKN Firmware Upgrade Upgraded to Libra Firmware version 5.31.
10 2021-07-14 20:10 ACKN Clean up Garbage removed from site.
11 2021-07-14 20:10 ACKN Feed Horn Box New Feed Horn protective box installed. New conduits installed.
12 2021-07-14 20:10 ACKN Feed Window Replaced.
13 2021-07-14 20:10 ACKN Firmware Upgrade Libra 5.6201, TS 1.23, TRD 1.701 running.
14 2021-07-14 20:10 ACKN Gain Check Software gain 1.048576, DC removal filter off, Input range 40 Vpp.
15 2021-07-14 20:10 ACKN Hub Swap Served by Ottawa Hub.
16 2021-07-14 20:10 ACKN Re-Point 35 dB co-pol 1 x-pol.
No Date Time UT Location Activity Comments
17 2021-07-14 20:10 ACKN SSPB Check Two watts sspb installed. S/N 16461.
18 2021-07-14 20:10 ACKN Service Visit Station was down. GPS cable was chewed up, feed window memberane damaged, water in feed horn assembly.
19 2021-08-12 18:38 ACKN DC Filter Check Unchanged from Off.
20 2021-08-12 18:38 ACKN Gain Change Software gain changed from 1.048576 to 1.0.
21 2021-09-30 18:00 ACKN DC Filter Check DC filter unchanged at off.
22 2021-09-30 18:00 ACKN Gain Check Software gain unchanged at 1.0 at 40 sps.
23 2021-09-30 18:00 ACKN Hardware Swap Cygnus 220 removed, 265 installed. 2 W SSPB removed, 1 W installed (SN A0208A).
24 2021-09-30 18:00 ACKN Re-Point Much improved. X-pol 0 Co-pol 28 dB. Additional bird protection (shark teeth) installed.
25 2021-09-30 18:00 ACKN SSPB Check All looked ok at station and at Hub (on analyzer), but no data received, even after a Cygnus swap. It was concluded that the new 2 W sspb was not working.
26 2021-09-30 18:00 ACKN Service Visit Visit initiated due to poor VSAT performance.
27 2021-09-30 18:00 ACKN VSAT Check Before re-pointing: X-pol 0 Co-pol 20 dB.
28 2021-11-03 19:31 ACKN Firmware Upgrade Trident 1.7210 in A and B. A is active.
29 2021-08-22 14:56 ACKN Restart Trident 141 shows bus gaps, rebooted to reset negative counter on "Bytes sent".
30 2021-09-12 17:19 ACKN Firmware Upgrade Libra 5.7005A installed.
31 2021-09-12 17:24 ACKN Firmware Upgrade TS 2.0000B installed. Trident 1.7210 already running.
32 2021-09-12 17:26 ACKN Restart Trident rebooted. Transmit errors before reboot 9640, after 0.
No Date Time UT Location Activity Comments
33 2021-09-12 17:29 ACKN Comments Monitoring station for bus gaps. Would firmware upgrade help?
34 2021-09-17 13:14 ACKN Comments Bus gaps continue. Trident 141 Instrument screen shows Transmit overruns of 3361 and errors of 3899 and receive overruns of 5708 after about 5 days since last reboot of TimeServer.
35 2021-07-11 14:40 ACKN Closed End of Operation.
36 2022-01-31 ACTO New Station Start SOSN/POLARIS. AC Powered.
37 2022-02-08 ACTO Seismo Swap Changed from CMG-3ESP to CMG-40T.
38 2022-02-22 ACTO Seismo Swap Changed from CMG-40T to CMG-3ESP.
39 2021-08-13 22:50 ACTO Firmware Upgrade Upgraded to Libra V5.20, TS V1.20, TR V1.50 EPLD V1.08
40 2021-09-23 22:15 ACTO Firmware Upgrade Upgraded to Libra Firmware version 5.31.
41 2022-04-11 16:56 ACTO Hardware Swap Now running CYG174 TRD77.
42 2022-04-11 16:56 ACTO Site Visit Maintenance visit.
43 2021-12-19 21:00 ACTO Service Visit Station down about 1/2 hour to swap Trident.
44 2021-12-19 21:30 ACTO Hardware Swap Trident 77 removed, 295 installed.
45 2021-12-19 21:31 ACTO DC Filter Check Unchanged from on at 1 mHz.
46 2021-12-19 21:32 ACTO Gain Check Unchanged from 0.5.
47 2021-06-22 18:24 ACTO Require GPS Lock GPS receiver installed on Libra com port, recording at London Hub.
48 2021-09-15 14:00 AKVQ New Station Code Internet, Ottawa Hub, at Akulivik, Hudson Bay.
No Date Time UT Location Activity Comments
49 2021-09-15 14:05 AKVQ Internet Station On VPN.
50 2021-09-15 14:10 AKVQ Hardware List Taurus 192, ESP T3D60, IP 86.18.
51 2021-09-15 14:20 AKVQ 40 sps, Gain 1, No DC filter, Require GPS Lock Checked.
52 2021-09-15 14:30 AKVQ Comments Internet TX problems, gaps in data.
53 2021-10-02 20:18 ALFO DC Filter Check Started with DC filter on at 0.001 Hz.
54 2021-10-02 20:18 ALFO Gain Check Started with software gain of 0.5 at 100 sps, input range 16 Vp-p.
55 2021-05-27 14:47 ALFO Service Visit To replace seismometer. Site looked ok.
56 2021-05-27 14:50 ALFO Seismo Swap Trillium 0075 removed. ESP T3J04 installed.
57 2021-05-27 14:51 ALFO Gain Check Unchaged at 0.5.
58 2021-05-27 14:52 ALFO DC Filter Check Unchaged at on and 1 mHz.
59 2021-05-27 14:53 ALFO Comments Trident Input range changed from 16 (for Trillium) to 40 Vpp(for ESP).
60 2021-05-27 16:10 ALFO Cold-Boot Standard procedure.
61 2021-05-27 16:15 ALFO Comments For next visit: Security cable for battery box and Proper upper pedestal to fit over leveling plate.
62 2021-12-14 16:00 ALFO Service Visit To re-point and charge batteries.
63 2021-12-14 18:20 ALFO Re-Point X/Co. 28/3 before, 32/2 dB after repoint.
64 2021-12-14 18:54 ALFO Comments All is ok. Trident screen was checked. Overruns looks ok.
No Date Time UT Location Activity Comments
65 2021-12-14 18:55 ALFO Battery Check Charging in progress.
66 2021-12-14 20:06 ALFO Service Visit End of Serivice. Batteries were charged for about 4 houurs. All ok.
67 2021-10-26 16:42 ALFO Service Visit Station has been dead, found Cygnus to be dead.
68 2021-10-26 16:43 ALFO Hardware Swap Dead Cygnus 175 removed, Cygnus 269 installed.
69 2021-10-26 16:48 ALFO Comments Dataserver restarted at London Hub (ALFO is served by this Hub).
70 2021-09-11 23:05 ALGO New Station Code Start recording in ALGO ringbuffers.
71 2021-09-17 00:25 ALGO DC Filter Change From off to on. At 0.001 mHz corner frequency.
72 2021-09-17 00:25 ALGO Gain Change From 1.048576 to 0.1.
73 2021-09-20 19:15 ALGO Firmware Upgrade Upgraded to Libra Firmware version 5.31.
74 2022-02-19 23:30 ALGO Hardware Swap Cygnus 266 replaced with Cygnus 185.
75 2022-05-07 17:39 ALGO Site Visit Cygnus installed in battery box. 2 more batteries. GPS cable in conduit. Blanket removed from Trident.
76 2022-05-07 17:39 ALGO Site Visit Maintenance visit.
77 2022-05-07 17:39 ALGO VSAT Check Co-pol 33, X-pol 2dB. Adjustment not needed.
78 2022-03-15 16:48 ALGO DC Filter Change Changed from ON @0.0010 Hz to OFF. Requested by Savka D. for calibration test. Will be restored to on later.
79 2022-03-15 20:30 ALGO DC Filter Change Changed from OFF to ON at 0.001 Hz. Savka's test finished.
80 2021-08-30 01:32 ALRB New Station Code Anahim Lake, Lehmann Ranch, BC, recording at Ottawa Hub, Naqs1,2, Telesat.
No Date Time UT Location Activity Comments
81 2021-08-30 01:36 ALRB Hardware List Cyg192, Trd91, TS37, SSPB A0177A, ESP T3B25.
82 2021-08-30 01:37 ALRB DC Filter Check Off.
83 2021-08-30 01:37 ALRB Gain Check 1.0 at 100 sps.
84 2021-08-30 01:38 ALRB Require GPS Lock On.
85 2021-08-30 01:38 ALRB VSAT Check 29.2 dB isolation.
86 2021-08-30 01:39 ALRB Comments Site installed without analyzer, helped by Telesat.
87 2021-09-06 16:58 ALRB Service Visit To fix bus gap problems.
88 2021-09-06 16:59 ALRB Comments Bus and grounding cables checked, all ok.
89 2021-09-06 16:59 ALRB Gain Check 1.0 at 100 sps, as before.
90 2021-09-06 16:59 ALRB Hardware Swap Trident 91 removed, Trd695 installed.
91 2021-09-06 17:00 ALRB Comments This is the end of Nechako basin installation/service.
92 2021-09-06 17:00 ALRB DC Filter Check Off, as before.
93 2021-09-06 17:00 ALRB Require GPS Lock On, as before.
94 2021-11-18 19:35 ANMB New Station Code Start recording in ANMB ringbuffers.
95 2021-06-18 ANMB Service Visit All ok. Fixed fence.
96 2021-08-19 17:31 ANMB Closed Station closed at 2004-08-19-17-31. Site owner requested removal of station since the property is sold.
No Date Time UT Location Activity Comments
97 2021-08-17 22:06 AP3N New Station Start Started new Hudson Bay station at Apex 3 site, recording on Naqs2, Ottawa Hub.
98 2021-08-17 22:07 AP3N Gain Check Gain of 1.0, sps 40.
99 2021-08-17 22:07 AP3N Hardware List Cyg 253, Trd 158, sspb a001a, 2w, seismo ESP T3B50.
100 2021-08-17 22:08 AP3N VSAT Check Telesat, first site on new bandwidth, 35 dB separation.
101 2021-08-17 22:09 AP3N DC Filter Check DC filter is not on.
102 2021-08-17 22:10 AP3N Require GPS Lock Require GPS Lock is on.
103 2021-08-17 22:11 AP3N Site Geology Solid bedrock at 69 27 31.1 N 84 27 41.7 W 437 m.
104 2021-08-17 22:20 AP3N Comments Started on Naqs1 at Ottawa Hub. Dataserver restarted.
105 2021-08-12 22:37 ARVN New Station Start Started new station at Arviat, HB, Ottawa Hub Naqs2, Telesat Carina 116.
106 2021-08-12 22:38 ARVN VSAT Check 31 db isolation.
107 2021-08-12 22:39 ARVN Hardware List Same as DVKN, Cyg243, Trd239, TS 056, 40T/T4187.
108 2021-08-12 22:40 ARVN Comments Libra VSAT IP is 10.15.255.243
109 2021-08-12 22:41 ARVN DC Filter Check Off.
110 2021-08-12 22:41 ARVN Gain Check Gain is 1.0 at 40 sps.
111 2021-08-12 22:42 ARVN Carina Hubs Both Cyg and Carina needed Multicast Forwarding at Naqs to work.
112 2021-08-12 22:42 ARVN Require GPS Lock On.
No Date Time UT Location Activity Comments
113 2021-08-12 22:44 ARVN Comments This is an AC station.
114 2021-08-12 22:56 ARVN New Station Start Started new station at Arviat, HB, Ottawa Hub Naqs1.
115 2021-08-12 22:57 ARVN Comments Naqs1 dataServer restarted.
116 2021-08-30 00:00 ARVN Comments Station moved by about one km to new locatoin for beeter AC connection. Old data to be junked, station code to be kept the same.
117 2021-10-27 12:51 ARVN TDMA Halt Removed from TDMA (CYG243, Carina 115) until next visit. Station down due to antenna shift.
118 2021-06-09 21:34 ATKO VSAT Check Co/X 29/0 dB. Ajusted IFL to 2 dB for long Cable. Correct SSPB profile entred.
119 2021-06-09 21:36 ATKO Hardware List CYG539, TS696, TRD059, Seismo. ESP T377, SSPB A0395A
120 2021-06-09 21:37 ATKO Gain Check 100 SPS, Gain 0.5, startup gain.
121 2021-06-09 21:38 ATKO DC Filter Check DC filter on at 0.001Hz. Startup.
122 2021-06-09 21:39 ATKO Site Geology Solid bedrock on hill top about 30 ft high about 40 yrds from access road.
123 2021-06-09 21:40 ATKO Comments Old (may be inactive) hydro wire nearby.
124 2021-06-09 21:53 ATKO New Station Code Started recording on Naqs2 in ATKO ringbuffers.
125 2021-06-09 21:58 ATKO New Station Code Started recording in Naqs1 in ATKO ringbuffers.
126 2022-02-27 15:05 B7-VAULT New Station Code O88T, Internet station. Started.
127 2022-02-27 15:07 B7-VAULT Hardware List CYG284 TIM228 TRD231, 3ESP
128 2022-02-27 15:09 B7-VAULT Gain Check Started at 100 SPS, 40 VPP, Gain of 0.5
No Date Time UT Location Activity Comments
129 2022-02-27 15:10 B7-VAULT DC Filter Check Started with NO DC filter.
130 2021-08-25 17:30 BANO New Station Code Start recording in BANO ringbuffers.
131 2021-09-23 23:40 BANO Firmware Upgrade Upgraded to Libra Firmware version 5.31.
132 2021-10-01 18:10 BANO Hardware Swap Cygnus 281 from 269, Trident 292 from 256
133 2021-10-22 17:43 BATG New Station Start Bathurst, NB, recording at London Hub.
134 2021-10-22 18:08 BATG Hardware List Cygnus 605, Trd 269, sspb 1w A0126A, seismo esp s/n T3D78.
135 2021-10-22 18:44 BATG DC Filter Check Off.
136 2021-10-22 18:44 BATG Gain Check 1.0 at 100 sps.
137 2021-10-22 18:44 BATG VSAT Check 26/0 dB.
138 2021-10-22 18:45 BATG Power Check AC site.
139 2021-10-22 18:45 BATG Require GPS Lock On.
140 2021-10-22 18:45 BATG Site Geology On edge of small dam bolted to bedrock.
141 2021-08-31 21:07 BATG Carina Swap Switched from SES-AM to Telesat on Ottawa Hub.
142 2021-08-31 21:08 BATG Hardware Swap Cygnus 422 installed, 605 removed.
143 2021-08-31 21:09 BATG VSAT Check 30 dB.
144 2021-08-31 21:15 BATG Comments Polarization was off by 90 deg during Telesat move, resulted in confusion. Fixed.
No Date Time UT Location Activity Comments
145 2021-08-02 23:50 BOXN New Station Code Start recording in BOXN ringbuffers.
146 2021-08-23 21:58 BOXN Firmware Upgrade Upgraded to Libra V5.20, TS V1.20, TR V1.50 EPLD V1.08
147 2021-09-14 22:45 BOXN SSPB Check Corrected SSPB temperature profile from flat 45 dB to manufacturer specs.
148 2021-09-21 12:37 BOXN Firmware Upgrade Upgraded to Libra Firmware version 5.31.
149 2021-07-20 00:19 BOXN Before Re-Point Before any re-adjustment, co-pol was 34 and x-pol was 10 dB, indicating antenna move.
150 2021-07-20 00:19 BOXN Feed Horn Box New Feed Horn protective box installed. New conduits installed.
151 2021-07-20 00:19 BOXN Firmware Upgrade Libra 5.6201 installed?
152 2021-07-20 00:19 BOXN Re-Point After re-pointing, co-pol was 38 with x-pol of 5 dB. Much improved.
153 2021-07-20 00:19 BOXN Service Visit Site was working OK. Stopped for service. Solar panel rack was loose, fixed.
154 2021-07-20 01:29 BOXN Restart Calibrated and mass-centred ok.
155 2021-09-17 01:35 BPRO DC Filter Change From off to on. At 0.001 mHz corner frequency.
156 2021-09-17 01:35 BPRO Gain Change From 1.048576 to 0.1.
157 2022-02-06 BRCO New Station Start SOSN/POLARIS. AC Powered.
158 2022-02-08 BRCO Seismo Swap Changed from CMG-3ESP to CMG-40T.
159 2022-03-01 BRCO Seismo Swap Changed from CMG-40T to CMG-3ESP.
160 2021-08-12 22:45 BRCO Firmware Upgrade Upgraded to Libra V5.20, TS V1.20, TR V1.50 EPLD V1.08
No Date Time UT Location Activity Comments
161 2021-09-16 21:10 BRCO DC Filter Change From off to on. At 0.001 mHz corner frequency.
162 2021-09-16 21:10 BRCO Gain Change From 1.048576 to 0.1.
163 2021-09-23 17:51 BRCO Firmware Upgrade Upgraded to Libra Firmware version 5.31.
164 2022-04-19 22:00 BRCO Firmware Upgrade TimeServer V2.00.00 upgrade to fix GPS SOH problem. Not sure if upgrade fixed the problem or simply the reboot did it. Served by BD, data entry IA.
165 2021-09-10 16:50 BRPO New Station Code Start recording in BRPO ringbuffers.
166 2021-09-20 22:10 BRPO Firmware Upgrade Upgraded to Libra Firmware version 5.31.
167 2021-10-15 09:?? BRPO Closed Moving station to safer location.
168 2021-11-02 18:00 BUKO New Station Start Started new statation (moved hardware from PTCO).
169 2021-11-02 18:01 BUKO VSAT Start 28/1 dB.
170 2021-11-02 18:02 BUKO Hardware List CYG179 TRD071 ESP-T3B18
171 2021-11-02 18:03 BUKO Site Geology Bedrock (outcrop).
172 2021-11-02 18:04 BUKO Power Upgrade Needed on next visit. Only 5 solar panels installed.
173 2021-11-03 13:50 BUKO Gain Check Started with software gain of 0.5.
174 2021-11-03 13:51 BUKO DC Filter Check Started with DC filter on at 1 mHz.
175 2021-11-17 19:08 BUKO Service Visit Added one solar panel, replaced some battery terminals
176 2021-11-17 19:12 BUKO Battery Check Starting to charge batteries with generator.
No Date Time UT Location Activity Comments
177 2021-11-17 19:13 BUKO Restart Site restarting.
178 2021-11-18 13:00 BUKO Battery Check Generator stopped. Lasted about 18 hours??
179 2021-11-19 22:24 BUKO Battery Check Generator started again. Battery volts 12.6 VDC
180 2021-08-17 01:58 BULN New Station Start Bullion, Hudson Bay, starting Ottawa Hub Naqs1, Naqs2.
181 2021-08-17 01:59 BULN Hardware List Cyg246, Trd235, TS105, ESP T3A86, SSPB H013A, redeployed from YMBN.
182 2021-08-17 02:01 BULN Battery Check Only 11 batteries (new) were available.
183 2021-08-17 02:01 BULN VSAT Check 31 dB Isolation, Telesat, Carina 116.
184 2021-08-17 02:02 BULN DC Filter Check Off.
185 2021-08-17 02:02 BULN Gain Check Gain 1, sp 40 sps.
186 2021-08-17 02:02 BULN Require GPS Lock On.
187 2021-08-17 02:03 BULN Site Geology 200 m off runway.
188 2021-09-07 18:27 BULN VSAT Check IFL loss faked to 4.0 db to increase power. EIRP inc is already at max. 10 dB. This improved performance.
189 2021-07-30 03:45 CAMN New Station Code Start recording in CAMN ringbuffers.
190 2021-08-12 21:45 CAMN Firmware Upgrade Upgraded to Libra V5.20, TS V1.20, TR V1.50 EPLD V1.08
191 2021-09-14 22:10 CAMN SSPB Check Corrected SSPB temperature profile from flat 45 dB to manufacturer specs.
192 2021-07-19 21:47 CAMN Before Re-Point Before any re-adjustment, co-pol was 35 and x-pol was 15 dB, indicating antenna move.
No Date Time UT Location Activity Comments
193 2021-07-19 21:47 CAMN Feed Horn Box New Feed Horn protective box installed. New conduits installed.
194 2021-07-19 21:47 CAMN Firmware Upgrade Libra 5.6201 installed?
195 2021-07-19 21:47 CAMN Re-Point After re-pointing, co-pol was 35 with x-pol of 2 dB. Improved.
196 2021-07-19 21:47 CAMN Service Visit Site was working OK. Stopped for service.
197 2021-07-19 22:54 CAMN Restart Calibrated & mass-centred ok.
198 2021-07-15 17:41 CAMN Closed End of operation.
199 2022-01-25 14:00 CARINA 92 Comments Ottawa Hub, Telesat antenna, stations on Hub 92 not working, see SOH graphs.
200 2022-01-25 15:00 CARINA 92 VSAT Check Ottawa Hub, Telesat antenna, No RX.
201 2022-01-25 15:10 CARINA 92 SSPB Check SSPB and LNB replaced.
202 2022-01-25 15:30 CARINA 92 Re-Point Telesat antenna needed repoint, x/c before 5/25, after 0/32 dB.
203 2022-01-25 15:40 CARINA 92 Comments Repoint needed early spring and mid-December.
204 2022-01-25 16:00 CARINA 92 VSAT Start Ottawa Hub, Telesat antenna up and ok.
205 2021-07-30 20:17 CBRQ New Station Code Started first Quebec station at Cabonga Reservoir.
206 2021-07-30 20:18 CBRQ Comments Served by Ottawa Hub. Naqs2 recording started 20:17, Naqs1 at 20:21.
207 2021-07-30 20:19 CBRQ Hardware List Cyg 561, TS 713, TRD 754, ESP Seism. T3K12, sspb a0496a.
208 2021-07-30 20:20 CBRQ VSAT Check Co/X 35/4 dB.
No Date Time UT Location Activity Comments
209 2021-07-30 20:21 CBRQ Site Geology Solid bedrock.
210 2021-07-30 20:22 CBRQ Cold-Boot Before taking screen captures.
211 2021-07-30 20:23 CBRQ Gain Check Started with gain of 0.5 at 100 sps.
212 2021-07-30 20:24 CBRQ DC Filter Check Started with dc filter on at 0.001 hz.
213 2021-07-30 20:25 CBRQ Comments Ringbuffers restarted.
214 2021-05-19 04:46 CBRQ Closed Station closed.
215 2021-10-19 15:32 CHEG New Station Start Ottawa Hub, Naqs2, Telesat, Cheticamp, NS
216 2021-10-19 15:34 CHEG Hardware List Cygnus 562, Trd 760, sspb H0014a 1w, esp seismo.
217 2021-10-19 15:38 CHEG DC Filter Check Off.
218 2021-10-19 15:38 CHEG Gain Check 1.0 at 100 sps.
219 2021-10-19 15:38 CHEG Require GPS Lock On.
220 2021-10-19 15:38 CHEG VSAT Check 37 dB isolation.
221 2021-10-19 15:39 CHEG Site Geology Seis on patch of bedrock.
222 2021-10-19 15:40 CHEG Comments VSAT on mount bolted to bedrock.
223 2021-10-19 15:40 CHEG Seismo Check Seismo at 46.80836 N 60.67447 W 1463 ft
224 2021-10-19 15:42 CHEG Comments Mass centred and calibrated ok.
No Date Time UT Location Activity Comments
225 2021-10-19 15:50 CHEG Comments Started recording at Ottawa Naqs1.
226 2021-10-19 15:51 CHEG Comments Dataserver restarted on London Hub, Naqs1.
227 2021-10-19 15:51 CHEG Comments Used Telesat here since GE had poor coverage.
228 2021-08-29 19:10 CHEG Service Visit All ok, solar panel rack leg fixed.
229 2021-08-29 19:11 CHEG SSPB Check Profile adjusted from 45 to 50 dB flat, EIRP increase adjusted from -5 to 0 to balance.
230 2021-08-29 19:11 CHEG VSAT Check 31.5 dB isolation.
231 2021-08-29 19:12 CHEG Comments Site in very good condition, bolts on solar panel secured.
232 2021-10-11 20:15 CLPO New Station Code Start recording in CLPO ringbuffers.
233 2021-10-11 20:16 CLPO Seismo Check Started with CMG-ESP (standard vault), Trillium (Test vault).
234 2021-10-11 20:17 CLPO Gain Check This station started with gain 0.1, iir on at 0.001 on both Tridents? Check with CA
235 2022-05-06 14:49 CLPO Seismo Check Trillium Seismometer and its Trident removed.
236 2022-05-06 14:49 CLPO Site Visit Maintenance visit.
237 2022-05-06 14:49 CLPO VSAT Check Co-pol 32, X-pol 9dB. Adjusted, co-pol 34, X-pol 2 dB.
238 2021-09-01 22:00 CLSB New Station Code Cack Lake, Smith's Ranch, SW of Quesnel, BC. Started on Ottawa Hub, Telesat.
239 2021-09-02 14:38 CLSB Hardware List Cyg273, Trd425, TS170, ESP T3B26, SSPB H0020A.
240 2021-09-02 14:39 CLSB VSAT Check 30 dB, strong signal on Telesat.
No Date Time UT Location Activity Comments
241 2021-09-02 14:40 CLSB DC Filter Check Off.
242 2021-09-02 14:40 CLSB Gain Check 1.0 at 100 sps.
243 2021-09-02 14:40 CLSB Require GPS Lock On.
244 2021-09-02 14:41 CLSB Comments GPS slow to connect. Bad seismo spotted, replaced.
245 2021-09-02 14:41 CLSB Site Geology Packed clay, trees nearby.
246 2021-11-17 18:50 CLVB New Station Code Start recording in CLVB ringbuffers.
247 2021-12-06 21:30 CLVB Battery swap. All six batteries swapped with new ones. All new firmware installed.
248 2021-11-21 00:47 CLVB Battery Check Charged for many hours using generator.
249 2021-11-21 00:48 CLVB Restart Station restarted after battery charged.
250 2021-11-21 00:49 CLVB Comments LVD set to 10.5 VDC.
251 2021-11-21 00:50 CLVB Battery Check Batteries charged for 5 hours.
252 2021-11-22 21:33 CLVB Battery Check Low battery volts, 11.92 VDC. 11.67 VDC.
253 2021-11-23 18:05 CLVB Temp Halt Stopped at 11.52 VDC due to low battery power.
254 2021-12-09 19:45 CLVB Service Visit 8 New fully charged Batteries installed, old batteries removed.
255 2021-12-09 20:03 CLVB Battery Check All batteries are ok now. VDC is 12.92.
256 2021-12-09 20:03 CLVB Restart Site restarting.
No Date Time UT Location Activity Comments
257 2021-12-09 20:05 CLVB Battery Check Solar panel is charging ok too. 20 VDC.
258 2022-01-16 18:00 CLVB Service Visit To swap batteries.
259 2022-01-16 18:10 CLVB Battery Check 8 charged batteries installed.
260 2022-01-16 18:20 CLVB VSAT Check Co/X-pol 33/4 dB.
261 2022-01-16 19:30 CLVB Cold-Boot All started after cold boot.
262 2022-01-16 19:36 CLVB Restart Station restarted.
263 2021-12-31 20:00 CLVB Service Visit To add a second battery box and 6 batteries.
264 2021-12-31 20:10 CLVB VSAT Check Was good at 35/5 dB.
265 2022-05-12 23:30 CLVB Service Visit Station was down.
266 2022-05-12 23:31 CLVB Wildlife Cables were chewed by bear. All cables including power, bus, RF, etc were repaired or replaced.
267 2022-05-12 23:45 CLVB VSAT Check X/Co pol 2/34 dB.
268 2022-05-12 23:47 CLVB Restart After cold boot.
269 2021-08-04 15:19 CLVB Closed End of operation.
270 2021-07-25 17:25 CLWO DC Filter Check On at 0.001 Hz.
271 2021-07-25 17:25 CLWO Gain Check Station Started at software gain of 0.5, sample rate 100 sps.
272 2021-07-25 17:25 CLWO Hardware List CYG183 TRD85 CMG-40T 4595, SSPB A0306A
No Date Time UT Location Activity Comments
273 2021-07-25 17:25 CLWO New Station Code Start recording in CLWO ringbuffers.
274 2021-07-25 17:25 CLWO Site Geology Limestone bedrock beneath shallow overburden.
275 2021-07-25 17:25 CLWO VSAT Start Co-pol 25 X-pol 0
276 2022-03-26 17:00 CLWO Service Visit To replace seismo. Station found in good and secure condition. Snow in the field. Walked to the site.
277 2022-03-26 17:10 CLWO Hardware Swap 40T removed, ESP (T3G91) installed.
278 2022-03-26 17:20 CLWO VSAT Check Antenna re-pointed. X/Co: 1/29 dB.
279 2022-03-26 17:25 CLWO DC Filter Check No change.
280 2022-03-26 17:30 CLWO Gain Check No change.
281 2022-03-26 19:30 CLWO Comments End of Service. Sericed by KA. Database entry IA.
282 2021-12-22 16:00 CLWO Service Visit To clear antenna of ice and snow, check power.
283 2021-12-22 18:00 CLWO Comments Too hard to walk alone to the site (KA) with no snow shoes.
284 2021-12-22 18:30 CLWO Power Check Battery volts low, need sun for charging.
285 2021-11-02 21:42 CLWO MT Station Started MT recording, live to London Hub.
286 2021-10-06 12:53 CODG New Station Start Started at London Hub, station Codroy, NF.
287 2021-10-06 12:56 CODG Hardware List Cyg 262, Trd 293, esp seism. T3B19.
288 2021-10-06 12:57 CODG Gain Check Gain 1.0, SPS 100.
No Date Time UT Location Activity Comments
289 2021-10-06 13:02 CODG DC Filter Check Off.
290 2021-10-06 13:02 CODG Require GPS Lock On.
291 2021-10-06 13:03 CODG Comments IP 62. We had trouble with Cygnus 283 at this location. Phase lock timing was not working. Used another cygnus, 262 instead.
292 2021-10-06 13:04 CODG VSAT Check 0/16 dB. Poor.
293 2021-10-06 13:37 CODG SSPB Check H0003A. Not sure if profile is correct.
294 2021-10-06 13:39 CODG Comments 47.84056 N 59.25352 W 50 m (seismo location).
295 2021-10-06 13:40 CODG Comments DC site, 8 solar panels, 6 batteries.
296 2021-10-06 13:40 CODG Site Geology Clay.
297 2021-10-06 13:41 CODG Comments Poor VSAT signal, needed high Erip increase.
298 2021-08-28 14:32 CODG Feed Horn Box Replaced.
299 2021-08-28 14:32 CODG Service Visit To repaire animal damage.
300 2021-08-28 14:33 CODG Cables LNB cable repaired, SSPB cable replaced.
301 2021-08-28 14:33 CODG Carina Swap Switched from SES-AM to Telesat, Ottawa Hub, Carina 116.
302 2021-08-28 14:34 CODG VSAT Check 39 dB isolation!, running hot.
303 2022-04-06 COQB New Station Start AC powered.
304 2021-08-27 COQB Firmware Upgrade Upgraded to Libra V5.20, TS V1.20, TR V1.50 EPLD V1.08
No Date Time UT Location Activity Comments
305 2021-09-24 12:16 COQB Firmware Upgrade Upgraded to Libra Firmware version 5.31.
306 2021-09-25 01:50 COQB DC Filter Change From off to on. At 0.001 mHz corner frequency.
307 2021-09-25 01:50 COQB Gain Change From 1.048576 to 0.1.
308 2022-01-28 18:45 COQB Hardware Swap Cygnus 187 replaced with 269, TRD 77 with 214.
309 2021-06-18 21:45 COQB Service Visit Grounded both Cygnus & Trident. All else ok.
310 2021-11-23 21:55 COQB Comments Station off. VSAT problem??
311 2021-08-03 15:03 COQB Closed End of operation.
312 2021-07-19 23:00 COWN New Station Code Start recording in COWN ringbuffers.
313 2021-08-23 22:03 COWN Firmware Upgrade Upgraded to Libra V5.20, TS V1.20, TR V1.50 EPLD V1.08
314 2021-09-14 19:17 COWN SSPB Check Corrected SSPB temperature profile from flat 45 dB to manufacturer specs.
315 2021-09-22 20:55 COWN Firmware Upgrade Upgraded to Libra Firmware version 5.31.
316 2021-09-22 20:55 COWN Re-Point VSAT antenna repointed, tightened. Wildlife damage repaired.
317 2021-07-06 22:00 COWN Service Visit Site is working. Stopped for service.
318 2021-07-06 23:15 COWN Feed Horn Box New Feed Horn protective box installed. New conduits installed.
319 2021-07-06 23:15 COWN Feed Window New Feed Window installed.
320 2021-07-06 23:15 COWN Firmware Upgrade Libra 5.6201 in both partitions.
No Date Time UT Location Activity Comments
321 2021-07-06 23:15 COWN Re-Point Co-pol 28, X-pol 0 dB.
322 2021-09-30 13:50 COWN Re-Point Due to lack of communication to GE, x-pol/co-pol not available.
323 2021-09-30 13:50 COWN Service Visit Station was down due to shifting of antenna. No wildlife damage observed but additional bird protection (shark teeth) installed on feed horn box.
324 2021-11-22 21:32 COWN Battery Check Low battery volts, 11.92 VDC.
325 2021-07-18 16:30 CTLN New Station Code Started recording at Ottawa Naqs2 in CTLN ringbuffers (temp site was O28T).
326 2021-07-18 16:40 CTLN New Station Code Started recording at Ottawa Naqs1 in CTLN ringbuffers (temp site was O28T).
327 2021-08-27 17:15 DELO DC Filter Change From off to on. At 0.001 mHz corner frequency.
328 2021-08-27 17:15 DELO Gain Check Kept gain at 1.048576.
329 2021-08-27 17:15 DELO New Station Code Start recording in DELO ringbuffers.
330 2021-09-19 18:30 DELO DC Filter Check Kept on at 0.001 mHz corner frequency.
331 2021-09-19 18:30 DELO Gain Change From 1.048576 to 0.1.
332 2021-12-07 19:00 DELO Service Visit To upgrade power system solar panels.
333 2021-06-22 19:51 DREO New Station Start Darlington East, Libra VSAT recording at Ottawa Hub Naqs1.
334 2021-06-22 19:53 DREO VSAT Check 34.5 dB isolation for Telesat.
335 2021-06-22 19:55 DREO Gain Check Software gain 1.0, Vpp 40, at 100 sps.
336 2021-06-22 19:56 DREO DC Filter Check Off.
No Date Time UT Location Activity Comments
337 2021-06-22 19:56 DREO Hardware List Cygnus 802, TrD 1144, seismometer Geotech, 568.
338 2021-06-22 19:56 DREO Require GPS to Lock On.
339 2021-06-22 19:58 DREO Comments N43.87244 w 78.70396 309 FT
340 2021-06-22 19:59 DREO Comments A0017498 sspb, 2 w
341 2021-06-22 20:01 DREO Power Check 16 batteries, 12 solar panels.
342 2021-06-20 20:42 DRWO Internet Station New station, Darlington West, Ontario, Ottawa Naqs1.
343 2021-06-20 20:43 DRWO Internet Station 100 sps, No DC filter, 40 Vpp, Low Impedance, Software Gain 1.
344 2021-06-20 20:45 DRWO Hardware List Taurus 1013, KS-2000M seismometer, S/N 569.
345 2021-06-20 20:46 DRWO Require GPS to Lock On. GPS always on, No Alignment.
346 2021-06-20 20:47 DRWO Comments Darlington Vault.
347 2021-07-18 03:02 DSMN Hardware List CYG396 TRD513 CMG-3ESP T3F12, SPS 40, Gain 1., DC filter On at 1 mHz.
348 2021-07-18 03:02 DSMN New Station Code Start recording in DSMN ringbuffers.
349 2021-07-18 03:02 DSMN New Station Start Start recording in DSMN ringbuffers at DSMN. No data from this location in temp name O78T which was previously allocated to electronics.
350 2021-07-18 03:02 DSMN SSPB Check 2 Watts SSPB, S/N 16463
351 2021-07-18 03:02 DSMN Site Geology Solid Bedrock. Station rock-bolted. Discovery camp nearby, about 300 yards, generator running.
352 2021-07-18 03:02 DSMN VSAT Start Co-pol 35 X-pol 3
No Date Time UT Location Activity Comments
353 2021-07-18 13:28 DSMN DC Filter Check DC filter is turned off.
354 2021-07-18 16:00 DSMN Seismo Check Vertical component calibration signal clipped. Probably due to bad calibration coil. Seismometer locked/unlocked a few times. Calibration appear to unsettle seismometer. DO NOT Calibrate this seismometer until further notice. Noisy test data between 16:00 and 17:30.
355 2021-08-12 13:37 DSMN DC Filter Check Unchanged from Off.
356 2021-08-12 13:37 DSMN Gain Check Software gain unchanged from 1.
357 2021-11-03 14:16 DSMN Firmware Upgrade Trident 1.7210 in A and B. A is active.
358 2021-07-18 20:00 DSMN Service Visit To swap seismometer. Problem loging into Cygnus due to IP conflict (cyg ip same as laptop).
359 2021-07-18 20:28 DSMN Seismo Swap Installed ESP T3J66 removed ESP T3F12(confirm later).
360 2021-07-18 20:30 DSMN Seismo Check Mass centered ok.
361 2021-07-18 20:31 DSMN VSAT Check Repointed. Before co/x 35/10 after 36/5 dB. It is a 2 Watt sspb.
362 2021-07-18 20:49 DSMN Seismo Check Seismo calibration ok. New seismo ok, then.
363 2021-07-18 20:53 DSMN Seismo Check Calibration at 10 sec worked as well as 5 sec cal. All seem to be ok with new seism.
364 2021-07-18 20:55 DSMN Comments Refer to Activity line 539 on July 18, 03 for seismometer history.
365 2021-07-23 16:05 DSMN Closed End of operation.
366 2021-11-09 17:15 DSNO Internet Station Started at deep SNO, Ottawa Hub Naqs1,2.
367 2021-11-09 17:16 DSNO Hardware List Taurus 294. Trillium 120P s/n to be reported.
368 2021-11-09 17:17 DSNO Gain Check Gain 1.0 at 200 sps.
No Date Time UT Location Activity Comments
369 2021-11-09 17:18 DSNO DC Filter Check Off.
370 2021-11-09 17:19 DSNO Require GPS to Lock Off. No GPS at this site.
371 2021-07-26 03:00 DVKN New Station Code Start recording in DVKN ringbuffers.
372 2021-07-26 04:52 DVKN New channel prototype Channel prototype to Broad band, Trident and CMG-40T seismometer (BHZ-2 to BHZ-3, BHN-2 to BHN-3, BHE-2 to BHE-3).
373 2021-08-26 20:31 DVKN Firmware Upgrade Upgraded to Libra V5.20, TS V1.20, TR V1.50 EPLD V1.08
374 2021-09-20 00:20 DVKN Firmware Upgrade Upgraded to Libra Firmware version 5.31.
375 2021-10-24 DVKN VSAT icing fixed. Icing formed on vsat feed window were cleared. Station Started transmitting again.
376 2021-07-08 22:10 DVKN Before Re-Point Before any change to the station, it was noted that the elevation nut on antenna was loose. Before any re-adjustment, co-pol was 28 and x-pol was 5 dB. This could explain bad TX in recent times.
377 2021-07-08 22:10 DVKN Feed Horn Box New Feed Horn protective box installed. New conduits installed.
378 2021-07-08 22:10 DVKN Firmware Upgrade Libra 5.6201 installed. SSPB S/N H0016A.
379 2021-07-08 22:10 DVKN Re-Point After antenna adjustments and re-pointing, co-pol was 38 with x-pol of 2 dB. Much improved.
380 2021-07-08 22:10 DVKN Seismo Check 40T seismometer was off-level probably due to shifting of ground. Re-leveled.
381 2021-07-08 22:10 DVKN Service Visit Site was working generally OK.
382 2021-07-17 14:21 DVKN Closed End of Operation.
383 2021-09-14 EKTN New Station Start First AC station of NWT array.
384 2021-07-18 21:00 EKTN Temp Stop Stopped recording in EKTN ringbuffers to swap electronics.
No Date Time UT Location Activity Comments
385 2021-07-18 21:10 EKTN Hardware Swap Cygnus 156/HRD 207 replaced with Cygnus 222/Trident 143.
386 2021-07-19 00:00 EKTN SR Change Sampling rate changed from 20 sps to 40 sps.
387 2021-07-19 15:00 EKTN Restart Started recording in new ringbuffers at EKTN.
388 2021-09-14 15:30 EKTN SSPB Check Corrected SSPB temperature profile from flat 45 dB to manufacturer specs.
389 2021-09-19 16:30 EKTN Site Visit All ok.
390 2021-09-19 16:31 EKTN Firmware Upgrade Upgraded to Libra Firmware version 5.31.
391 2021-07-13 18:30? EKTN Service Visit Staion was not working OK even after a previous service visit last week. The LNB cable was found to be loose. All cables re-checked.
392 2021-07-13 19:20 EKTN Firmware Upgrade Libra 5.6201, TS 1.23, TRD 1.701 running.
393 2021-07-13 19:20 EKTN Gain Check Software gain 1.0, DC removal filter off, Input range 40 Vpp.
394 2021-07-13 19:20 EKTN Hardware Swap Trident 417 installed, 143 removed. 2 Watts SSPB installed, S/N 16465.
395 2021-07-13 19:20 EKTN Hub Swap Served by Ottawa Hub.
396 2021-07-13 19:20 EKTN Re-Point 24 dB co-pol 0 x-pol.
397 2021-07-13 19:30 EKTN Feed Horn Box New Feed Horn protective box installed. RF cables near feed horn placed in conduit.
398 2021-08-12 13:21 EKTN DC Filter Check Unchanged from Off.
399 2021-08-12 13:21 EKTN Gain Check Software gain unchanged from 1.
400 2021-10-02 19:25 EKTN Hardware Swap Cygnus 222 removed, 392 installed. All else the same.
No Date Time UT Location Activity Comments
401 2021-10-02 19:25 EKTN Service Visit Visit initiated to swap Cygnus.
402 2021-10-02 20:01 EKTN Gain Check Software gain unchanged at 1.0 at 40 sps.
403 2021-11-03 15:12 EKTN Firmware Upgrade Trident 1.7210 in A and B. B is active.
404 2021-07-14 14:04 EKTN Closed End of Operation.
405 2022-02-01 ELAO New Station Start First DC (Solar Panels) station of Ontario array.
406 2022-02-07 ELAO Seismo Swap Changed from CMG-3ESP to CMG-40T.
407 2021-07-23 16:50 ELFO New Station Code New Internet station started, served by UWO hub.
408 2021-07-23 16:51 ELFO Hardware List Cyg 564, Trd 775, ESP seism. T3J88.
409 2021-07-23 16:52 ELFO Gain Check Started with standard gain of 0.5 at 100 sps.
410 2021-07-23 16:53 ELFO DC Filter Check Started with standard Ontario DC filter on at 0.001 Hz.
411 2021-07-23 16:54 ELFO Comments Mass center and calibration ok but not able to log to the station.
412 2021-08-20 21:10 ELFO Comments Cygnus#564 for this Internet node was incorrectly left in TDMA for Ottawa Hub. Removed.
413 2022-01-14 15:00 ELFO Service Visit To install new charge controller and 2 batteries.
414 2021-10-13 18:00 ELFO Internet Station Converted to Taurus Internet at old location, served by Ottawa Hub.
415 2021-10-13 18:10 ELFO Hardware List Taurus 121, ESP T3J88 (as before).
416 2021-10-13 18:20 ELFO Gain Change Gain is 1.0 (was 0.5 before) at 100 sps.
No Date Time UT Location Activity Comments
417 2021-10-13 18:30 ELFO DC Filter Check On at 0.001 Hz.
418 2021-10-13 18:40 ELFO Require GPS to Lock On.
419 2021-10-13 18:50 ELFO Internet Station On VPN.
420 2022-02-08 ELGO Name Change Ringbuffer and station code changed from ELAO to ELGO.
421 2022-02-22 ELGO Seismo Swap Changed from CMG-40T to CMG-3ESP.
422 2021-08-21 19:18 ELGO Firmware Upgrade Upgraded to Libra V5.20, TS V1.20, TR V1.50 EPLD V1.08
423 2021-09-17 02:20 ELGO DC Filter Change From off to on. At 0.001 mHz corner frequency.
424 2021-09-17 02:20 ELGO Gain Change From 1.048576 to 0.1.
425 2021-09-24 09:22 ELGO Firmware Upgrade Upgraded to Libra Firmware version 5.31.
426 2021-11-15 20:37 ELGO Comments Rebooted Time Server to fix GPS SAT not showing problem.
427 2021-11-15 20:40 ELGO Comments Now showing Instrument, Timing and GPS Sat SOH correctly. Needed a reboot.
428 2022-04-28 12:55 ELGO Closed End of Operation.
429 2022-04-04 ENGB New Station Start AC powered.
430 2021-08-14 20:12 ENGB Firmware Upgrade Upgraded to Libra V5.20, TS V1.20, TR V1.50 EPLD V1.08
431 2021-09-12 09:36 ENGB DC Filter Change From off to on. At 0.001 mHz corner frequency.
432 2021-09-12 09:36 ENGB Gain Change From 1.048576 to 0.1.
No Date Time UT Location Activity Comments
433 2021-09-23 20:14 ENGB Firmware Upgrade Upgraded to Libra Firmware version 5.31.
434 2021-06-17 18:34 ENGB Gain Change Software gain is set to 1.0 now but sps kept at 100 as before.
435 2021-06-17 18:34 ENGB Hardware Swap Installed Cyg398, Trd418, kept seismo the same, T3B20.
436 2021-06-17 18:34 ENGB Hub Swap Station is served by Ottawa Hub.
437 2021-08-11 18:31 ENGB DC Filter Check Unchanged from ON at 0.001 Hz.
438 2021-08-11 18:31 ENGB Gain Change Software gain changed from 1.0 to 0.5.
439 2021-11-03 14:53 ENGB Firmware Upgrade Trident 1.7210 in A and B. B is active.
440 2022-03-02 17:06 ENGB Service Visit Stopped station for maintenance visit.
441 2022-03-02 17:13 ENGB VSAT Check X/Co: 0/27 dB.
442 2022-03-02 17:17 ENGB Restart Station restarted, all OK.
443 2021-08-16 17:34 ENGB Comments GPS sats were not visible, but system worked.
444 2021-08-16 17:34 ENGB Firmware Upgrade Upgrade to Libra V57005B and TS2000.
445 2021-08-16 17:34 ENGB Service Visit To figure out the gap problem.
446 2021-08-16 17:36 ENGB Comments GPS stas are visible now.
447 2021-08-16 17:36 ENGB Comments No change to Trident firmware.
448 2021-08-16 17:37 ENGB Cold-Boot Standard procedure.
No Date Time UT Location Activity Comments
449 2021-08-16 17:37 ENGB VSAT Check X/Co 0/22 dB. Unchaged but checked.
450 2021-08-16 17:47 ENGB Gain Check Unchanged at 0.5.
451 2021-08-16 17:48 ENGB Comments Firmware upgrade appear to have helped.
452 2021-08-16 17:48 ENGB DC Filter Check Unchaged from On at 1 mHz.
453 2021-08-30 17:14 ENGB Closed End of operation.
454 2021-06-11 18:33 EPLO New Station Start Started on Naqs2, Ottawa Hub, in EPLO buffers.
455 2021-06-11 18:34 EPLO New Station Start Started on Naqs1, Ottawa Hub, in EPLO buffers.
456 2021-06-18 18:03 EPLO VSAT Check Increased throughput from 9000 to 12000 in TDMA.
457 2021-06-18 18:03 EPLO VSAT Check ReTX too high, getting too many bad packets. Why?
458 2021-06-18 18:10 EPLO VSAT Check RX: good/bad packes is now 1075/143.
459 2021-06-18 18:20 EPLO VSAT Check RX: good/bad packes is now 1782/841. In 10 mins, we had about 700 bad packets (but about 10 bad bursts only).
460 2021-06-18 18:24 EPLO VSAT Check TDMA Eirp was reduced by 2 dB.
461 2021-06-18 18:46 EPLO VSAT Check Too many bad bursts. TDMA Erip was increased by 6 dB. Now it is at 8 dB, 3 dB higher than original value before this set of tests.
462 2021-06-18 18:54 EPLO VSAT Check Far too many bad bursts. Eirp was reduced to 0.0 dB.
463 2021-06-18 19:07 EPLO VSAT Check Giving up on EPLO. VSAT behaving badly.
464 2021-07-18 11:30 FEDNOR Comments Started field work for installation of FedNor 3 stations.
No Date Time UT Location Activity Comments
465 2021-06-18 21:58 FEDNOR2 Comments Team just finished the last of FedNor2 installation. On their way home CA, JD, MP, CB. Already home: AS, David and Coleman and IA.
466 2021-08-24 20:06 FLLB New Station Code Fletcher Lake Big Creek, Nechako Basin, BC, Ottawa Hub Naqs1,2, Telesat.
467 2021-08-24 20:07 FLLB Hardware List Cyg266, Trd76, 3ESP T3D84, redeploed from TSJB.
468 2021-08-24 20:08 FLLB DC Filter Check Off.
469 2021-08-24 20:08 FLLB Gain Check Gain of 1 at 100 sps.
470 2021-08-24 20:09 FLLB Battery Check 6 new batteries, 8 solar panels.
471 2021-08-24 20:09 FLLB Require GPS Lock On.
472 2021-08-24 20:09 FLLB Site Geology River bed deposits.
473 2021-08-24 20:09 FLLB VSAT Check 29.5 dB isolation.
474 2021-08-24 20:10 FLLB Cold-Boot Was done before all this.
475 2021-06-26 17:46 FLLB Comments Mass-centered OK.
476 2021-06-26 17:46 FLLB Service Visit Replacing bad seismometer.
477 2021-06-26 17:47 FLLB Cold-Boot OK.
478 2021-06-26 18:21 FLLB Comments Calibration OK.
479 2021-06-26 18:23 FLLB Hardware Swap T3D84 removed, T3K45 installed.
480 2021-06-26 18:24 FLLB Hardware Swap Seismometer cable changed due to condensation in connector.
No Date Time UT Location Activity Comments
481 2021-06-26 18:25 FLLB Comments All is OK.
482 2021-06-26 18:25 FLLB Comments Padlock changed.
483 2021-06-26 18:25 FLLB VSAT Check 30 dB, no change.
484 2021-06-26 18:26 FLLB Gain Check Gain 1 at 40 Vpp 100 sps.
485 2021-06-26 18:29 FLLB DC Filter Check Off.
486 2021-07-20 20:24 GALN New Station Code Started on Naqs2, Ottawa Hub in GALN ringbuffers.
487 2021-07-20 20:29 GALN New Station Code Started on Naqs1, Ottawa Hub in GALN ringbuffers.
488 2021-10-25 14:42 GASG New Station Start Gaspe, Quebec, on Telesat, recording at Ottawa Hub.
489 2021-10-25 14:47 GASG Hardware List Cyg 419, Trd 296, seismo esp T3D85, SSPB H001a 1w, TS 547.
490 2021-10-25 14:48 GASG Comments 48.94626 N 66.11613 W 855 ft.
491 2021-10-25 14:49 GASG Site Geology Glacial.
492 2021-10-25 14:50 GASG Comments Movement of material in nearby gravel pit for a couple of days.
493 2021-10-25 14:50 GASG Gain Check 1.0 at 100 sps.
494 2021-10-25 14:51 GASG DC Filter Check Off.
495 2021-10-25 14:51 GASG Require GPS Lock On.
496 2021-10-25 14:51 GASG VSAT Check 33 dB isolation on Telesat.
No Date Time UT Location Activity Comments
497 2021-10-25 14:52 GASG Power Check DC site on 6 batteries and 10 panels, but limited view of sun.
498 2021-09-08 18:00 GASG Service Visit Routine maintenance.
499 2021-09-08 18:10 GASG VSAT Check Repointing gained 2 dB, isolation is 40 dB.
500 2021-09-24 21:28 GBLN DC Filter Check Started with DC filter off.
501 2021-09-24 21:28 GBLN Gain Check Started with software gain of 1.0 at 40 sps.
502 2021-09-24 21:28 GBLN New Station Start Start recording in GBLN ringbuffers.
503 2021-11-03 20:23 GBLN Firmware Upgrade Trident 1.7210 in A and B. B is active.
504 2021-08-24 15:00 GBLN Service Visit To bring the station up and install MT.
505 2021-08-24 15:01 GBLN Wildlife Tarps missing on battery boxes, one shredded nearby suspect animal attack, big hole in feed window.
506 2021-08-24 15:30 GBLN Repair Feed window replaced with a reconditioned one.
507 2021-08-24 15:32 GBLN VSAT Check Co/x improved from 17/0 to 20/0 dB.
508 2021-08-24 15:33 GBLN Seismo Swap ESP T3K45 installed, 40T T4316 removed.
509 2021-08-24 15:34 GBLN Restart Ottawa Hub Naqs2 restarted.
510 2021-08-24 15:35 GBLN Restart Ottawa Hub Naqs1 restarted.
511 2021-08-24 15:47 GBLN Comments Both Naqs1 and Naqs2 of Ottawa Hub ready for MT.
512 2021-08-24 15:50 GBLN Comments Calibration and Mass-centre ok.
No Date Time UT Location Activity Comments
513 2021-08-24 16:39 GBLN Comments MT installation in progress.
514 2021-07-24 15:30 GBLN Service Visit Seismo vault tipped on its side. MT equipment damaged but may be fixable. Repair in progress.
515 2021-07-24 15:40 GBLN Hardware Swap TRD 143 removed, may be damagd.
516 2021-07-24 15:40 GBLN Hardware Swap Trd 513 installed.
517 2021-07-24 16:00 GBLN Seismo Check Three attempts failed to mass-center seismometer. May be damaged.
518 2021-07-24 16:30 GBLN Seismo Swap Bad (damaged?) seismometer T3K45 replaced. ESP T3J66 installed.
519 2021-07-24 16:55 GBLN Seismo Check New seismometer responded ok to mass-center and calibration.
520 2021-07-24 16:57 GBLN Comments MT repair in progress.
521 2021-07-25 17:34 GBLN Service Visit Site was down but nothing was broken. Feed horn was replaced.
522 2021-07-25 17:40 GBLN Comments Raining at the site.
523 2021-07-25 17:40 GBLN SSPB Check New SSPB is A0176A.
524 2021-07-25 17:40 GBLN VSAT Check 25/0 dB.
525 2021-07-25 17:41 GBLN Comments Sensor calibration ok.
526 2021-10-17 12:33 GBN New Station Start Started new Atlantic station at Guysborough, recording at London Hub.
527 2021-10-17 12:34 GBN Comments Dataserver restarted on London Hub, Naqs1.
528 2021-10-18 19:29 GBN Hardware List Cygnus 535, Trident 214, Carina 29, IP 39, Seismo ESP, s/n T3D80, sspb H0002A.
No Date Time UT Location Activity Comments
529 2021-10-18 19:33 GBN DC Filter Check Off.
530 2021-10-18 19:33 GBN Gain Check Gain 1.0, SPS 100.
531 2021-10-18 19:33 GBN Require GPS Lock On.
532 2021-10-18 19:34 GBN Site Geology Solid bedrock.
533 2021-10-18 19:34 GBN VSAT Check 26/0 dB.
534 2021-10-18 19:35 GBN Comments This is AC site with NPM mount bolted to bedrock.
535 2021-08-24 18:10 GBN Comments Switched from SES-AM to Telesat.
536 2021-08-24 18:11 GBN VSAT Check At least 28 dB isolation, very strong signal.
537 2021-08-19 21:48 GIFN New Station Start Gifford Fjord, Baffin Island, recording at Naqs2, Ottawa.
538 2021-08-19 21:52 GIFN New Station Start Gifford Fjord, Baffin Island, recording at Naqs1, Ottawa.
539 2021-08-19 21:53 GIFN Hardware List Cyg 609 Trd 848 sspb a0002a, Seismometer esp T3p20.
540 2021-08-19 21:54 GIFN Gain Check Gain of 1.0, sps 40.
541 2021-08-19 21:55 GIFN DC Filter Check DC filter is not on.
542 2021-08-19 21:56 GIFN VSAT Check 30 dB of isolation on Telesat.
543 2021-08-19 21:57 GIFN Site Geology Sand.
544 2021-08-22 14:56 GIFN Require GPS Lock Require GPS Lock is on.
No Date Time UT Location Activity Comments
545 2021-08-22 14:57 GIFN Seismo Check New ESP type with auto mass lock.
546 2021-07-23 17:00 GLWN New Station Code Start recording in GLWN ringbuffers. 2021-07-24 16:00 P25T New Station Start Start recording in P25T ringbuffers at YNEN with CMG-40T.
547 2021-08-12 21:25 GLWN Firmware Upgrade Upgraded to Libra V5.20, TS V1.20, TR V1.50 EPLD V1.08
548 2021-09-23 18:01 GLWN Repair Snowstorm at the site. RF cables repaired.
549 2021-09-25 17:10 GLWN SSPB Check Corrected SSPB temperature profile from flat 45 dB to manufacturer specs.
550 2021-07-06 16:22 GLWN Feed Horn Box New Feed Horn protective box installed. New conduits installed.
551 2021-07-06 16:22 GLWN Feed Window Small bite in Feed Window repaired.
552 2021-07-06 16:22 GLWN Firmware Upgrade Libra 5.6201, TS 1.23, Trd 1.7201, EPLD 1.23
553 2021-07-06 16:22 GLWN Hardware Swap Cyg 251replaced with Cyg421. All other hardware the same.
554 2021-07-06 16:22 GLWN Hub Swap Served by Ottawa Hub.
555 2021-07-06 16:22 GLWN Re-Point Co-pol 20 dB. Low signal.
556 2021-07-06 16:22 GLWN Service Visit Could not communicate with Cygnus 251. Something wrong with Cygnus. Need to be checked later. SSPB was loose.
557 2021-08-12 11:55 GLWN DC Filter Check Unchanged from Off.
558 2021-08-12 11:55 GLWN Gain Change Software gain changed from 1.048576 to 1.0.
559 2021-11-03 21:02 GLWN Firmware Upgrade Trident EPLD 1.23 installed in B.
560 2021-11-03 21:03 GLWN Firmware Upgrade Trident 1.7210 in A and B. A is active.
No Date Time UT Location Activity Comments
561 2022-04-10 GOWB New Station Start AC powered.
562 2021-08-13 20:11 GOWB Firmware Upgrade Upgraded to Libra V5.20, TS V1.20, TR V1.50 EPLD V1.08
563 2021-09-12 10:12 GOWB DC Filter Change From off to on. At 0.001 mHz corner frequency.
564 2021-09-12 10:12 GOWB Gain Change From 1.048576 to 0.1.
565 2021-09-20 22:12 GOWB Firmware Upgrade Upgraded to Libra Firmware version 5.31.
566 2021-11-05 18:45 GOWB Service Visit Station was going down repeatedly. Visit to find the problem.
567 2021-11-05 18:46 GOWB VSAT Check Before repoint and after repoint attemp 27/0.
568 2021-11-05 18:47 GOWB Comments No problem found at the site.
569 2021-08-19 17:03 GOWB Closed End of operation.
570 2021-05-29 21:00 HBDB Hardware List Guralp ESP T3A81?? Cygnus 394 Trident 425
571 2021-05-29 21:00 HBDB Site Geology Soft Clay Soil
572 2021-05-29 21:00 HBDB VSAT Start X-pol 2, Co-pol 34 dB.
573 2021-05-29 21:?? HBDB New Station Code Start recording in HBDB ringbuffers.
574 2021-12-19 18:44 HBDB Service Visit Station shut down to swap electronics.
575 2021-12-19 18:45 HBDB Hardware List Old hardware to be removed: CYG394 TRD425
576 2021-12-19 19:20 HBDB Hardware Swap CYG394 TRD425 removed, CYG161 TRD064 installed
No Date Time UT Location Activity Comments
577 2021-12-19 20:01 HBDB Restart Back on air.
578 2021-12-19 20:02 HBDB Hub Swap Moved to Ottawa Hub.
579 2021-12-19 20:10 HBDB DC Filter Check Unchanged, On at 1 mHz.
580 2021-12-19 20:12 HBDB Gain Check Unchanged at 0.5.
581 2021-12-19 20:13 HBDB Comments IP is at 161 on Hub 5.
582 2022-03-02 22:13 HBDB Service Visit Stopped station for maintenance visit.
583 2022-03-02 22:15 HBDB VSAT Check X/Co. 2/32 dB.
584 2022-03-02 22:18 HBDB Cold-Boot Restarted OK.
585 2021-07-31 18:56 HBDB Hardware Swap Installed Cyg 562, TS 711, Trd 760. Removed Cyg 161, Trd 064.
586 2021-07-31 18:57 HBDB DC Filter Check Unchaged at on at 0.001 Hz.
587 2021-07-31 18:57 HBDB Gain Check Unchaged at 0.5.
588 2021-07-31 18:58 HBDB Comments Condition of the site was ok at the time of visit.
589 2021-07-31 18:58 HBDB Comments Ringbuffers restarted.
590 2021-07-31 19:10 HBDB Comments Switched O30T to HBDB in naqs.stn files on Ottawa Naqs2 and Naqs1. Station currently down for final adjustments.
591 2021-07-31 19:13 HBDB Comments Cold boot. Ground slump near the base on the King Post. May be problematic.
592 2021-07-28 14:58 HBDB Closed End of operation.
No Date Time UT Location Activity Comments
593 2022-02-04 HGVO New Station Start DC (Solar Panels).
594 2021-08-21 HGVO Firmware Upgrade Upgraded to Libra V5.20
595 2021-09-25 03:00 HGVO DC Filter Change From off to on. At 0.001 mHz corner frequency.
596 2021-09-25 03:00 HGVO Gain Change From 1.048576 to 0.1.
597 2022-04-15 HGVO Power Upgrade Added 2 more batteries.
598 2022-04-15 HGVO Site Visit Maintenance visit.
599 2021-08-27 16:30 HGVO DC Filter Check Unchanged at 0.001 Hz and on.
600 2021-08-27 16:30 HGVO Gain Check Unchanged at 0.5, sample rate 100.
601 2021-08-27 16:30 HGVO Hardware Swap CYG175, TRD073 removed, CYG173, TRD059 installed.
602 2021-08-27 16:30 HGVO Service Visit
603 2021-10-22 16:15 HGVO Seismo Swap ESP-T3B16 Removed. 40T-T4770 insatalled.
604 2021-11-10 15:57 HGVO Temp Halt Cyg, Tri, seismo returned to UWO for repair.
605 2021-11-18 16:00 HGVO Service Visit New hardware, trying to find source of problem.
606 2021-11-18 16:10 HGVO Cables Bus cable checks ok. Not replaced. Grounding ok.
607 2021-11-18 16:20 HGVO Hardware List CYG422 TIM528 TRD237 T3B16(ESP)
608 2021-11-18 16:21 HGVO VSAT Check 35/1.5.
No Date Time UT Location Activity Comments
609 2021-11-18 16:24 HGVO Hub Swap Served by Ottawa Hub.
610 2021-11-18 16:25 HGVO Restart Restarted with new hardware.
611 2021-11-18 16:26 HGVO SR Change Incorrect at 40 sps.
612 2021-11-18 16:35 HGVO Gain Check Unchanged at 0.5.
613 2021-11-18 16:36 HGVO DC Filter Check Unchaged at on and 1 mHz.
614 2021-11-18 16:37 HGVO SR Change Corrected from 40 to 100 sps.
615 2021-11-18 17:01 HGVO Comments VSAT log in ok at ip 10.6.255.33
616 2021-11-18 17:02 HGVO Battery Check DC volts ok at 13.5
617 2021-11-21 14:01 HGVO VSAT Check VSAT power level increased by 2 dB at Ottawa Hub.
618 2021-12-01 12:39 HGVO VSAT Check Hub power level increased from 5 to 7 dB, throughput from 8000 to 11000.
619 2021-12-01 12:40 HGVO Comments Hub tweaked to improve gap situation.
620 2021-12-19 15:00 HGVO Service Visit Station down about 1/2 hour to swap bus cable.
621 2021-12-19 16:00 HGVO Cables Bus cable replaced to fix gap problems.
622 2022-04-29 19:00 HGVO Closed End of Operation.
623 2021-11-10 19:00 HLSB New Station Code Start recording in HLSB ringbuffers.
624 2022-03-02 19:44 HLSB Service Visit Stopped station for maintenance visit.
No Date Time UT Location Activity Comments
625 2022-03-02 19:55 HLSB VSAT Check X/Co: 5/34 dB.
626 2022-03-02 19:58 HLSB Cold-Boot Restarted OK.
627 2022-03-02 20:00 HLSB DC Filter Check Unchanged at On.
628 2022-03-02 20:00 HLSB Gain Check Unchanged at 0.5.
629 2021-06-09 18:30 HLSB Service Visit To investigate VSAT problems. No damage to site but tall grass in view of antenna. Will take pics.
630 2021-06-09 18:31 HLSB VSAT Check No signal from station. Increased VSAT power at hub by 2 dB and did cold boot on station. No signal yet.
631 2021-06-09 18:37 HLSB VSAT Check Increased VSAT by another 5 dB (in total, changed from -7dB to 0 dB). No change. Asked operator to check all RF cables and VSAT.
632 2021-06-09 18:40 HLSB VSAT Check RF cable appears to have been damaged. Moving cable a little helped and station came back up. Shutting down to fix cable.
633 2021-06-09 18:50 HLSB VSAT Check SSPB cable has small scratch on the tape at the sspb end. Suspecting damage to feed horn. Plan to re-point the antenna.
634 2021-06-09 19:26 HLSB VSAT Check Replace feed horn. Contacted GE. They report signal is trash. Suspect bad Cygnus.
635 2021-06-09 19:50 HLSB Temp Halt Shut down to replace Cygnus. Need to bring a Cygnus from test site at PGC tomorrow.
636 2021-06-10 19:00 HLSB VSAT Check Co/X pol 25/0 dB.
637 2021-06-10 19:40 HLSB Comments New feedhorn installed with new rf cables.
638 2021-06-10 19:40 HLSB Hardware Swap Cygnus 217 removed, 172 installed to fix VSAT problem.
639 2021-06-10 19:41 HLSB Comments Station operation on London Carina 3 now.
640 2021-06-10 19:45 HLSB Hardware Swap New SSPB H0018A installed. Temperature profile needs to be updated.
No Date Time UT Location Activity Comments
641 2021-06-10 19:46 HLSB Cold-Boot Done.
642 2021-06-10 19:47 HLSB Comments Cygnus was from test site at PGC, was configured for GPS on serial port. Serial port was removed.
643 2021-06-10 19:50 HLSB DC Filter Check On at 0.001 Hz. Unchanged.
644 2021-06-10 19:51 HLSB Gain Check Unchanged at 0.5.
645 2021-07-26 18:25 HLSB Closed End of operation.
646 2021-11-12 02:35 HRMB New Station Code Start recording in HRMB ringbuffers.
647 2022-03-02 20:58 HRMB Service Visit Stopped station for maintenance visit.
648 2022-03-02 21:00 HRMB VSAT Check X/Co. 3/33 dB.
649 2022-03-02 21:08 HRMB Cold-Boot Restarted OK.
650 2022-03-02 21:10 HRMB DC Filter Check Unchaged from On at 1 mHz.
651 2022-03-02 21:10 HRMB Gain Check Unchanged at 0.5.
652 2021-07-27 15:25 HRMB Closed End of operation.
653 2021-06-17 HSLB Service Visit
654 2021-07-20 18:50 HSMO New Station Code New Station started on both Ottawa Hub NaqsServer. Switched from O64T to HSMO.
655 2021-07-20 19:28 HSMO SSPB Check SSPB profile corrected. All modem parameters checked, were ok. Screen captures done. Leaving the site.
656 2021-09-02 21:08 IGLN New Station Code Started recording at Ottawa Naqs2 in IGLN ringbuffers. Served by Telesat Anik E2R.
No Date Time UT Location Activity Comments
657 2021-09-02 21:09 IGLN DC Filter Check Started with DC filter off.
658 2021-09-02 21:09 IGLN Gain Check Started with gain of 1.0 at 40 sps.
659 2021-09-02 21:10 IGLN VSAT Check Co/X at 31/1 dB.
660 2021-09-02 21:17 IGLN Seismo Check Mass Center and Calibration OK.
661 2021-09-02 21:18 IGLN Hardware List Cyg 284, Trd 770, ESP T3K44.
662 2021-09-02 21:53 IGLN VSAT Check Operating with BPSK, 32 KB. Eirp increase 6 dB, bad/good packet ration 118/6792.
663 2021-09-02 21:57 IGLN New Station Code Started recording at Ottawa Naqs1 in IGLN ringbuffers.
664 2021-09-02 22:03 IGLN VSAT Check Bad/good ratio 156/7703. Eirp increased by 1 dB.
665 2021-10-06 16:40 IGLN Service Visit To upgrade SSPB from 1 W to 2 Watt and re-point dish.
666 2021-10-06 16:41 IGLN SSPB Check 2 Watt installed, 1 Watt removed.
667 2021-10-06 16:42 IGLN VSAT Check Isolation 30.0 dB. Eirp increase on 7 dB on Hub 9.
668 2021-10-06 16:45 IGLN Gain Check Unchaged at 1 and 40 sps.
669 2021-10-06 16:46 IGLN DC Filter Check Unchaged at no DC filter.
670 2021-06-18 00:01 IGLN Restart Station is back after down during winter months due to snow blcking the VSAT antenna.
671 2021-06-20 16:18 IGLN Seismo Check BHE component off-center. Signal on all channels look odd.
672 2021-08-22 20:16 IGLN Closed Moved, due to flooding of seismometer vault, to ILON.
No Date Time UT Location Activity Comments
673 2021-07-31 23:25 IHLN New Station Code Start recording in IHLN ringbuffers.
674 2021-08-13 18:35 IHLN Firmware Upgrade Upgraded to Libra V5.20, TS V1.20, TR V1.50 EPLD V1.08
675 2021-09-14 21:50 IHLN SSPB Check Corrected SSPB temperature profile from flat 45 dB to manufacturer specs.
676 2021-09-21 16:40 IHLN Firmware Upgrade Upgraded to Libra Firmware version 5.31.
677 2021-07-19 17:28 IHLN Feed Horn Box New Feed Horn protective box installed. New cables and conduits installed.
678 2021-07-19 17:28 IHLN Feed Window Broken, replaced.
679 2021-07-19 17:28 IHLN Firmware Upgrade Libra 5.6201 installed.
680 2021-07-19 17:28 IHLN Re-Point Co-pol 25, x-pol 0 dB.
681 2021-07-19 17:28 IHLN Service Visit Site was working generally OK. Stopped for service.
682 2021-07-19 19:38 IHLN Restart Calibrated OK.
683 2021-07-18 15:42 IHLN Closed End of operation.
684 2021-09-27 23:54 ILKN DC Filter Check Started with DC filter off.
685 2021-09-27 23:54 ILKN Gain Check Started with software gain of 1.0 at 40 sps.
686 2021-09-27 23:54 ILKN New Station Start Start recording in ILKN ringbuffers.
687 2021-11-03 21:21 ILKN Firmware Upgrade Trident 1.7210 in A and B. A is active.
688 2021-07-18 16:00 ILKN VSAT Check Co/x was unchanged at 27/0. No re-point needed.
No Date Time UT Location Activity Comments
689 2021-07-18 16:19 ILKN Gain Check Unchanged at 1.0.
690 2021-07-18 16:21 ILKN DC Filter Check Unchanged off.
691 2021-07-18 16:35 ILKN Cold-Boot Station started after cold-boot test.
692 2021-07-18 17:50 ILKN Service Visit To check VSAT. Site was working before. All apper ok.
693 2021-08-23 21:14 ILON Comments Old code of IGLN is retired due to change of location.
694 2021-08-23 21:17 ILON Hardware List CYG620, TRD871, ESP Seismo T3P18 (auto-mass-lock)
695 2021-08-23 21:18 ILON DC Filter Check DC filter is not on.
696 2021-08-23 21:18 ILON Gain Check Gain of 1.0, sps 40.
697 2021-08-23 21:18 ILON Site Geology Solid bedrock a couple of feet below seismometer pier.
698 2021-08-23 21:19 ILON Seismo Check Waiting for pier to dry out, seismometer not connected yet.
699 2021-08-23 21:20 ILON Require GPS Lock Require GPS Lock is on.
700 2021-08-23 21:20 ILON VSAT Check 30 dB of isolation on Telesat.
701 2021-08-23 23:00 ILON New Station Start AT new location, Igloolik Airport Weather Station, recording at Ottawa Hub Naqs2 and Naqs1.
702 2021-08-24 13:30 ILON Seismo Swap T3P18 has bad Z component, removed, T3K44 installed.
703 2021-08-24 13:49 ILON Comments End of new installation.
704 2021-09-12 20:00 INUQ New Station Code Internet, Ottawa Hub, at Inukjuak, Hudson Bay.
No Date Time UT Location Activity Comments
705 2021-09-12 20:05 INUQ Internet Station On VPN.
706 2021-09-12 20:10 INUQ Hardware List Taurus 270, ESP T3B16, IP 86.34
707 2021-09-12 20:20 INUQ 40 sps, Gain 1, No DC filter, Require GPS Lock Checked.
708 2021-09-16 22:00 IVKQ New Station Code Internet, Ottawa Hub, at Ivujivik, Hudson Bay.
709 2021-09-16 22:10 IVKQ Hardware List Taurus 201, ESP T3P18.
710 2021-09-16 22:20 IVKQ 40 sps, Gain 1, No DC filter, Require GPS Lock Checked.
711 2021-09-16 22:25 IVKQ Internet Station On VPN.
712 2021-07-21 23:30 JERN New Station Code LUPN moved to this location. Station started with same hardware but on Telesat.
713 2021-07-22 01:05 JERN Seismo Check Mass-center and calibration ok.
714 2021-07-24 13:58 JERN Gain Check Gain is set to 1.0
715 2021-07-24 13:58 JERN Hardware List CYG389 TRD437 CMG-3ESP T3E89.
716 2021-07-24 13:59 JERN DC Filter Check Off.
717 2021-07-24 13:59 JERN Require GPS Lock On.
718 2021-07-24 13:59 JERN VSAT Check Isolation 30 dB. On Telesat Hub 9.
719 2021-07-24 14:00 JERN Comments Operation at 40 SPS.
720 2021-08-25 11:50 JOSN New Station Code Second Sedna station, started at Ottawa Hub, Naqs1,2, Telesat.
No Date Time UT Location Activity Comments
721 2021-08-25 11:52 JOSN Hardware List Cyg392, Trd141, TS 373, 40T T47773, SSPB A0208A 1W.
722 2021-08-25 11:54 JOSN DC Filter Check Off.
723 2021-08-25 11:54 JOSN Gain Check 1.0 at 40 sps.
724 2021-08-25 11:54 JOSN Require GPS Lock On.
725 2021-08-25 11:55 JOSN Comments All hardware except for Cygnus redeployed from EKTN.
726 2021-08-25 11:55 JOSN VSAT Check 31 dB isolation.
727 2021-08-03 15:17 KASO New Station Code O54T is now recording in KASO ringbuffers at Ottawa Hub Naqs2.
728 2021-08-03 15:21 KASO New Station Code O54T is now recording in KASO ringbuffers at Ottawa Hub Naqs1.
729 2021-07-10 21:41 KASO Service Visit Just to adjust VSAT.
730 2021-07-10 22:18 KASO Comments Service continues.
731 2021-07-11 00:46 KASO VSAT Check Isolation improved from 10 to 16 dB.
732 2021-07-11 00:47 KASO Comments Feedhorn arms were upside down, corrected.
733 2021-07-11 00:49 KASO Comments General condition of site was fine.
734 2022-04-08 KELB New Station Start AC powered.
735 2021-08-23 22:17 KELB Firmware Upgrade Upgraded to Libra V5.20, TS V1.20, TR V1.50 EPLD V1.08
736 2021-09-06 16:30 KELB DC Filter Change From off to on. At 0.001 mHz corner frequency.
No Date Time UT Location Activity Comments
737 2021-09-06 16:30 KELB Gain Change From 1.048576 to 0.5. Vpp unchanged at 40.
738 2021-09-12 09:25 KELB DC Filter Change Unchanged from on at 0.001 mHz corner frequency.
739 2021-09-12 09:25 KELB Gain Change From 0.5 to 0.1. Vpp unchanged at 40.
740 2021-09-23 22:25 KELB Firmware Upgrade Upgraded to Libra Firmware version 5.31.
741 2021-06-22 13:04 KILO Hardware List CYG397 TRD369 CMG-40T T4168. SPS 100, Gain 1.
742 2021-06-22 13:04 KILO New Station Code Start recording in KILO ringbuffers.
743 2021-06-22 13:04 KILO Site Geology Bedrock
744 2021-06-22 13:04 KILO VSAT Start x/c ??
745 2021-08-11 19:46 KILO DC Filter Check Unchanged from ON at 0.001 Hz.
746 2021-08-11 19:46 KILO Gain Check Software gain unchanged from 1.0.
747 2021-08-12 19:20 KILO Seismo Check Confirmed to be T4168.
748 2021-10-30 16:08 KILO Service Visit Station working ok but Trident did not connect after reboot.
749 2021-10-30 16:09 KILO Wildlife Live and dead mouce in seismo vault.
750 2021-10-30 16:10 KILO Comments First time Trident failed to connect due to bad 1.7201 firmware.
751 2021-10-30 16:18 KILO Hardware Swap Trident 516 installed, 369 removed.
752 2021-10-30 17:35 KILO SR Change Incorrectly set to 40 sps.
No Date Time UT Location Activity Comments
753 2021-10-30 17:36 KILO VSAT Check Before 30/4 now 30/1 dB.
754 2021-10-30 17:37 KILO Gain Check Remained at 1, as before.
755 2021-10-30 17:38 KILO DC Filter Check Remained on at 1 mHz as before.
756 2021-10-30 17:40 KILO Repair Seismo vault sealed.
757 2021-11-01 13:55 KILO SR Change Restored from 40 to 100 sps.
758 2021-11-01 13:56 KILO Firmware Upgrade Trident 1.7210 in B.
759 2021-11-01 14:23 KILO Firmware Upgrade Trident 1.7210 in A and active.
760 2021-08-19 13:11 KILO Service Visit To upgrade seismometer.
761 2021-08-19 13:12 KILO Seismo Swap ESP T3D69 installed, 40T T4186 removed.
762 2021-08-19 14:25 KILO Cold-Boot Standard procedure.
763 2021-08-19 14:26 KILO Restart With new seismometer, recording at Ottawa Hub on Naqs2.
764 2021-08-19 14:27 KILO Gain Check Unchanged at 1.
765 2021-08-19 14:28 KILO DC Filter Check Unchanged at on and 0.001 Hz.
766 2021-08-19 14:29 KILO VSAT Check Co/x at 20/0 dB, unchanged.
767 2021-08-19 14:47 KILO Seismo Check Vertical component of new seismometer appeared to have been unlocked during transportation. However, all channels mass-centred ok and calibrated ok. Calibration singals at 1, 5 and 10 seconds around 14:47 UT look normal.
768 2021-08-19 14:50 KILO Comments Mouse droppings in seismo vault.
No Date Time UT Location Activity Comments
769 2021-08-19 14:51 KILO Comments Seismo may have a bad bubble, used external bubble to level. Inner vault cover was not installable but insolation covers installed. Should be adequate.
770 2021-08-19 14:52 KILO Comments Black bears spotted nearby but not bothering the site.
771 2021-08-19 15:13 KILO Comments Naqs1 restarted with new seismomter.
772 2021-07-28 13:45 KLBO New Station Code Start recording in KLBO ringbuffers.
773 2021-08-25 20:36 KLBO Service Visit Station was not transmitting, Cygnus found in Test mode.
774 2021-08-25 21:30 KLBO Seismo Swap CMG-40T T4567 removed, 40T T4317 installed.
775 2021-08-25 21:30 KLBO VSAT Start Co-pol 29, X-pol 0.
776 2021-08-26 12:25 KLBO Service Visit Stopped to replace Trident due to noise on n/s comp.
777 2021-08-26 12:40 KLBO DC Filter Check Unchanged at 0.001 Hz and on.
778 2021-08-26 12:40 KLBO Gain Check Unchanged at 0.5, sample rate 100.
779 2021-08-26 12:40 KLBO Hardware Swap Trident 490 installed. TRD501 removed. Restarted. Still noisy.
780 2021-08-26 12:58 KLBO Seismo Swap CMG-40T T4317 removed, 40T T4567 installed.
781 2022-04-01 17:15 KLBO Seismo Swap CMG-40T removed, ESP T3H26 installed.
782 2022-04-01 17:20 KLBO VSAT Check X/Co pol 3/34 dB.
783 2022-04-01 17:30 KLBO Comments No change to gain of dc filter, station was good and secure.
784 2021-06-03 17:00 KLNB Comments VSAT errors must be due to antenna pointed while Eirp power was at -7dB. Power is restored to 0 dB at Hub and a re-point is needed.
No Date Time UT Location Activity Comments
785 2021-06-11 17:12 KLNB Service Visit To improve VSAT.
786 2021-06-11 17:14 KLNB VSAT Check Re-point. Co/X now at 26/0 dB. Was 22/0 dB before.
787 2021-06-11 17:15 KLNB Comments Earlier pointing was done when raining. No good. Today is cloudy but ok.
788 2021-06-11 17:16 KLNB Comments In Trident Operationl screen, RX overrun increased from 22 to 30 in about 10 mins. No good. Bus problem?
789 2021-06-11 17:43 KLNB Cold-Boot All is ok. Trident screen was checked. Overruns look ok.
790 2021-06-11 17:44 KLNB DC Filter Check DC filter is on at 0.001 HZ. Unchanged.
791 2021-06-11 17:45 KLNB Comments Construction work in underway near the site. Plan to finish in a few months.
792 2021-06-11 17:45 KLNB Gain Check Is 0.5, unchanged.
793 2021-06-11 17:46 KLNB Comments When construction is done, school buses will park near the site. Expect noise during their work hours.
794 2022-02-24 18:21 KLNB Service Visit Due to bad seismic data.
795 2022-02-24 18:22 KLNB Comments Flooded seismo, water level 5 cm above connector. Cleaned up, changed seismo cable, restarted.
796 2022-02-24 18:23 KLNB Seismo Check Mass-Centered ok.
797 2022-02-24 18:26 KLNB Seismo Check Calibrated OK. End of service visit.
798 2021-06-22 15:55 KLNB Closed End of Operation.
799 2021-08-04 16:30 KNDN New Station Code Start recording in KNDN ringbuffers.
800 2021-08-21 19:24 KNDN Firmware Upgrade Upgraded to Libra V5.20, TS V1.20, TR V1.50 EPLD V1.08
No Date Time UT Location Activity Comments
801 2021-09-14 21:10 KNDN SSPB Check Corrected SSPB temperature profile from flat 45 dB to manufacturer specs.
802 2021-09-21 22:26 KNDN Firmware Upgrade Upgraded to Libra Firmware version 5.31.
803 2021-07-20 19:19 KNDN DC Filter Check DC removal filter is off.
804 2021-07-20 19:19 KNDN Feed Horn Box New Feed Horn protective box installed. New calbes and conduits installed.
805 2021-07-20 19:19 KNDN Feed Window Replaced.
806 2021-07-20 19:19 KNDN Firmware Upgrade Libra 5.6201, TS 1.23, Trident 1.7201, 1.23.
807 2021-07-20 19:19 KNDN Gain Change Software gain changed to 1.0.
808 2021-07-20 19:19 KNDN Hardware Swap Cygnus 419 installed, 260 removed.
809 2021-07-20 19:19 KNDN Hub Swap Served by Ottawa Hub.
810 2021-07-20 19:19 KNDN Re-Point Co-pol 30, X-pol 0 dB.
811 2021-07-20 19:19 KNDN Restart Calibrated & mass-centred ok.
812 2021-07-20 19:19 KNDN Service Visit Site was down (dead Cygnus). Pinhole in feed window memberane.
813 2021-08-12 17:57 KNDN DC Filter Check Unchanged from Off.
814 2021-08-12 17:57 KNDN Gain Check Software gain unchanged from 1.
815 2021-11-04 15:24 KNDN Firmware Upgrade Trident 1.7210 in A and B. B is active.
816 2021-08-20 20:20 KNDN Service Visit To re-point antenna and borrow Cygnus.
No Date Time UT Location Activity Comments
817 2021-08-20 20:21 KNDN Re-Point Co/X improved from 30/5 to 30/1 dB.
818 2021-08-20 20:25 KNDN Hardware Swap Cygnus 419 removed, Cyg 557 installed.
819 2021-08-20 20:26 KNDN Comments Cyg 557 has bad com ports but they are not used here.
820 2021-08-20 20:27 KNDN Comments New Cyg added to Naqs2, Ottawa Hub.
821 2021-08-20 20:31 KNDN Comments New Cyg added to Naqs1, Ottawa Hub.
822 2021-08-20 20:42 KNDN Restart With new Cygnus 557.
823 2021-08-20 21:43 KNDN Comments New Cyg added to Naqs1, Ottawa Hub.
824 2021-08-20 21:45 KNDN Comments Lost some data and is not doing ReTX right now. Is it due to Cygnus 557 buffers not flushed?
825 2021-07-16 19:30 KNDN Service Visit To restore dead station.
826 2021-07-16 19:40 KNDN Wildlife Damage to power cable was fixed.
827 2021-07-16 19:50 KNDN Restart Station restarted
828 2022-04-27 KSVO New Station Start DC (Solar Panels). Installed in NY state.
829 2021-09-24 KSVO Closed Plan to move station.
830 2021-09-24 12:00 KSVO Firmware Upgrade Upgraded to Libra Firmware version 5.31.
831 2021-11-15 20:30 KSVO Hardware Swap Removed Cygnus 184, installed 171. Kept same Trident.
832 2021-11-15 22:00 KSVO Reopened Libra 5.31, Tsr1.21, Trd. 16202, epsd 123, gain 0.1, DC removal filter on at 0.001 Hz.
No Date Time UT Location Activity Comments
833 2021-11-05 19:00 KSVO Service Visit Station went down after days of acting up.
834 2021-11-05 19:01 KSVO VSAT Check Before repoint attemp, 32/2, after 32/2 dB. No chagne.
835 2021-11-05 19:02 KSVO Battery Check All batteries and charge controller ok.
836 2021-11-05 19:03 KSVO Cables LNB cable found loose. All cables tightened.
837 2021-11-05 19:05 KSVO Comments Was LNB cable the only reason for problems??
838 2021-08-18 17:56 LAIN New Station Start HB, Lailor River, recording on Naqs1 and Naqs2, Ottawa Hub.
839 2021-08-18 17:59 LAIN DC Filter Check DC filter is not on.
840 2021-08-18 17:59 LAIN Gain Check Gain of 1.0, sps 40.
841 2021-08-18 17:59 LAIN Hardware List Cyg 261, Trd 248, sspb A0011A, 2 W, seismo, ESP, T3b27.
842 2021-08-18 18:00 LAIN Comments Com port is not working on Cyg 261 but is not needed here.
843 2021-08-18 18:01 LAIN Comments Mass centred and calibrated ok.
844 2021-08-18 18:01 LAIN Require GPS Lock Require GPS Lock is on.
845 2021-08-18 18:01 LAIN VSAT Check On telesat, 30 dB of isolation.
846 2021-08-18 18:02 LAIN Comments Team are finishing off seismometer vault.
847 2021-08-18 18:02 LAIN Site Geology Bedrock site.
848 2021-08-31 11:57 LAIN Comments Unable to do scheduled service. Trident and /or seismometer are broken. No seismic data for months.
No Date Time UT Location Activity Comments
849 2021-08-31 12:02 LAIN TDMA Halt Removed from TDMA (Cygnus 261, carina 115) until next visit. Station down.
850 2021-09-13 LDGN New Station Start DC (Solar Panels).
851 2021-07-22 17:00 LDGN Hardware Swap Cygnus 158/HRD 209 replaced with Cygnus 245/Trident 237.
852 2021-07-22 17:00 LDGN Temp Stop Stopped recording in ringbuffers LDGN to swap electronics.
853 2021-07-23 01:00 LDGN SR Change Sampling rate from to 20 sps to 40 sps.
854 2021-07-23 16:00 LDGN Restart started recording in new LDGN ringbuffers.
855 2021-08-27 LDGN Firmware Upgrade Upgraded to Libra V5.20, TS V1.20, TR V1.50 EPLD V1.08
856 2021-09-21 15:56 LDGN Firmware Upgrade Upgraded to Libra Firmware version 5.31.
857 2021-07-09 19:30 LDGN Feed Horn Box New Feed Horn protective box installed. New conduits installed.
858 2021-07-09 19:30 LDGN Feed Window Replaced.
859 2021-07-09 19:30 LDGN Firmware Upgrade Libra 5.6201 installed and other firmware.
860 2021-07-09 19:30 LDGN Gain Check Software gain 1.0, DC removal filter on at 1 mHz, Input range 40 vPP.
861 2021-07-09 19:30 LDGN Hardware Swap Trident 423 installed, 237 removed.
862 2021-07-09 19:30 LDGN Hub Swap Served by Ottawa Hub.
863 2021-07-09 19:30 LDGN Re-Point SSPB was not replaced (no spare available). Cleaned-up old SSPB installed. Co-pol 17-20, X-pol 0 dB. IP 245. Pinged ok.
864 2021-07-09 19:30 LDGN Service Visit Site was down. Memberane on feed window was damaged. SSPB full of water.
No Date Time UT Location Activity Comments
865 2021-08-12 13:00 LDGN DC Filter Change Changed from On to Off.
866 2021-08-12 13:00 LDGN Gain Check Unchanged from 1.0.
867 2021-11-04 14:48 LDGN Firmware Upgrade Trident 1.7210 in A and B. B is active.
868 2021-06-11 19:42 LDGN VSAT Check Problem with VSAT: Dish out or Cygnus problem. TX is down by 5 dB but shape is ok. No data coming.
869 2021-06-11 19:44 LDGN Comments Site probably need a visit.
870 2022-05-03 01:22 LDGN Closed End of Operation.
871 2021-06-16 21:30 LDIO New Station Code Started recording on Naqs2 in LDIO ringbuffers.
872 2021-06-16 21:35 LDIO New Station Code Started recording on Ottawa Naqs1 in LDIO ringbuffers.
873 2021-06-16 21:38 LDIO Seismo Check Low frequency ride on Vertical comp.
874 2021-06-07 16:00 LDIO Service Visit To replace seismo and repoint.
875 2021-06-07 16:30 LDIO Re-Point X/co 0/28 dB, 2 dB improvement.
876 2021-06-07 16:45 LDIO Seismo Swap 3T seismometer removed due to LP noise, 3ESP s/n T3K42 installed.
877 2021-06-07 16:57 LDIO Seismo Check Mass centred and calibrated ok.
878 2021-07-25 01:00 LGSN New channel prototype Channel prototype to Broad band, Trident and CMG-40T seismometer (BHZ-2 to BHZ-3, BHN-2 to BHN-3, BHE-2 to BHE-3).
879 2021-08-13 23:08 LGSN Firmware Upgrade Upgraded to Libra V5.20, TS V1.20, TR V1.50 EPLD V1.08
880 2021-09-14 18:40 LGSN SSPB Check Corrected SSPB temperature profile from flat 45 dB to manufacturer specs.
No Date Time UT Location Activity Comments
881 2021-09-20 14:45 LGSN Firmware Upgrade Upgraded to Libra Firmware version 5.31.
882 2021-07-08 16:00 LGSN Feed Horn Box New Feed Horn protective box installed. New conduits installed.
883 2021-07-08 16:00 LGSN Feed Window Feed window replaced.
884 2021-07-08 16:00 LGSN Firmware Upgrade Libra 5.6201 installed.
885 2021-07-08 16:00 LGSN Re-Point Co-pol 25, X-pol 0 dB.
886 2021-07-08 16:00 LGSN Service Visit Site was working OK.
887 2021-08-13 14:59 LGSN Hardware Swap Switched from London to Ottawa Hub.
888 2021-08-13 16:00 LGSN Service Visit Animal damage. Feed horn bent, feed window pecked. Both fixed.
889 2021-08-13 16:01 LGSN VSAT Check Co/X 23/0 dB.
890 2021-08-13 16:02 LGSN Comments Need large wrench for next visit to tighten antenna.
891 2021-08-13 16:43 LGSN Restart Station restarted, Ottawa Hub, both Naqsservers.
892 2021-07-17 16:25 LGSN Closed End of Operation.
893 2021-09-08 17:00 LINO New Station Code Start recording in LINO ringbuffers.
894 2021-09-17 01:50 LINO DC Filter Change From off to on. At 0.001 mHz corner frequency.
895 2021-09-17 01:50 LINO Gain Change From 1.048576 to 0.1.
896 2022-05-04 14:20 LINO Closed End of Operation.
No Date Time UT Location Activity Comments
897 2022-05-07 15:59 LITE VSAT Check X-pol is off by 5 dB. Site visit is required.
898 2022-05-07 16:00 LITE Comments Eirp Inc. changed from -2 to -7 dB.
899 2021-06-10 17:26 LONDON Carina Hubs 3 out of 4 were hung due to power switch. Cold booted to fix problem.
900 2021-06-10 17:27 LONDON Carina Hubs Carina Hubs and fans will be moved to UPS tomorrow.
901 2021-10-31 15:10 LONDON Hub Swap All stations but TOBO moved from Carina 29 to Carina30 and 31.
902 2021-06-10 17:25 LONDON HUB Comments Moved NAQS1 and Polaris firewall, 2 network switches, modem, monitor and KVM switch to 24 hour UPS.
903 2021-05-25 15:00 LONDON HUB Re-Point Co/X improved by 2 dB to 30/0 dB.
904 2021-10-31 01:00 LONDON HUB Hub Swap All stations moved from Carina 28 to Carina30. Carina 28 powered down.
905 2022-04-10 17:00 LONDON HUB VSAT Check X-pol 30 dB, power improved by 2 dB.
906 2021-07-11 22:22 LUPN Hardware List CYG389 TRD437 CMG-3ESP T3E89, SPS 40, Gain 1., DC filter on at 1 mHz.
907 2021-07-11 22:22 LUPN New Station Code Start recording in LUPN ringbuffers.
908 2021-07-11 22:22 LUPN SSPB Check 2 Watts SSPB, S/N 16462
909 2021-07-11 22:22 LUPN Site Geology
910 2021-07-11 22:22 LUPN VSAT Start co-pol 25 x-pol 0
911 2021-07-12 19:17 LUPN DC Filter Check DC filter is turned off.
912 2021-07-12 21:31 LUPN Restart Initial installation of battery charger at the back of the antenna created spikes on CSA CANOPUS Riometer nearby. The battery charger was moved away from antenna to inside the CSA building and station restarted at 21:31 UT. Data prior to this time are disrupted frequently.
No Date Time UT Location Activity Comments
913 2021-08-12 18:06 LUPN DC Filter Check Unchanged from Off.
914 2021-08-12 18:06 LUPN Gain Check Software gain unchanged from 1.
915 2021-11-04 15:26 LUPN Firmware Upgrade Trident 1.7210 in A and B. A is active.
916 2021-10-01 19:18 MADG New Station Start At Ile d'Entree, near Magdelan Is., NS.
917 2021-10-02 23:37 MADG Recording Hub London.
918 2021-10-02 23:38 MADG DC Filter Check DC filter is off.
919 2021-10-02 23:38 MADG Gain Check Gain of 1.0 at 100 sps.
920 2021-10-02 23:38 MADG SR Change Incorrectly reported 100 sps, it remains at 40 sps until Oct. 21 at 12:48 UT.
921 2021-10-02 23:39 MADG Require GPS Lock On.
922 2021-10-02 23:39 MADG VSAT Check 2/30 dB.
923 2021-10-02 23:40 MADG Hardware List Cyg254, TRD767, sspb h0019a, ip 254.
924 2021-10-02 23:41 MADG SSPB Check Soil/clay.
925 2021-10-02 23:41 MADG Seismo Check Seismometer T3A81 is damaged, needs replacement.
926 2021-10-02 23:42 MADG Comments Antenna is mounted on King Post. Site is solar-powered.
927 2021-10-21 12:48 MADG SR Change From incorrect 40 sps to correct 100 sps.
928 2021-10-22 14:25 MADG Seismo Swap Seismo with bad Z component swapped with good seismo.
No Date Time UT Location Activity Comments
929 2021-10-22 14:25 MADG Seismo Swap T3A81 removed, T3B22 installed.
930 2021-09-06 14:43 MADG Carina Swap Switched from SES-AM to Telesat, Carina 92, London to Ottawa Hub.
931 2021-09-06 14:44 MADG Comments Corrosion problem, feed horn need to be replaced on next service.
932 2021-09-06 14:44 MADG VSAT Check 30 dB isolation, strong signl.
933 2021-09-06 14:45 MADG Comments Cell phone coverage from site, but poor line.
934 2021-10-15 14:30 MALG New Station Start Started station Malagash, NS recording at London Hub.
935 2021-10-19 19:01 MALG Hardware List Cygnus 398, Trd 418, seismo ESP, s/n T3B20.
936 2021-10-19 19:03 MALG DC Filter Check Off.
937 2021-10-19 19:03 MALG Gain Check Gain 1.0, SPS 100.
938 2021-10-19 19:03 MALG VSAT Check 29/0 dB.
939 2021-10-19 19:04 MALG Require GPS Lock On.
940 2021-08-23 16:53 MALG Comments Switched from SES-AM to Telesat.
941 2021-08-23 16:54 MALG Comments F-Connector repaired on LNB.
942 2021-08-23 16:54 MALG VSAT Check 34.5 dB isolation.
943 2021-06-20 19:18 MALO Hardware List Cyg392 Trd361 CMG-40T T4774, SPS 40, Gain 1.
944 2021-06-20 19:18 MALO New Station Code Start recording in MALO ringbuffers.
No Date Time UT Location Activity Comments
945 2021-06-20 19:18 MALO SR Change SR change to100 sps did not register. SR is still at 40.
946 2021-06-20 19:18 MALO Site Geology Bedrock. About 1.5 km from lake.
947 2021-06-20 19:18 MALO VSAT Start Xpol 1, Co-pol 31
948 2021-07-17 20:55 MALO Service Visit Station was OK but Hub communication to it was not possible due to unknown problem. Power was turned off.
949 2021-07-17 21:01 MALO Restart Restarted without any hardware change.
950 2021-07-17 21:15 MALO Hub Login Hub communication restored. Logged in an change sample rate from 40 to 100 sps on Cygnus 392.
951 2021-07-17 22:00 MALO Temp Stop Station stopped. End of Cygnus 392.
952 2021-07-17 22:22 MALO Hardware Swap Cygnus 422 installed, 392 removed.
953 2021-07-17 22:34 MALO Temp Stop Stopped to re-connect temperature sensor.
954 2021-07-17 22:39 MALO Restart Restarted.
955 2021-07-17 22:52 MALO Gain Check Software gain 1.0 , DC removal filter ON at 1 mHz, Input range 40 Vpp.
956 2021-07-17 22:52 MALO Sample Rate Change Sample rate is set to 100 SPS, on Cygnus 422. This is start of good data with new sample rate.
957 2021-08-11 19:42 MALO DC Filter Check Unchanged from ON at 0.001 Hz.
958 2021-08-11 19:42 MALO Gain Check Software gain unchanged from 1.0.
959 2021-10-29 20:30 MALO Service Visit Station not working ok due to bad Cygnus? Visit indicated old Cygnus not working in Test mode. No physical damage observed at structures.
960 2021-10-29 20:31 MALO Hardware Swap Cyg 422 removed, 220 installed.
No Date Time UT Location Activity Comments
961 2021-10-29 20:32 MALO VSAT Check 28/1. No change from installation time.
962 2021-10-29 20:34 MALO Restart All OK with new Cygnus.
963 2021-10-29 20:40 MALO Gain Check Unchanged from 1.
964 2021-10-29 20:41 MALO DC Filter Check Unchaged from On at 1 mHz.
965 2021-11-03 16:04 MALO Firmware Upgrade Trident 1.7210 in A and B. A is active.
966 2021-11-10 21:05 MALO Comments Rebooted Time Server to fix gps sat problem.
967 2021-08-17 14:00 MALO DC Filter Check Unchaged at on and 1 mHz.
968 2021-08-17 14:00 MALO VSAT Check Co/x unchanged at 29/0 dB.
969 2021-08-17 14:18 MALO Service Visit To upgrade seismometer.
970 2021-08-17 14:19 MALO Comments There is a new locked gate on the way to the site. However, battery box security cable was cut open but box was ok. Wire to one solar panel was cut, fixed.
971 2021-08-17 14:20 MALO Seismo Swap ESP T3J87 installed, 40T T4774 removed.
972 2021-08-17 14:52 MALO Cold-Boot To make sure all config. parameters are saved correctly.
973 2021-08-17 14:53 MALO Restart Station restarted. Both Ottawa Hub Naqsservers receive data.
974 2021-08-17 14:55 MALO Gain Check Unchaged at 1.0 (this is default for FedNor1).
975 2021-08-17 15:34 MALO Comments Some new mining/drilling activity near by (about 2 km).
976 2021-08-18 16:08 MALO Comments Security screws installed on battery boxes.
No Date Time UT Location Activity Comments
977 2021-07-26 20:00 MCKN New Station Code Start recording in MCKN ringbuffers.
978 2021-07-26 20:01 MCKN LVD Check The Low Volt Disconnect set to 11.0 VDC instead of 11.5.
979 2021-09-20 16:30 MCKN Firmware Upgrade Upgraded to Libra Firmware version 5.31.
980 2021-09-20 16:30 MCKN Repair Wildlife damage repaired. Station re-started.
981 2021-09-25 16:40 MCKN SSPB Check Corrected SSPB temperature profile from flat 45 dB to manufacturer specs.
982 2021-07-07 18:11 MCKN Feed Horn Box New Feed Horn protective box installed. New conduits installed.
983 2021-07-07 18:11 MCKN Firmware Upgrade Libra 5.6201 installed.
984 2021-07-07 18:11 MCKN Re-Point Co-pol 22, X-pol 0 dB. IFL loss at 3 db.
985 2021-07-07 18:11 MCKN Service Visit Cables were damaged but feed window and sspb ok.
986 2021-07-07 18:30 MCKN Hub Swap Served by Ottawa Hub.
987 2021-08-12 12:30 MCKN DC Filter Check Unchanged from Off.
988 2021-08-12 12:30 MCKN Gain Change Software gain changed from 1.048576 to 1.0.
989 2021-11-04 15:08 MCKN Firmware Upgrade Trident 1.7210 in A and B. B is active.
990 2021-08-14 14:30 MCKN Service Visit Trident not communicating to NMX bus, replaced.
991 2021-08-14 14:31 MCKN Comments Trd 137 had failed on July 30, 2004. Never recovered.
992 2021-08-14 14:31 MCKN Hardware Swap Trident 137 removed, 746 installed.
No Date Time UT Location Activity Comments
993 2021-08-14 14:32 MCKN Seismo Swap ESP T3J30 installed. 40T T4938 removed
994 2021-08-14 15:50 MCKN Restart Restarted recording, Ottawa.
995 2021-08-14 15:51 MCKN Cold-Boot Standard procedure.
996 2021-08-14 15:52 MCKN Gain Check Unchanged at 1.0 and sps of 40.
997 2021-08-14 15:53 MCKN DC Filter Check Unchanged at not on.
998 2021-08-14 18:17 MCKN Comments Started recording MT data on both Naqsservers at Ottawa.
999 2021-08-14 18:38 MCKN Comments MT Nims instrument configured at low gain with filter on DC.
1000 2021-08-14 18:44 MCKN Comments Old 486 computer used for MT setup has failed.
1001 2021-09-14 03:00 MCKN Comments Trident 746 is not communicating to Cygnus. No seismic data. This is second time that Cygnus 219 at this site is failing to service the NMX bus or causing Trident to die.
1002 2021-09-15 18:00 MCKN Restart Logged to MCKN over VSAT, Trident not showing. Reboot TimeServer in an attempt to power-cycle NMX bus.
1003 2021-09-15 18:07 MCKN Restart Reboot Cygnus in attempt to restart the Trident.
1004 2021-09-15 18:20 MCKN Comments Booting TimeServer and Cygnus did not fix the problem. Trident is not communicating. On TS Instrument screen, bytes sent are increasing, bytes/packets received continue to be zero.
1005 2021-07-20 18:00 MCKN Service Visit To restore seismic and service MT.
1006 2021-07-20 18:01 MCKN Hardware Swap Cygnus 550, Trdident 245 installed on Hub 6.
1007 2021-07-20 18:20 MCKN Comments Servicing MT equipment.
1008 2021-07-20 20:15 MCKN Gain Check Confirm gain is at 1.0.
No Date Time UT Location Activity Comments
1009 2021-07-20 20:16 MCKN Comments Finished MT, replace electords and checked Mag orientation and replace gps antenna mast. GPS was locked on MT.
1010 2021-07-20 20:16 MCKN Comments Require GPS Lock is on.
1011 2021-07-20 20:16 MCKN DC Filter Check Is not one.
1012 2021-07-20 20:17 MCKN Comments Team leaving site for Snap Lake. Due to leave for YK tomorrow.
1013 2022-05-07 21:00 MEDO New Station Code Started recording at MEDO ringbuffers.
1014 2022-05-07 21:01 MEDO Hardware List Cyg 394, TRD095, TS375, ESP T3B91, SSPB a0117a.
1015 2022-05-07 21:02 MEDO VSAT Check Co/X pol 36/6 dB.
1016 2022-05-07 21:03 MEDO Gain Check Started at Gain 0.5, sps 100.
1017 2022-05-07 21:05 MEDO DC Filter Check Started with DC filter on at 1 mHz.
1018 2022-05-07 21:07 MEDO Comments Fence installed to stop cows.
1019 2022-05-10 20:00 MEDO Firmware Upgrade Libra firmware to V6.65.01.
1020 2022-05-10 20:05 MEDO Firmware Upgrade TS upgraded to V2.00.00.
1021 2022-05-10 20:10 MEDO SSPB Check SSPB gain table set ftat to 45.0 dB.
1022 2022-05-10 20:15 MEDO Comments TX IFL loss set to 2.0 dB for 6 m cable.
1023 2022-05-10 20:20 MEDO Comments Ringbuffer sizes changed (BDunn).
1024 2022-01-13 18:00 MEDO Service Visit To upgrade and fix power system
No Date Time UT Location Activity Comments
1025 2022-01-13 18:10 MEDO Power Check Added 2 batteries. LVD reset from 12 VDC to 11.5 VDC.
1026 2022-01-13 18:20 MEDO Re-Point Before x/c was 32/2, after repoint it is 32/1 dB.
1027 2022-01-13 18:30 MEDO Comments Electric fence installed by land owner.
1028 2022-04-07 MGCB New Station Start DC (Solar Panels).
1029 2021-08-26 20:30 MGCB Firmware Upgrade Upgraded to Libra V5.20, TS V1.20, TR V1.50 EPLD V1.08
1030 2021-09-12 09:44 MGCB DC Filter Change From off to on. At 0.001 mHz corner frequency.
1031 2021-09-12 09:44 MGCB Gain Change From 1.048576 to 0.1.
1032 2021-09-24 09:24 MGCB Firmware Upgrade Upgraded to Libra Firmware version 5.31.
1033 2021-06-27 16:04 MGCB Gain Change Software gain at 1.0 (IA: check with IAK records)
1034 2021-06-27 16:04 MGCB Hardware Swap Cygn180 replaced with Cyg327. All othe hardware the same.
1035 2021-08-11 18:54 MGCB DC Filter Check Unchanged from ON at 0.001 Hz.
1036 2021-08-11 18:54 MGCB Gain Change Software gain changed from 0.1 to 0.5.
1037 2021-11-04 16:42 MGCB Firmware Upgrade Trident 1.7210 in A and B. A is active.
1038 2022-05-13 12:52 MGCB Closed End of Operation.
1039 2021-11-23 01:33 MGCB Hardware Swap Trident 080 installed, 083 removed.
1040 2021-08-04 01:25 MGTN New Station Code Start recording in MGTN ringbuffers.
No Date Time UT Location Activity Comments
1041 2021-08-04 16:20 MGTN Trident swap Changed Trident from 122 to 227.
1042 2021-08-27 21:42 MGTN Firmware Upgrade Upgraded to Libra V5.20, TS V1.20, TR V1.50 EPLD V1.08
1043 2021-09-23 21:?? MGTN Firmware Upgrade Upgraded to Libra Firmware version 5.31.
1044 2021-09-23 21:?? MGTN Software bug. Station had lost TDMA. Stopped.
1045 2021-07-20 15:31 MGTN Before Re-Point Before any re-adjustment, co-pol was low at 15 dB and no x-pol detected, indicating unstable antenna base.
1046 2021-07-20 15:31 MGTN Feed Horn Box New Feed Horn protective box installed. New cables and conduits installed.
1047 2021-07-20 15:31 MGTN Firmware Upgrade Libra 5.6201 installed?
1048 2021-07-20 15:31 MGTN Re-Point After re-pointing, co-pol was 21 with x-pol of 0 dB, improved as expected.
1049 2021-07-20 15:31 MGTN SSPB Check 2 Watts SSPB, S/N 16460 installed. Profile (25/46.9, -40/46.0, 50/47.5 entered).
1050 2021-07-20 15:31 MGTN Service Visit Site was working OK. Stopped for service. SSPB cable was loose, connector fell off. Minor damage to tarp on battery box.
1051 2021-07-20 16:39 MGTN Restart Calibrated & mass-centred ok.
1052 2022-02-23 19:03 MGTN Comments Station was not sending seismic signal. VSAT login showed no Trident. Rebooted.
1053 2022-02-23 19:05 MGTN Comments Reboot did not solve problem. Trident not showing. No seismic signal.
1054 2022-05-13 20:30 MGTN Temp Halt Station was removed from carina 31, TDMA. No data coming from Trident.
1055 2021-08-18 16:30 MGTN Service Visit Trident found dead. Swapped.
1056 2021-08-18 17:05 MGTN Hardware Swap Trident 694 installed, 178 removed.
No Date Time UT Location Activity Comments
1057 2021-08-18 17:08 MGTN Hub Swap Station moved from London to Ottawa Hub.
1058 2021-08-18 20:00 MGTN Hardware Swap Cygnus 261 removed, 557 installed.
1059 2021-08-18 20:01 MGTN VSAT Check Poor VSAT. Swapped Cygnus, no help. Did a re-point, improved Co/x from 13/0 to 26/0.
1060 2021-08-18 20:02 MGTN Comments Cyg 261 removed from MGTN might have been ok.
1061 2021-08-19 20:18 MGTN Comments Starting MT data at Ottawa Hub on Naqs2.
1062 2021-08-19 20:25 MGTN Comments Starting MT data at Ottawa Hub on Naqs1.
1063 2021-07-20 14:40 MGTN Service Visit To fix MT equipment.
1064 2021-07-20 14:41 MGTN Wildlife Sever damage, beyone repair, to MT equipment.
1065 2021-07-20 14:45 MGTN Comments Seismic is ok, MT can not be repair. Closed.
1066 2021-08-31 12:17 MGTN Comments Unable to do scheduled service. Station has been down since 2005.
1067 2021-08-31 12:18 MGTN TDMA Halt Removed from TDMA (Cygnus 559, Carina 62) until next visit. Station down.
1068 2022-03-22 23:48 MH1B Closed End of Operation.
1069 2022-03-20 22:16 MH2B Closed End of Operation.
1070 2021-08-13 23:14 MH3B New Station Start Started recording at Ottawa Hub Naqs1 and 2.
1071 2021-08-13 23:15 MH3B Hardware List Cyg 616, Trd 854, Trd 856, sspb a0401a, 40T T4776.
1072 2021-08-13 23:15 MH3B VSAT Check 40/3 dB.
No Date Time UT Location Activity Comments
1073 2021-08-13 23:16 MH3B Comments Seismometer is oriented to True North.
1074 2021-08-13 23:17 MH3B Gain Check Trd 854, Gain 0.5, sps 40.
1075 2021-08-13 23:18 MH3B DC Filter Check Unticked on Trd 854.
1076 2021-08-13 23:18 MH3B Gain Check Trd 856, Gain 0.5, sps 40.
1077 2021-08-13 23:19 MH3B DC Filter Check Trd 856, unticked.
1078 2021-08-13 23:20 MH3B Require GPS Lock Require GPS lock to lock is on on both Tridents.
1079 2021-08-13 23:20 MH3B Site Geology Riverbed silt, seismo on concrete slab.
1080 2021-08-13 23:21 MH3B Comments Golf course site.
1081 2021-08-13 23:21 MH3B Comments Seismo is at 46.30846 N 119.33798 W 117 m
1082 2021-08-14 00:12 MH3B Comments Wind sensor data are received at the Hub after a power cycle of the WS425 Wind Sensor. Team about to leave the site.
1083 2021-08-11 18:28 MH4B New Station Start Started at MSH site close to volcano
1084 2021-08-11 18:29 MH4B Hardware List Cyg 619, Trd 064, TRD 137, sspb a0406a, seismometer 40T, T4777
1085 2021-08-11 18:31 MH4B Comments Seismometer is oriented to North.
1086 2021-08-11 18:32 MH4B Gain Check TRD 064, SPS 40, Gain 0.5
1087 2021-08-11 18:32 MH4B VSAT Check 36/5 dB.
1088 2021-08-11 18:36 MH4B DC Filter Check TRd 064, DC filter in not on.
No Date Time UT Location Activity Comments
1089 2021-08-11 18:37 MH4B DC Filter Check Trd 137, DC is not on.
1090 2021-08-11 18:37 MH4B Gain Check Trd 137, sps 40, Gain 0.5
1091 2021-08-11 18:38 MH4B Require GPS Lock Require GPS Lock is on on both Tridents.
1092 2021-08-11 18:38 MH4B Site Geology Volcanic ash. Seismo on cement block.
1093 2021-08-11 18:40 MH4B Comments Team is taking GPS locations of all sensors.
1094 2021-08-11 19:20 MH4B Comments GPS location of seismo 46.30791 N 122.26601 W 1051 m
1095 2021-08-11 19:21 MH4B New Station Code Started at both Naqs1 and Naqs2, Ottawa Hub at18:27.
1096 2021-08-03 20:30 MLON New Station Code Start recording in MLON ringbuffers.
1097 2021-09-03 18:01 MLON Firmware Upgrade Upgraded to Libra V5.20, TS V1.20, TR V1.50 EPLD V1.08
1098 2021-09-14 22:30 MLON SSPB Check Corrected SSPB temperature profile from flat 45 dB to manufacturer specs.
1099 2021-09-20 21:10 MLON Firmware Upgrade Upgraded to Libra Firmware version 5.31.
1100 2021-09-20 21:10 MLON Site Visit All appear to be ok at this station.
1101 2021-07-08 18:52 MLON Feed Horn Box New Feed Horn protective box installed. New conduits installed.
1102 2021-07-08 18:52 MLON Firmware Upgrade Libra 5.6201 installed.
1103 2021-07-08 18:52 MLON Re-Point Co-pol 22 X-pol 2 dB.
1104 2021-07-08 18:52 MLON Service Visit Site was working OK.
No Date Time UT Location Activity Comments
1105 2021-08-21 23:55 MPPO New Station Code Start recording in MPPO ringbuffers.
1106 2021-09-16 21:20 MPPO DC Filter Change From off to on. At 0.001 mHz corner frequency.
1107 2021-09-16 21:20 MPPO Gain Change From 1.048576 to 0.1.
1108 2022-05-06 17:39 MPPO Site Visit Maintenance visit. Site in perfect condition.
1109 2022-05-06 17:39 MPPO VSAT Check Co-pol 32, X-pol 9dB. Adjusted, co-pol 33, X-pol 3 dB.
1110 2021-08-02 21:30 MPPO Service Visit To fix VSAT problem. GPS engine suspected defective. Swapping Cygnus.
1111 2021-08-02 21:35 MPPO Hardware Swap Cyg 268 removed, Cyg 563 installed.
1112 2021-08-02 21:35 MPPO Hub Swap Moved from London to Ottawa Hub.
1113 2021-08-02 22:06 MPPO Restart Station started at Ottawa Hub, Naqs2.
1114 2021-08-02 22:08 MPPO Cold-Boot Standard procedure.
1115 2021-08-02 22:15 MPPO Comments Site in perfect condition. Old Cygnus GPS tracking but not locking.
1116 2021-08-02 22:15 MPPO DC Filter Check Unchaged at on at 0.001 Hz.
1117 2021-08-02 22:15 MPPO Gain Check Unchanged at 0.5
1118 2021-08-03 00:58 MPPO New Station Code Started recording in Naqs1 ringbuffers.
1119 2022-04-02 15:04 MPPO Service Visit Charge controller was broken, replaced.
1120 2022-04-02 15:06 MPPO VSAT Check 30/0 dB.
No Date Time UT Location Activity Comments
1121 2022-04-02 15:10 MPPO Comments Mass-center and calibration ok.
1122 2021-08-07 19:45 MRHQ New Station Code Started recording at Ottawa Naqs2 in MRHQ ringbuffers.
1123 2021-08-07 19:50 MRHQ New Station Code Started recording at Ottawa Naqs1 in MRHQ ringbuffers.
1124 2021-05-24 15:00 MRHQ Service Visit To replace bad seismometer.
1125 2021-05-24 15:10 MRHQ Seismo Swap Seismometer was flooded. Replaced with ESP T3K95.
1126 2021-05-24 15:15 MRHQ Power Check Braker was incorrectly installed. Corrected.
1127 2021-05-24 15:20 MRHQ VSAT Check Antenna gain was set incorrectly to 45.5, was corrected to 46.5 dB.
1128 2021-05-24 15:25 MRHQ VSAT Check SSPB profile corrected.
1129 2021-05-24 15:30 MRHQ Re-Point Before 32/8, now 35/0 dB Co/X-pol.
1130 2022-04-04 18:30 MRHQ Service Visit Needed repoint.
1131 2022-04-04 18:35 MRHQ VSAT Check 26/0 dB.
1132 2022-04-04 18:36 MRHQ SSPB Check New 2 W SSPB.
1133 2022-04-04 18:37 MRHQ Power Check LVD added to AC power system.
1134 2022-04-04 18:40 MRHQ Comments Snowing heavily at the site.
1135 2021-07-23 15:17 MSNO New Station Code Started at Ottawa Hub Naqs2, Moosonee.
1136 2021-07-23 15:23 MSNO New Station Start Started at Ottawa Hub Naqs1, Moosonee.
No Date Time UT Location Activity Comments
1137 2021-07-23 15:26 MSNO Gain Check SPS 100, Gain 1.
1138 2021-07-23 15:27 MSNO Comments Require GPS lock to lock is on.
1139 2021-07-23 15:27 MSNO DC Filter Check DC filer on at 1 mHz.
1140 2021-07-23 15:28 MSNO Comments Finished installation, team to install Timmins Sunday.
1141 2021-06-15 00:01 MUMO Hardware List Cyg390 Trd364 3ESP T3F10 check w CA. SPS 100, Gain 1.
1142 2021-06-15 00:01 MUMO New Station Code Start recording in MUMO ringbuffers.
1143 2021-06-15 00:01 MUMO Site Geology Solid bedrock, about 100 m off the end of the runway.
1144 2021-06-15 00:01 MUMO VSAT Start X-pol 0 Co-pol 35 dB
1145 2021-08-11 19:35 MUMO DC Filter Check Unchanged from ON at 0.001 Hz.
1146 2021-08-11 19:35 MUMO Gain Check Software gain unchanged from 1.0.
1147 2021-11-03 16:05 MUMO Firmware Upgrade Trident 1.7210 in A and B. B is active.
1148 2021-08-04 19:44 NANO New Station Code Started recording with code NANO at Aroland/Nakina in Ottawa Hub Naqs2.
1149 2021-08-04 19:48 NANO New Station Code Started recording with code NANO at Aroland/Nakina in Ottawa Hub Naqs1.
1150 2021-08-04 19:50 NANO Comments End of FedNor3 installation.
1151 2021-10-05 14:00 NANO Service Visit Vandalism, station turned off. Electronics replaced.
1152 2021-10-05 14:10 NANO VSAT Check 30 dB xpol.
No Date Time UT Location Activity Comments
1153 2021-07-25 19:29 NAQS1 Comments Dataserver restarted on Ottawa Hub Naqs1.
1154 2021-11-26 15:00 NATG Comments Recording at Ottawa Hub, Telesat.
1155 2021-11-26 15:00 NATG New Station Start Last of the Atlantic 2005, about 40 km from Natashquan, Quebec.
1156 2021-11-28 14:32 NATG DC Filter Check Off.
1157 2021-11-28 14:32 NATG Gain Check 1.0 at 100 sps.
1158 2021-11-28 14:32 NATG Require GPS Lock On.
1159 2021-11-28 14:33 NATG Hardware List Cyg 161, Trd 913, TS 075, esp T3P19 with auto mass-lock.
1160 2021-11-28 14:33 NATG Site Geology Solid Bedrock.
1161 2021-11-28 14:33 NATG VSAT Check 34 dB isolation, Telesat.
1162 2021-11-28 14:34 NATG Comments King Post bolted to rock face. Need more bolts on next visit.
1163 2021-11-28 14:34 NATG Power Check AC site. No grounding in supplied power.
1164 2021-11-28 14:35 NATG Comments Cygnus and charging system in battery box with two AC-type batteries.
1165 2021-09-10 16:44 NATG Service Visit To fix bad VSAT.
1166 2021-09-10 16:45 NATG Hardware Swap LNB replaced before bad Cygnus was spotted.
1167 2021-09-10 16:45 NATG Hardware Swap TimeServer problem with Cygnus 161. Replaced with Cygnus 605.
1168 2021-09-10 16:46 NATG Seismo Check Calibrated OK.
No Date Time UT Location Activity Comments
1169 2021-09-10 16:46 NATG VSAT Check 32.5 dB, 2 dB improvement.
1170 2021-09-10 16:47 NATG Comments VSAT bolts to rock were ok but secured.
1171 2021-07-29 20:30 NODN New Station Code Start recording in NODN ringbuffers.
1172 2021-08-28 18:53 NODN Firmware Upgrade Upgraded to Libra V5.20, TS V1.20, TR V1.50 EPLD V1.08
1173 2021-09-14 23:00 NODN SSPB Check Corrected SSPB temperature profile from flat 45 dB to manufacturer specs.
1174 2021-09-23 19:44 NODN Firmware Upgrade Upgraded to Libra Firmware version 5.31.
1175 2021-07-07 21:34 NODN Feed Horn Box New Feed Horn protective box installed. New conduits installed.
1176 2021-07-07 21:34 NODN Feed Window Feed window replaced.
1177 2021-07-07 21:34 NODN Firmware Upgrade All new firware (due to out-of-sync LibGui and firmware, initially had trouble sending test signal)
1178 2021-07-07 21:34 NODN Hardware Swap Cygnus 395 installed, 265 removed.
1179 2021-07-07 21:34 NODN Re-Point Co-pol 38, X-pol 2 dB. Strong Ontario-like signal.
1180 2021-07-07 21:34 NODN SSPB Check New 2 watts SSPB SN 16464 installed. Profile (25/46.93, -40/45.73,50/47.73).
1181 2021-07-07 21:34 NODN Service Visit Damage to feed horn mount. Looked like bear visit.
1182 2021-07-07 22:49 NODN Hub Swap Served by Ottawa Hub.
1183 2021-08-12 12:40 NODN DC Filter Check Unchanged from Off.
1184 2021-08-12 12:40 NODN Gain Change Software gain changed from 1.048576 to 1.0.
No Date Time UT Location Activity Comments
1185 2021-11-04 16:24 NODN Firmware Upgrade Trident 1.7210 in A and B. A is active.
1186 2021-08-20 17:30 NODN Service Visit Site was down due to a dead Cygnus. No other damage to anything.
1187 2021-08-20 17:31 NODN Hardware Swap Dead Cygnus 395 removed, Cyg261 installed.
1188 2021-08-20 17:32 NODN Comments Com ports not working for Cyg261 but they are not in use.
1189 2021-08-20 17:33 NODN Re-Point Co/X was 18/0 before re-point and is now 23/0 dB.
1190 2021-08-20 17:34 NODN Restart Recording at Ottawa Hub, Naqs2.
1191 2021-08-20 18:27 NODN Restart Ottawa Hub Naqs1 restarted.
1192 2021-08-20 18:37 NODN Comments No change to gain or dc filter.
1193 2021-07-17 15:51 NODN Closed End of Operation.
1194 2021-07-29 13:33 NSKO New Station Code Started at Ottawa Hub Naqs 2.
1195 2021-07-29 13:36 NSKO New Station Code Started at Ottawa Hub Naqs 1.
1196 2021-09-17 13:16 NWT VSAT Check Bandwidth increased for YNEN and GBLN, decreased for MCKN.
1197 2021-07-20 19:59 O21T New Station Start Started new station at Gameti Lake.
1198 2021-07-20 20:02 O21T Hardware List Cyg 558, TS 759, TRD 757, SSPB A0407A, seism. ESP, T3J18.
1199 2021-07-20 20:04 O21T DC Filter Check Started with DC Filter off, NWT default.
1200 2021-07-20 20:04 O21T Gain Check Started with sps 40 and gain 1.0 default nwt.
No Date Time UT Location Activity Comments
1201 2021-07-20 20:04 O21T VSAT Check Co/x 25/2 dB.
1202 2021-07-20 20:05 O21T Site Geology Near airport, sand, rock, esker.
1203 2021-07-20 20:06 O21T Comments Airport nearby is used a few times/day.
1204 2021-07-20 20:06 O21T Comments One of the 20 amp breaker on solar input is faulty.
1205 2021-07-20 20:07 O21T SSPB Check SSPB gain table set according to sspb profile.
1206 2021-08-26 20:56 O22T VSAT Start Started recording test station in Ottawa on E2R satellite.
1207 2021-09-06 22:08 O22T New Station Start Started new AC station at Rankin Inlet Airport, YRTN, served by Ottawa Hub, Antenna 2, Telesat Anik E2R.
1208 2021-09-06 22:10 O22T VSAT Check 32 dB of isolation. Good.
1209 2021-08-05 17:45 O23T New Station Start Started new station at Restoule Prov. Park.
1210 2021-08-05 17:46 O23T Gain Check Cyg. 553, Trd. 750, TS 667, sspb A0399A, seis. esp. T3K46.
1211 2021-08-05 17:47 O23T Gain Check Started with gain of 0.5 sps 100.
1212 2021-08-05 17:47 O23T VSAT Check Co/x: 34/4 dB.
1213 2021-08-05 17:48 O23T DC Filter Check Started with On at 0.001 Hz.
1214 2021-08-05 17:49 O23T Cold-Boot Did as standard procedure.
1215 2021-08-05 17:49 O23T Comments Cleared ringbuffers.
1216 2021-08-05 17:49 O23T Site Geology Solid bedrock.
No Date Time UT Location Activity Comments
1217 2021-08-05 17:50 O23T Comments Started at Ottawa Hub.
1218 2021-08-05 17:50 O23T Power Check 8 solar panels and 10 batteries installed.
1219 2021-08-07 19:21 O25T New Station Start Started new station at Morin Heights, MRHQ.
1220 2021-08-07 19:22 O25T Hardware List Cyg 556, ts 615, trd 772, sspb a0391a, esp. T3K13.
1221 2021-08-07 19:23 O25T Site Geology Solid bedrock.
1222 2021-08-07 19:24 O25T Cold-Boot Standard procedure.
1223 2021-08-07 19:25 O25T VSAT Check Co/x 35/5 dB.
1224 2021-08-07 19:26 O25T Gain Check Started with gain of 0.5 at 100 sps.
1225 2021-08-07 19:27 O25T DC Filter Check Started with dc filter on at 0.001 hz.
1226 2021-08-07 19:28 O25T Comments This is an AC site, served by Ottawa Hub.
1227 2021-08-07 19:29 O25T Comments Short trees planted to onscure the VSAT dish.
1228 2021-07-18 01:40 O28T New Station Start At location about 10 miles N of Castor Lake
1229 2021-07-18 01:44 O28T Hardware List Cyg 560, TRD 752, TS757, SSPB A093A,Seis. esp T3J33
1230 2021-07-18 01:45 O28T DC Filter Check Off at 0.001 startup.
1231 2021-07-18 01:45 O28T Gain Check Gain 1.0 at 40 sps startup.
1232 2021-07-18 01:45 O28T VSAT Check Co/x 33/2 dB.
No Date Time UT Location Activity Comments
1233 2021-07-18 01:46 O28T Comments No good location near original Castor lake, so moved to this location.
1234 2021-07-18 01:46 O28T Site Geology Sand, few trees around. A few trees cut to produce view.
1235 2021-07-18 01:47 O28T Seismo Check Mass centered ok.
1236 2021-07-18 01:49 O28T Cold-Boot No problem.
1237 2021-07-18 01:52 O28T Comments On Sunday, plan to visit ILKN and GBLN and DSMN, then to Gameti lake to install new site.
1238 2021-07-31 18:54 O30T Service Visit Visited HBDB to swap bad cygnus. Recording at O30T ringbuffer for now.
1239 2021-07-31 18:55 O30T VSAT Check Co/X 33/2 dB. Minor adjustment was done.
1240 2021-09-04 21:17 O42T New Station Start Started new DC station at Qililugaq Exploration Camp, QILN, served by Ottawa Hub Antenna 2, Telesat.
1241 2021-09-04 21:18 O42T VSAT Check Co-pol isolation 30-32 dB. Good.
1242 2021-09-04 21:19 O42T Comments Minor damage to antenna edge and back during transportation. Front ok.
1243 2021-09-02 18:30 O43T New Station Start Starting new station at Igloolik Airport. Installation in progress.
1244 2021-09-02 18:45 O43T Comments Too cold for cement to set on seismo vault. Waiting.
1245 2021-09-02 19:00 O43T Power Check Powered by AC from nearby airport shack.
1246 2021-09-02 19:10 O43T Repair 2 W SSPB failed. Replaced with one Watt unit.
1247 2021-07-29 11:53 O47T New Station Start Started at Neskantaga First Nation Airport.
1248 2021-07-29 12:44 O47T Hardware List Cyg 268, trd 396, sspb a0015a, esp seismo t3p34
No Date Time UT Location Activity Comments
1249 2021-07-29 13:16 O47T Gain Check Sps 100, gain 1.0
1250 2021-07-29 13:16 O47T VSAT Check 22/2 dB.
1251 2021-07-29 13:17 O47T Comments Require GPS Lock is on.
1252 2021-07-29 13:17 O47T Gain Check On at 0.001 Hz.
1253 2021-07-29 13:17 O47T Site Geology Soft sandy soil.
1254 2021-07-29 13:18 O47T Comments Location 52.19650 N 87.93053 W 241 m
1255 2021-07-29 13:22 O47T Comments Site is 100 m from runway at small airport with about 5 daily flights.
1256 2021-06-26 19:51 O50T New Station Start Recording at YOSQ in O50T ringbuffers.
1257 2021-07-31 15:05 O50T Hardware List Cyg 665, Trd 747.
1258 2021-07-31 15:05 O50T New Station Code New test station started for MACE replacement.
1259 2021-06-23 16:00 O51T New Station Start At Wemindji. Recording in O51T ringbuffers at Ottawa.
1260 2021-07-23 14:55 O53T New Station Code Starting new station at Ottawa Hub, Moosonee location.
1261 2021-07-23 14:56 O53T Hardware List Cyg 606, trd 851, seismo 3esp t3p31, sspb a0016a
1262 2021-07-23 14:58 O53T Site Geology Soft sand, about 100 m from Airport runway.
1263 2021-07-23 14:58 O53T VSAT Check 30/0 dB.
1264 2021-08-02 18:28 O54T New Station Start Started at kasabonika Lake, KASO recording at Ottawa Hub.
No Date Time UT Location Activity Comments
1265 2021-08-02 18:30 O54T Hardware List Cyg 607, Trd 852, TS 748, sspb a0019a, esp seismo T3P61, IP 78
1266 2021-08-02 18:40 O54T VSAT Check 35/5 dB.
1267 2021-08-02 18:50 O54T Comments Require GPS Lock is on.
1268 2021-08-02 18:50 O54T DC Filter Check On at 0.001 Hz.
1269 2021-08-02 18:50 O54T Gain Check SPS 100, Gain 1.0, Vpp 40.
1270 2021-08-02 18:55 O54T Comments About 80 m from runway.
1271 2021-08-02 18:55 O54T Comments Site is at small airport with 5-7 flights a day. Expect high wind and snow.
1272 2021-08-02 18:55 O54T Site Geology Seismometer in sand, at 53.52785N 88.64139W 192 m.
1273 2021-07-25 18:55 O62T New Station Start Started Timmins stations in O62T at Ottawa Hub.
1274 2021-07-25 18:56 O62T Hardware List Cyg 610, Trd 227, ESP T3F12, SSPB A0013A
1275 2021-07-25 19:05 O62T Seismo Check Bedrock at 48.46587 N 81.30321 W, 392 m
1276 2021-07-25 19:06 O62T Gain Check 100 SPS at 1.0.
1277 2021-07-25 19:06 O62T VSAT Check 30/0 dB. IP 79.
1278 2021-07-25 19:07 O62T Comments Require GPS Lock is on.
1279 2021-07-25 19:07 O62T DC Filter Check On at 0.001 Hz.
1280 2021-07-25 19:08 O62T Seismo Check Mass centred and calibrated ok.
No Date Time UT Location Activity Comments
1281 2021-08-04 19:13 O62T New Station Code New station started at Aroland Nakina, recording at Ottawa Hub.
1282 2021-08-04 19:27 O62T Hardware List Cyg 395, TRD 853, TS 480, sspb A0403A, ESP seismo T3K12.
1283 2021-08-04 19:28 O62T Gain Check SPS 100, Gain 1.0, Vpp 40.
1284 2021-08-04 19:28 O62T VSAT Check 31/0 dB.
1285 2021-08-04 19:29 O62T Comments Require GPS Lock is on.
1286 2021-08-04 19:29 O62T DC Filter Check On at 0.001 Hz.
1287 2021-08-04 19:29 O62T Site Geology Bedrock at 50.35429 N, 86.96841 W, 309 m
1288 2021-08-04 19:30 O62T Comments Site is about 1/2 hours from the Band.
1289 2021-07-20 17:50 O64T New Station Start New station started near Cobalt.
1290 2021-07-20 17:51 O64T Hardware List Cyg 561, Trident 754, sspb A0014a, esp T3P33.
1291 2021-07-20 17:52 O64T Comments At 47.37082 N, 79.66566 W, 306 m.
1292 2021-07-20 17:52 O64T VSAT Check Co/x 32/2 dB.
1293 2021-07-20 17:54 O64T Gain Check 1.0 at 100 sps.
1294 2021-07-20 17:54 O64T Site Geology Solid bedrock.
1295 2021-07-20 17:55 O64T Comments Require GPS Lock is on.
1296 2021-07-20 17:55 O64T DC Filter Check On at 0.001 Hz.
No Date Time UT Location Activity Comments
1297 2021-07-20 17:56 O64T Cold-Boot To restart and capture screens.
1298 2021-12-04 00:00 O66T Comments Started long-term test of Trillium 120P in Ottawa insulated vault.
1299 2021-12-06 14:38 O66T Comments Trident moved into the insulated cavity near seismometer to register seismometer temperature more accurately.
1300 2021-12-06 14:38 O66T Hardware List Trillium s/n is 121.
1301 2022-04-05 17:25 O66T VSAT Check Needed repoint. Now 29/0 dB, improved by 10 dB.
1302 2022-04-05 17:30 O66T Comments Testing Trillium 120P here, antenna movement consistent with 120P showing effect of tilt.
1303 2022-04-12 18:23 O66T Comments Trident digitizer configuration corrected from +12V, active low, High, High to +5V, active High, Low, Low, Low to match Trillium seismometer.
1304 2022-04-12 18:25 O66T Comments Naqs Staion file at Ottawa corrected for Trident connected to Trillium.
1305 2021-07-09 12:50 O66T VSAT Check 30 dB. Station is up.
1306 2021-07-09 12:51 O66T Comments Mag connected to this station.
1307 2021-06-21 19:56 O70T New Station Start Start recording in O70T ringbuffers at KILO.
1308 2021-06-14 19:10 O71T New Station Start Start recording in O71T ringbuffers at MUMO.
1309 2021-06-10 21:45 O72T New Station Start Start recording in O72T ringbuffers at VIMO.
1310 2021-06-20 15:04 O73T New Station Start Start recording in O73T ringbuffers at MALO.
1311 2021-06-23 20:35 O74T New Station Start Start recording in O74T ringbuffers at SUNO.
1312 2021-06-18 19:58 O77T New Station Start Start recording in O77T ringbuffers at OTRO.
No Date Time UT Location Activity Comments
1313 2021-07-11 22:00 O79T New Station Start Start recording in O79T ring buffers at LUPN. Station co-installed with Canadian Space Agency (CSA) CANOPUS station at Contort Lake, Lupin where Magnetometer and Riometer are operating.
1314 2021-06-09 00:21 O81T New Station Start Start recording in O81T ringbuffers at SILO.
1315 2021-09-24 20:40 O82T Hardware List Cyg 260, Trd 143, CMG-40T T4769, SSPB H0028A.
1316 2021-09-24 20:40 O82T New Station Start Start recording in O82T ringbuffers at GBLN.
1317 2021-09-24 20:40 O82T Site Geology
1318 2021-09-24 20:40 O82T VSAT Start Co-pol 20 X-pol 0
1319 2021-09-27 18:10 O84T Hardware List Cyg 251, Trd 501, CMG-40T T4316, SSPB A0147A.
1320 2021-09-27 18:10 O84T New Station Start Start recording in O84T ringbuffers at ILKN.
1321 2021-09-27 18:10 O84T Site Geology
1322 2021-09-27 18:10 O84T VSAT Start Co-pol 27 X-pol 0
1323 2022-02-27 15:11 O88T New Station Start O88T, Ottawa Vault is now under Test.
1324 2022-02-02 17:41 O91T Comments Temp Station O91T, at least, on Jan. 31 and Feb. 1, 04 shows Trident SOH temperature is about 10 deg warmer than the SSPB temperature. Both sensors are in the same test vault.
1325 2021-06-11 18:12 O93T Hardware List Sspb a0340a, cyg 523, ts 653, trd 693. seis t379 esp.
1326 2021-06-11 18:12 O93T New Station Start Started new site at Exp. Lake.
1327 2021-06-11 18:13 O93T VSAT Check Co/x 25/0 dB.
1328 2021-06-11 18:14 O93T DC Filter Check On is at 0.001 starting.
No Date Time UT Location Activity Comments
1329 2021-06-11 18:14 O93T Gain Check Sps 100, gain is 0.5 starting.
1330 2021-06-11 18:14 O93T Site Geology Cleared forested area, exposed bedrock.
1331 2021-06-11 18:15 O93T Comments Trident input range is at 40 pp.
1332 2021-06-11 18:16 O93T Comments Finishing off the site in the next hour.
1333 2021-06-11 18:17 O93T Comments Cold boot was done before latest checks.
1334 2021-06-13 16:51 O94T New Station Start Installed and starting at Red Lake but not connected to Hub yet. VSAT problems.
1335 2021-06-13 17:21 O94T New Station Start Cygnus 536 is installed, 535(not working) removed. VSAT is on now.
1336 2021-06-13 17:35 O94T Comments Mass centered ok.
1337 2021-06-13 17:37 O94T Site Geology Bedrock, spotty. Seismo on bedrock. VSAT is bolted. Solar panels not.
1338 2021-06-13 17:38 O94T Hardware List Cyg 536, ts 706, trd 695, seism esp T3762, 3T seismo. sspb a0389a.
1339 2021-06-13 17:38 O94T VSAT Check Co/X 32/0 dB.
1340 2021-06-13 17:40 O94T DC Filter Check On at 0.001Hz startup.
1341 2021-06-13 17:40 O94T Gain Check 100 sps, gain 0.5, startup gain.
1342 2021-06-13 17:41 O94T Comments On a decommissioned gold mine. About a KM from some workshop with minor activity.
1343 2021-06-18 20:34 O95T New Station Start New station at pnpo Pukaskwa park.
1344 2021-06-18 20:35 O95T Comments Test name is not matched correctly with Cyg. Trident.
No Date Time UT Location Activity Comments
1345 2021-06-18 20:36 O95T Hardware List Cyg 178, TS 043, Trd. 730, sspb a0396a, seismo. 3T, T3189.
1346 2021-06-18 20:36 O95T VSAT Check Co/x 35/5 dB.
1347 2021-06-18 20:37 O95T DC Filter Check On at 0.001 Hz.
1348 2021-06-18 20:37 O95T Gain Check Gain is 0.5 startup. sps is at 100.
1349 2021-06-18 20:38 O95T Cold-Boot Issued. OK.
1350 2021-06-18 20:38 O95T Comments One of the 12 solar panels broken but usable, it is in use.
1351 2021-06-18 20:38 O95T Site Geology Solid bedrock.
1352 2021-06-09 21:30 O96T New Station Start Started first of FedNor2 stations at Atikokan Iron Mine, ATKO, recoding in O96T ringbuffers for the next 23 minutes.
1353 2021-06-16 21:03 O97T New Station Start Started at LDIO in O97T ringbuffers.
1354 2021-06-16 21:04 O97T Hardware List Cyg540, ts697, trd 736, sspb a0394a, seism. 3T, T3190.
1355 2021-06-16 21:05 O97T Gain Check Gain of 0.5 at sps of 100 at start.
1356 2021-06-16 21:05 O97T VSAT Check Co/x: 35/4 dB.
1357 2021-06-16 21:06 O97T DC Filter Check On at 0,001 Hz.
1358 2021-06-16 21:06 O97T Site Geology Exposed bedrock (cleared for exploration, no longer in use).
1359 2021-06-16 21:07 O97T Comments Hilty drill broke on last hole (overheated). Truck window shattered. Trailer tire burst.
1360 2021-06-16 21:14 O97T Seismo Check Seismo calibration ok.
No Date Time UT Location Activity Comments
1361 2021-06-15 14:54 O99T New Station Start Starting in O99T buffers at Pickle Lake, PKLO
1362 2021-06-15 14:55 O99T Hardware List SSPB A0393A, CYG542, TS693, TRD738, Seism. 3T, T3249
1363 2021-06-15 14:56 O99T DC Filter Check On at 0.001 Hz.
1364 2021-06-15 14:56 O99T Gain Check SPS 100, gain 0.5 starting.
1365 2021-06-15 14:57 O99T Cold-Boot Done.
1366 2021-06-15 14:57 O99T Site Geology Exposed bedrock near old quarry (being used for gravel, no crushing at this time).
1367 2021-06-15 14:57 O99T VSAT Check Co/x 32/2 dB.
1368 2021-06-15 14:58 O99T Comments Gate, but open mostly. MNR people to check on the site. Appear not to be a security issue.
1369 2021-06-15 14:59 O99T Comments Mass Centred ok.
1370 2021-06-15 15:00 O99T Comments Driving to next site today. Be there tonight.
1371 2021-06-19 17:55 ORHO Internet Station New internet station at Orleans (Heritage park) Ontario, Ottawa Naqs1.
1372 2021-06-19 17:57 ORHO Internet Station 100 sps, 40Vpp, Software Gain 1, No DC filter, Low Impedance.
1373 2021-06-19 18:01 ORHO Hardware List Taurus 267, Guralp T3, T3762.
1374 2021-06-19 18:01 ORHO Require GPS to Lock On.
1375 2021-06-19 18:54 ORHO Comments Station off-line until June 22.
1376 2021-06-23 03:00 ORHO Restart Station restarted after Internet connection was serviced.
No Date Time UT Location Activity Comments
1377 2021-06-27 14:57 ORHO Seismo Swap T3762 (3T) removed, T3A81 (3ESP) installed.
1378 2021-06-27 15:00 ORHO Seismo Check Mass-centered OK.
1379 2021-06-27 22:48 ORHO Comments Auto Mass-centering is turned on, on Red at 4 V.
1380 2021-06-18 23:47 OTRO Hardware List Cyg391 Trd466 CMG-40T T4767. SPS 40, Gain 1.
1381 2021-06-18 23:47 OTRO New Station Code Start recording in OTRO ringbuffers.
1382 2021-06-18 23:47 OTRO Site Geology Glacial Moraine (similar to NWT) about 100 m from river
1383 2021-06-18 23:47 OTRO VSAT Start Xpol 5, Co-pol 35
1384 2021-06-19 00:33 OTRO SR Change SR changed to 100 sps, Gain 1.
1385 2021-08-11 19:56 OTRO DC Filter Check Unchanged from ON at 0.001 Hz.
1386 2021-08-11 19:56 OTRO Gain Check Software gain unchanged from 1.0.
1387 2021-10-29 16:30 OTRO Service Visit Just to check VSAT. Solar panel racks and antenna loose, tightened. Added bird deterent on feed horn. Local person checks once in a while.
1388 2021-10-29 16:32 OTRO VSAT Check Before 30/0, after 30/0.
1389 2021-11-03 17:35 OTRO Firmware Upgrade Trident 1.7210 in A and B. B is active.
1390 2021-11-24 08:11 OTRO Temp Halt Stopped due to heavy snow & ice on antenna?
1391 2021-11-24 19:20 OTRO VSAT Check Increased VSAT power at hub by 6 dB and changed modulation to 64 KB to get the site back on.
1392 2021-11-24 19:42 OTRO VSAT Check Remote Tx power too low.
No Date Time UT Location Activity Comments
1393 2021-11-25 19:01 OTRO VSAT Check Restored VSAT power at Hub to original level of 2dB and modulation to 112KB.
1394 2021-11-25 19:03 OTRO Comments VSAT is back to normal, all appear OK.
1395 2021-08-18 14:00 OTRO Service Visit To swap seismometer. Site condition good.
1396 2021-08-18 14:00 OTRO VSAT Check Co/x at 29/0, unchanged.
1397 2021-08-18 14:30 OTRO Seismo Swap ESP T3H24 installed, 40T T4767 removed.
1398 2021-08-18 15:06 OTRO Cold-Boot Standard procedure.
1399 2021-08-18 15:07 OTRO Restart Station restarted, recording on Ottawa Hub Naqs2.
1400 2021-08-18 16:02 OTRO Gain Check Unchanged at 1.0, default for FedNor1 sites.
1401 2021-08-18 16:03 OTRO Comments Hunters camping nearby.
1402 2021-08-18 16:03 OTRO DC Filter Check Unchanged at on at 0.001 Hz.
1403 2021-08-18 16:04 OTRO Comments Site gurad hired to clear snow from dish in the winter, watch site.
1404 2021-08-18 16:07 OTRO Restart Station recording on Ottawa Hub Naqs2 with the new seismometer.
1405 2021-11-21 13:59 OTTAWA Hub Login NTP server is now 132.156.100.63.
1406 2021-11-21 14:10 OTTAWA Hub Login Firewall NTP server set to 132.156.100.63.
1407 2021-11-26 17:44 OTTAWA Carina Hubs Carina 62 and 63 were powered down briefly to switch to temp gps antenna.
1408 2021-11-26 19:25 OTTAWA Carina Hubs Carina 62 and 63 were powered down briefly to switch to new gps antenna.
No Date Time UT Location Activity Comments
1409 2021-11-26 20:06 OTTAWA Carina Hubs Original GPS antenna cables were found to have been damaged during Hub installation. All cables were replaced. Improvements observed in GPS S/N on all Carinas.
1410 2021-08-02 21:45 OTTAWA Comments Naqs1 is down.
1411 2021-08-03 00:57 OTTAWA Comments Naqs1 restarted after system fixup.
1412 2021-08-26 18:10 OTTAWA Comments Tx power adjusted to 28.4 dB as per Telesat contract. Sent a test signal with BPSK32_1_2 and power of 39 to reach Tx gain of 4.75 for this calibration of power.
1413 2021-11-19 00:00 OTTAWA Comments Activity log is out of date. Need update.
1414 2022-03-29 16:16 OTTAWA Comments DataServer Version 1.05.01 Running on Naqs1, port 29001.
1415 2022-03-29 16:17 OTTAWA Comments NaqsServer 1.90 continues to run.
1416 2022-03-29 16:18 OTTAWA Comments NptoNMXP Version 1.00.08 is running. Converting Taurus data to NaqsStreams.
1417 2022-03-29 16:19 OTTAWA Comments Data from Taurus 204 in name 204T are streamed to Naqs1.
1418 2022-04-08 16:58 OTTAWA Restart Re-point of Telesat antenna completed, Carina 92 restarted.
1419 2022-04-12 15:00 OTTAWA Re-Point SES-AM antenna re-pointed. Improved Co/X from 27/0 to 30/0 dB.
1420 2022-04-15 17:48 OTTAWA Comments Multi-cast IP on Ottawa Hub6, Carina 63 was incorrect at 224.1.1.1, corrected to 225.1.1.1
1421 2022-05-04 18:00 OTTAWA Restart Test vault with Taurus 204 restarted in UBC-type vault.
1422 2022-05-04 18:10 OTTAWA Seismo Check ESP s/n T3F12 is installed in Taurus vault.
1423 2022-05-05 11:40 OTTAWA Comments Taurus 204 with ESP running ok. Mass-centered.
1424 2021-10-12 23:21 OTTAWA Comments New (third) 3.8 M Hub antenna installed. Served by SatMex.
No Date Time UT Location Activity Comments
1425 2021-06-02 18:05 OTTAWA Comments Activity File moved to new home.
1426 2021-08-02 14:00 OTTAWA Hub Swap Linux firewall replaced with a pair of SonicWall Pro 300 firewalls.
1427 2021-09-07 17:52 OTTAWA Comments Data directories on drive D: on Naqs Servers 1 and 2 cleaned up. Interent and Test site data ringbuffers moved to drive C: due to lack of space on D:
1428 2021-09-07 17:54 OTTAWA Comments SOH ringbuffers for Internet and Test sites remain on drive D:.
1429 2021-09-08 22:00 OTTAWA Comments Storm in Ottawa caused temporary outage on GBLN, GLWN, ILKN, MCKN, SNPN and SILO. Indicating TX power on these is low?
1430 2021-12-06 19:53 OTTAWA VSAT Check Bad burst Rx may continue on Telesat, even after repoint.
1431 2021-12-06 19:53 OTTAWA VSAT Check Telesat Antenna co-pol improved by 0.5 dB, isolation is now 32.5 dB.
1432 2022-01-01 00:00 OTTAWA MT Station New Scaling corrections started for Ontario MT stations.
1433 2022-02-06 18:00 OTTAWA Comments Naqs1 and Naqs2 server MTU restored to 1500.
1434 2022-02-15 17:15 OTTAWA Battery Check All 10 UPS batteries were replaced.
1435 2022-02-15 17:20 OTTAWA Comments Inverter glitch due to power disconnect.
1436 2022-03-15 17:34 OTTAWA VSAT Check Satmex antenna re-point, x-pol: before 25, now 32 dB.
1437 2021-08-07 23:24 OTTAWA HUB Comments ReTX turned on on Naqs1. It was already on on Naqs2.
1438 2021-08-09 17:59 OTTAWA HUB Comments Naqs1 was down for service from 13:59 to 14:35 ET.
1439 2021-08-26 18:00 OTTAWA HUB VSAT Start Started Anik E2R Service on second 3.8 m antenna, Carina 92.
1440 2021-08-26 18:10 OTTAWA HUB VSAT Check Co-Pol of 35 dB achieved.
No Date Time UT Location Activity Comments
1441 2022-01-31 17:30 OTTAWA HUB Reboot Naqs2 was rebooted after service maintenance.
1442 2022-02-01 18:31 OTTAWA HUB Reboot Naqs1 war rebooted after service maintenance.
1443 2022-02-01 19:12 OTTAWA HUB Reboot NaqsServer 1.9 and required DLL installed on Naqs1. Restarted at 19:12 UT.
1444 2022-03-29 15:45 OTTAWA HUB Comments DataServer Stopped on Naqs2.
1445 2022-04-08 16:28 OTTAWA HUB Re-Point Improved Co-pol by 3 and X-Pol by 2 dB for Telesat antenna. Isolastion is now 35 dB.
1446 2021-05-19 15:00 OTTAWA HUB Re-Point SES-AM antenna re-pointed. Improved Co/X from 28/10 to 30/0 dB.
1447 2021-05-30 17:37 OTTAWA HUB Comments NaqsServer 1.90 is installed and started on Naqs2.
1448 2021-06-20 12:45 OTTAWA HUB Reboot Naqs2 needed a re-boot to clear some data TX problems.
1449 2021-07-29 13:37 OTTAWA HUB Comments Dataserver restarted on Ottawa Hub Naqs1.
1450 2021-08-03 15:22 OTTAWA HUB Comments Dataserver restarted on Ottawa Hub Naqs1.
1451 2021-08-04 19:49 OTTAWA HUB Comments Dataserver restarted on Ottawa Hub Naqs1.
1452 2021-08-10 21:58 OTTAWA HUB Comments Naqs1, 2 and Dataserver restarted.
1453 2021-12-06 19:51 OTTAWA HUB VSAT Check SES-AM Antenna steady at 29.5 dB x-pol.
1454 2022-04-18 15:30 OTTAWA-HUB Comments Cables, switches and Internet connection upgraded. 8 and 24 managed switches installed.
1455 2022-05-02 18:00 OTTAWA-HUB VSAT Check Satmex antenna re-point, x-pol: improved by 16 dB, now it is 35 dB.
1456 2022-05-07 15:58 OTTAWA-HUB VSAT Check Satmex antenna re-point, x-pol: before 26, now 32 dB.
No Date Time UT Location Activity Comments
1457 2022-05-12 12:35 OTTAWA-HUB Comments NPtoNMP software v 2.2.1 installed on Naqs1.
1458 2021-08-20 OTTO New Station Start This test station is not registered in Station Book.
1459 2021-09-17 02:20 Ontario Resynch check Stations in 133a-133w: Trident Resynchronization mode = Discard samples
1460 2021-06-09 00:21 Ottawa Hub Start Start recording at Ottawa Hub.
1461 2021-07-07 17:14 Ottawa Hub Down Hub and Naqs shut down 17:14-17:35 to install AC switch on Inverter-charger.
1462 2021-11-09 23:45 P20T New Station Start Start recording in P20T ringbuffers at HSLB.
1463 2021-09-06 17:00 P21T New Station Start Start recording in P21T ringbuffers at PECO.
1464 2021-09-06 17:00 P21T New Station Start Start recording in P21T ringbuffers at PECO.
1465 2021-07-26 17:00 P22T New Station Start Start recording in P22T ringbuffers at MCKN with CMG-40T.
1466 2021-07-20 21:00 P23T New temp start Started recording in P23T ringbuffers at ACKN.
1467 2021-07-19 00:00 P24T New temp start Started recording in P24T ringbuffers at EKTN.
1468 2021-07-24 22:00 P26T New Station Start Start recording in P26T ringbuffers at LGSN with CMG-40T. 2021-07-25 00:00 LGSN New Station Code Start recording in LGSN ringbuffers.
1469 2021-07-26 00:00 P27T New Station Start Start recording in P27T ringbuffers DVKN with CMG-40T.
1470 2021-07-23 01:00 P28T New temp start Started recording in ringbuffers P28T at LDGN.
1471 2021-07-21 19:00 P29T New temp start Started recording in ringbuffers P29T at YMBN.
1472 2021-07-19 22:00 P30T New Station Start Start recording in P30T ringbuffers at COWN.
No Date Time UT Location Activity Comments
1473 2021-07-23 15:00 P31T New Station Start Start recording in P31T ringbuffers at GLWN.
1474 2021-11-18 18:00 P32T New Station Start Start recording in P32T ringbuffers at ANMB.
1475 2021-07-30 00:00 P33T New Station Start Start recording in P33T ringbuffers CAMN.
1476 2021-07-31 22:00 P34T New Station Start Start recording in P34T ringbuffers at IHLN.
1477 2021-07-31 18:00 P35T New Station Start Start recording in P35T ringbuffers at SNPN.
1478 2021-08-03 18:00 P36T New Station Start Start recording in P36T ringbuffers at MLON.
1479 2021-08-04 13:00 P37T New Station Start Start recording in P37T ringbuffers at KNDN.
1480 2021-08-02 22:00 P39T New Station Start Start recording in P39T ringbuffers BOXN.
1481 2021-07-29 17:00 P40T New Station Start Start recording in P40T ringbuffers at NODN.
1482 2021-08-03 21:00 P41T New Station Start Start recording in P41T ringbuffers at MGTN.
1483 2021-09-07 22:00 P42T New Station Start Start recording in P42T ringbuffers at LINO.
1484 2021-08-27 15:00 P43T New Station Start Start recording in P43T ringbuffers at DELO.
1485 2021-08-21 23:45 P44T New Station Start Start recording in P44T ringbuffers at MPPO.
1486 2021-08-24 23:00 P45T New Station Start Start recording in P45T ringbuffers at BANO.
1487 2021-09-11 21:00 P46T New Station Start Start recording in P46T ringbuffers at ALGO.
1488 2021-09-10 16:00 P47T New Station Start Start recording in P47T ringbuffers at BRPO.
No Date Time UT Location Activity Comments
1489 2021-11-20 22:00 P49T New Station Start Start recording in P49T ringbuffers at TWKB.
1490 2021-11-25 19:40 P50T New Station Start Start recording in P50T ringbuffers at TWGB.
1491 2021-11-16 23:37 P51T New Station Start Start recording in P51T ringbuffers at CLVB.
1492 2021-11-14 22:45 P53T New Station Start Start recording in P53T ringbuffers at PIMB.
1493 2021-10-11 18:00 P54T New Station Start Start recording in P54T ringbuffers at CLPO.
1494 2021-11-12 00:21 P55T New Station Start Start recording in P55T ringbuffers at HRMB.
1495 2021-11-27 00:10 P56T New Station Start Start recording in P56T ringbuffers at TWBB.
1496 2021-12-04 17:29 P57T New Station Start Start recording in P57T ringbuffers at SSIB.
1497 2021-11-13 22:00 P58T New Station Start Start recording in P58T ringbuffers at PEMO.
1498 2022-05-14 20:30 P59T New Station Start Start recording in P59T ringbuffers at YOUB.
1499 2022-02-22 21:30 P60T New Station Start Start recording in P60T ringbuffers at SILB. SILB New Station Code Start recording in SILB ringbuffers.
1500 2021-05-17 01:49 P62T New Station Start Start recording in P62T ringbuffers at TSJB.
1501 2021-05-29 21:00 P75T New Station Start Start recording in P57T ringbuffers at HBDB.
1502 2021-07-29 14:29 P81T New Station Start Start recording in P81T ringbuffers at TOBO.
1503 2021-07-25 17:25 P82T New Station Start Start recording in P82T ringbuffers at CLWO.
1504 2021-07-27 16:47 P83T DC Filter Check On at 0.001 Hz.
No Date Time UT Location Activity Comments
1505 2021-07-27 16:47 P83T Gain Check Station Started at software gain of 0.5, sample rate 100 sps.
1506 2021-07-27 16:47 P83T Hardware List CYG417 TRD501 CMG-40T 4567, SSPB A0305A
1507 2021-07-27 16:47 P83T New Station Start Start recording in P83T ringbuffers at KLBO.
1508 2021-07-27 16:47 P83T Site Geology Bedrock
1509 2021-07-27 16:47 P83T VSAT Start Co-pol 25 X-pol 0
1510 2021-08-21 16:17 P84T DC Filter Check On at 0.001 Hz.
1511 2021-08-21 16:17 P84T Gain Check Station Started at software gain of 0.5, sample rate 100 sps.
1512 2021-08-21 16:17 P84T Hardware List CYG180 TRD319 CMG-3ESP T3D83, SSPB A0303A
1513 2021-08-21 16:17 P84T New Station Start Start recording in P84T ringbuffers at PLIO.
1514 2021-08-21 16:17 P84T Site Geology Limestone with shallow overburden.
1515 2021-08-21 16:17 P84T VSAT Start Co-pol 25 X-pol 0
1516 2021-10-02 18:58 P85T Hardware List Cyg 175, Trd 073, Trillium 0075, SSPB A0171A.
1517 2021-10-02 18:58 P85T New Station Start Start recording in ALFO ringbuffers. Cold Boot ok.
1518 2021-10-02 18:58 P85T New Station Start Start recording in P85T ringbuffers at ALFO.
1519 2021-10-02 18:58 P85T Site Geology Bedrock
1520 2021-10-02 18:58 P85T VSAT Start Co-pol 31 X-pol 1.
No Date Time UT Location Activity Comments
1521 2021-11-10 21:20 P86T New Station Start Electronics from HGVO under test at UWO.
1522 2022-05-04 18:05 P99T New Station Start Started recording at TORO in temp P99T ringbuffers.
1523 2022-05-04 18:10 P99T VSAT Check Co/X Pol 35/2 dB.
1524 2021-10-31 14:41 PCTO Firmware Upgrade Installed Trident 1.7210 in A.
1525 2021-09-06 19:10 PECO New Station Code Start recording in PECO ringbuffers.
1526 2021-09-06 19:10 PECO New Station Code Start recording in PECO ringbuffers.
1527 2021-09-06 19:10 PECO Seismo Check Started the station with CMG-40T.
1528 2021-09-16 22:15 PECO DC Filter Change From off to on. At 0.001 mHz corner frequency.
1529 2021-09-16 22:15 PECO Gain Change From 1.048576 to 0.1.
1530 2021-09-20 17:04 PECO Gain Check Software gain was 0.1, IIR filter on at 0.001
1531 2021-09-20 17:04 PECO Seismo Swap Changed seismometer from CMG-40T to CMG-3ESP.
1532 2021-11-13 22:45 PEMO New Station Code Start recording in PEMO ringbuffers.
1533 2022-02-14 16:50 PEMO Hardware Swap Trident 076 replaced with TRD 122.
1534 2022-02-19 23:30 PEMO Hardware Swap Cygnus 174 replaced with Cygnus 184.
1535 2022-05-08 15:34 PEMO Site Visit Cygnus installed in battery box. more batteries? GPS cable in conduit. Battery box dug in a bit more.
1536 2022-05-08 15:34 PEMO Site Visit Maintenance visit.
No Date Time UT Location Activity Comments
1537 2022-05-08 15:34 PEMO VSAT Check Co-pol 34, X-pol 4dB. Adjusted, co-pol 33, X-pol 2 dB. IFL loss set to 2 dB.
1538 2021-05-26 14:30 PEMO Service Visit To upgrade seismometer.
1539 2021-05-26 14:31 PEMO Seismo Swap CMG-40T s/n T4185 removed. ESP s/n T3J31 installed.
1540 2021-05-26 14:32 PEMO Comments Evidence of mice in seismometer vault.
1541 2021-05-26 14:33 PEMO Seismo Check Mass centered and calibrated ok.
1542 2021-05-26 14:34 PEMO Gain Check Unchaged at 0.5.
1543 2021-05-26 14:35 PEMO DC Filter Check Unchaged from On at 1 mHz.
1544 2021-05-26 14:36 PEMO Firmware Upgrade Trident upgraded to 1.7201, epld 1.23. TS was running current v 1.23. Unable to upgrade Libra, kept old version 5.42 in A. Newer 5.6201 loaded in B but fails to test.
1545 2022-02-26 PGCB New Station Start First station of BC array. AC powered.
1546 2021-07-19 PGCB Closed Finished with training station.
1547 2022-02-20 11:00 PGCB VSAT Start Restarted test station, X/Co: 2/31dB. CMG40T 4688, cyg 172, trd 369
1548 2021-11-14 23:05 PIMB New Station Code Start recording in PIMB ringbuffers.
1549 2021-06-18 PIMB Service Visit All OK. Screw fixed in junction box.
1550 2021-07-28 01:38 PIMB Closed End of operation.
1551 2021-06-15 15:10 PKLO New Station Code Started on Naqs2, Ottawa Hub, in PKLO buffers.
1552 2021-06-15 15:14 PKLO New Station Code Started on Naqs1, Ottawa Hub, in PKLO buffers.
No Date Time UT Location Activity Comments
1553 2022-01-30 PKRO New Station Start SOSN/POLARIS. AC Powered.
1554 2021-08-13 18:44 PKRO Firmware Upgrade Upgraded to Libra V5.20, TS V1.20, TR V1.50 EPLD V1.08
1555 2021-09-17 02:00 PKRO DC Filter Change From off to on. At 0.001 mHz corner frequency.
1556 2021-09-17 02:00 PKRO Gain Change From 1.048576 to 0.1.
1557 2021-09-23 20:12 PKRO Firmware Upgrade Upgraded to Libra Firmware version 5.31.
1558 2022-05-05 22:15 PKRO Hardware Swap Now running CYG172 TRD66.
1559 2022-05-05 22:15 PKRO Site Visit Maintenance visit.
1560 2021-10-19 15:30 PKRO Service Visit Station was down due to bad Cygnus.
1561 2021-10-19 15:56 PKRO Hardware Swap Cyg 222 installed, 172 removed.
1562 2021-06-17 18:00 PKRO Service Visit To improve VSAT power.
1563 2021-06-17 18:10 PKRO Comments Weeds were blocking feedhorn and antenna. All cleaned up. Power went up by 5 dB.
1564 2021-06-17 18:20 PKRO VSAT Check Antenna re-pointed. X/Co: 0/30 dB.
1565 2021-06-17 18:21 PKRO Comments No additional improvement in power level.
1566 2021-06-17 18:22 PKRO Gain Check Unchanged at 0.5
1567 2021-06-17 18:23 PKRO DC Filter Check Unchanged at on and 0.001 Hz.
1568 2021-06-17 18:25 PKRO Firmware Upgrade All firmeware are up-to-date.
No Date Time UT Location Activity Comments
1569 2021-06-17 18:40 PKRO Cold-Boot Cold boot is standard procedure at the end of visit.
1570 2021-06-17 18:50 PKRO Comments Covered ground with rubber mat to prevent weeds growing.
1571 2021-09-01 18:00 PKRO Service Visit To fix battery charger. Fuse was blown in charger and charger was bust. Replacing fuse did not fix the problem. Charger replaced.
1572 2021-09-01 18:15 PKRO Battery Check Old, likely bad, DC-Type batteries removed, new AC-Type batteries installed.
1573 2021-09-01 19:14 PKRO Restart Station back on. Shutting down again to check VSAT.
1574 2021-09-01 19:40 PKRO VSAT Check X pol 0 dB, co-pol not measured due to modulated signal.
1575 2021-09-01 20:56 PKRO Gain Check Unchaged at 0.5.
1576 2021-09-01 20:57 PKRO DC Filter Check Unchaged at on at 0.001 Hz.
1577 2021-09-01 20:58 PKRO Comments Fuses left at the site for battery charger.
1578 2021-06-17 17:00 PKRO Re-Point X/Co 30/0, 6 dB improved by clearing weeds and re-pointing.
1579 2021-09-16 21:57 PLBC New Station Code PLCB is renamed to PLBC, Ottawa, Naqs2 data starts.
1580 2021-09-16 22:05 PLBC Comments Ottawa Hub, Dataserver restarted.
1581 2021-09-16 22:05 PLBC Name Change PLCB is renamed to PLBC, Ottawa, Naqs1 data starts.
1582 2021-06-11 21:26 PLBC Comments Stopped POLARIS VSAT, restarted as CNSN.
1583 2021-09-03 21:11 PLCB Comments Renamed to PLBC, later.
1584 2021-09-03 21:11 PLCB DC Filter Check Off.
No Date Time UT Location Activity Comments
1585 2021-09-03 21:11 PLCB Gain Check 1.0 at 100 sps.
1586 2021-09-03 21:11 PLCB Hardware List Cyg 277, Trd, 276, esp s/n t3n05
1587 2021-09-03 21:11 PLCB Require GPS Lock On.
1588 2021-09-03 21:11 PLCB VSAT Check 30 db Isolation on Telesat.
1589 2021-09-16 21:55 PLCB Closed Changed name to PLBC.
1590 2021-08-22 16:00 PLIO New Station Code Start recording in PLIO ringbuffers.
1591 2021-08-04 17:05 PLIO SSPB Check Damaged sspb repaired.
1592 2021-08-04 17:05 PLIO Service Visit To restore dead station.
1593 2021-08-04 17:10 PLIO Hardware Swap New SSPB is A0165A, LNB 12200294.
1594 2021-08-04 17:20 PLIO VSAT Check 26/0.
1595 2021-08-04 17:30 PLIO Seismo Check Seismo is about 15 cm above water level.
1596 2021-08-04 17:40 PLIO Comments First site serviced by Bicyle (KA, BD).
1597 2021-11-01 19:27 PLIO MT Station Started MT recording, live to London Hub.
1598 2022-05-13 20:30 PLVO DC Filter Check DC filter is on at 0.001 HZ, startup.
1599 2022-05-13 20:30 PLVO Gain Check Gain is at 0.5 for startup.
1600 2022-05-13 20:30 PLVO Hardware List CYG541, TS694, TRD733, SSPB A0172A, Seis. ESP T3H25 (from UWO).
No Date Time UT Location Activity Comments
1601 2022-05-13 20:30 PLVO VSAT Check Co/X 32/.4 dB.
1602 2022-05-13 22:30 PLVO New Station Code Started recording at PLVO ringbuffers.
1603 2021-10-18 16:00 PLVO Service Visit To connect MT station.
1604 2021-10-18 16:10 PLVO Comments MT battery box and recorder missing.
1605 2021-10-18 16:20 PLVO VSAT Check 25/0 dB, improved by 3.5 dB.
1606 2021-10-26 19:36 PLVO MT Station Started MT recording, live to London Hub.
1607 2021-12-11 17:00 PLVO MT Station North e-line was cut by animale bite. Added 25' splice, soldered, heat tape applied and bagged in silicon.
1608 2021-06-18 20:48 PNPO New Station Code Started recording at Ottawa Naqs2 in PNPO ringbuffers.
1609 2021-06-18 20:57 PNPO Comments Sensor Calibration ok.
1610 2021-06-18 21:03 PNPO New Station Code Started recording at Ottawa Naqs1 in PNPO ringbuffers.
1611 2021-06-18 21:04 PNPO Comments Observation: If Cyg/Trd do not match in naqs.stn, then the code attached to the correct Trident will tranport the data.
1612 2021-06-20 23:00 PNPO Comments Stopped sending GPS soh data. Why?
1613 2022-02-05 PTCO New Station Start DC (Solar Panels).
1614 2022-02-07 PTCO Seismo Swap Changed from CMG-3ESP to CMG-40T.
1615 2022-02-21 PTCO Seismo Swap Changed from CMG-40T to CMG-3ESP.
1616 2021-08-13 23:42 PTCO Firmware Upgrade Upgraded to Libra V5.20, TS V1.20, TR V1.50 EPLD V1.08
No Date Time UT Location Activity Comments
1617 2021-09-17 00:15 PTCO DC Filter Change From off to on. At 0.001 mHz corner frequency.
1618 2021-09-17 00:15 PTCO Gain Change From 1.048576 to 0.1.
1619 2021-09-20 19:30 PTCO Firmware Upgrade Upgraded to Libra Firmware version 5.31.
1620 2022-04-15 PTCO Power Upgrade Added 2 more batteries.
1621 2022-04-15 PTCO Site Visit Maintenance visit.
1622 2022-04-15 PTCO Wildlife Alert Snake visit (below battery box). Wasps in battery box.
1623 2021-10-31 18:07 PTCO Closed Moving to more secure location (BUKO).
1624 2021-10-31 18:08 PTCO Comments 2 solar panel stolen, one shattered.
1625 2021-10-31 18:09 PTCO Comments Old large seismo valut and fence left at the site.
1626 2021-07-11 19:18 PTCO Reopened In original location. Started recording in PTCO ringbuffers at Ottawa Hub on both NaqsServers.
1627 2021-07-11 19:23 PTCO Hardware List CYG171,TRD070, TS034, SSPB a0307A, SEISM. 40T, 4401
1628 2021-07-11 19:30 PTCO VSAT Check Co/x 35/3 dB.
1629 2021-07-11 19:32 PTCO Cold-Boot Will screen capture.
1630 2021-07-11 19:32 PTCO Comments Lots of people on road about 50 m from seism. Station visible from the road.
1631 2021-07-11 19:32 PTCO Site Geology Solid bedrock.
1632 2021-07-11 19:42 PTCO Gain Check Restarted with Ontario default gain of 0.5 at 100 sps.
No Date Time UT Location Activity Comments
1633 2021-07-11 19:43 PTCO DC Filter Check Restarted with Ontario default DC filter on at 0.001Hz.
1634 2021-07-21 17:55 PTCO Service Visit Stoped to change seismo and trident.
1635 2021-07-21 18:00 PTCO Seismo Swap ESP T3K14 installed. CMG-40T 4401 removed.
1636 2021-07-21 18:10 PTCO DC Filter Check DC filter is on at 0.001 HZ, unchanged.
1637 2021-07-21 18:10 PTCO Gain Check Gais is 0.5 is unchanged.
1638 2021-07-21 18:10 PTCO Hardware Swap Trident 070 removed, TRD747 installed.
1639 2021-07-21 18:47 PTCO Comments Cygnus ringbuffers restarted.
1640 2021-07-21 18:49 PTCO Cold-Boot Cold Boot.
1641 2021-07-21 18:50 PTCO Comments Problem restarting VSAT after cold boot. Source unknown. Increasing TX power did not help.
1642 2021-07-21 19:10 PTCO Restart All OK and station started.
1643 2021-07-21 19:12 PTCO Comments Trident gain and DC filter checked again, they are unchanged after cold boot.
1644 2022-01-25 16:10 PTCO Service Visit To restore power and check VSAT. Ice cleanded off antenna and solar panels.
1645 2022-01-25 16:20 PTCO VSAT Check Co/x at 28/0 dB, no adjustment needed.
1646 2022-01-25 16:30 PTCO Power Check Loose wire fixed in charge controller.
1647 2022-01-25 19:03 PTCO Restart Station is back up. Signal looks good.
1648 2022-04-30 15:30 PTCO Closed End of Operation.
No Date Time UT Location Activity Comments
1649 2021-09-04 22:11 QILN New Station Code Started recording at Ottawa Naqs2 in QILN ringbuffers.
1650 2021-09-04 22:14 QILN New Station Code Started recording at Ottawa Naqs1 in QILN ringbuffers.
1651 2021-09-04 22:15 QILN Gain Check Gain of 1.0 at 40 sps.
1652 2021-09-04 22:16 QILN DC Filter Check DC filter not on.
1653 2021-09-04 22:54 QILN VSAT Check Eirp increase up from 4 to 6 dB to improve TX power. TX DAC increased from 10.50 to 16.25.
1654 2021-09-04 22:56 QILN Hardware List Cygnus 555, Trident 737, TimeServer 675. Seismo. ESP T3K94, SSPB A0397A.
1655 2021-09-03 21:04 RAMB New Station Code South West of Quesnel, Ramsey Rd, BC, Ottawa Hub, Telesat.
1656 2021-09-03 21:05 RAMB Hardware List Cyg182, Trd63, TS53, seismometer 40T, s/n T4769, sspb A0160A.
1657 2021-09-03 21:06 RAMB DC Filter Check Off.
1658 2021-09-03 21:06 RAMB Gain Check 1.0 at 100 sps.
1659 2021-09-03 21:07 RAMB Battery Check 6 new batteries, 8 panels.
1660 2021-09-03 21:07 RAMB Require GPS Lock On.
1661 2021-09-03 21:07 RAMB VSAT Check 31 db isolation.
1662 2021-09-03 21:08 RAMB Site Geology Silt.
1663 2021-06-13 17:47 RDLO New Station Code Started recordoring on Naqs2 in RDLO ringbuffers.
1664 2021-06-13 17:51 RDLO New Station Code Started recording on Naqs1 in RDLO ringbuffers.
No Date Time UT Location Activity Comments
1665 2021-07-07 12:04 RDLO Seismo Check HHN component was noisy yesterday, looks ok today.
1666 2021-09-29 16:00 RDLO Closed Moved to new location RLKO on Nov. 19, 2004.
1667 2021-11-19 15:36 RLKO New Station Code Started recording at both Naqs in Ottawa Hub.
1668 2021-08-10 21:30 RLKO Service Visit To fix lack of data problem.
1669 2021-08-10 21:31 RLKO Hardware Swap Cyguns 621, trd 867, seismo T3P07 installed.
1670 2021-08-10 21:33 RLKO Hardware List Cygnus 536, trd 695 removed.
1671 2021-08-10 21:34 RLKO Comments Old cygnus was not providing power on the bus. Cable was tested ok, cygnus faulty.
1672 2021-08-10 21:35 RLKO VSAT Check 30/0 dB. No adjustment needed.
1673 2021-08-10 21:47 RLKO Comments Site undisturbed.
1674 2021-08-10 21:57 RLKO Comments Finished service. New ESP seismo mass-centered ok.
1675 2021-08-10 22:51 RLKO Gain Check 100 SPS, Gain 0.5
1676 2021-08-10 22:52 RLKO Comments Require GPS lock to lock is on.
1677 2021-08-10 22:52 RLKO DC Filter Check DC filer on at 1 mHz.
1678 2021-08-11 14:55 RLKO Seismo Swap Bad HHE component with new T3P07 seismo, replaced with original T3762, 3T seismo.
1679 2021-08-11 15:02 RLKO Comments End of Service visit. Site is back to normal.
1680 2021-11-01 21:00 RLKO Comments New seismometer T3P06 has electonics mass unlock.
No Date Time UT Location Activity Comments
1681 2021-11-01 21:00 RLKO Seismo Swap T3762 removed, ESP T3P06 installed.
1682 2021-11-01 21:00 RLKO Service Visit To replace seismometer with bad ns component.
1683 2021-10-03 13:50 RLKO Service Visit TimeServer failure, both Cygnus and Trident replaced.
1684 2021-10-04 13:50 RLKO VSAT Check X-pol 29 dB.
1685 2021-08-05 17:59 RSPO New Station Code Started at Ottawa Hub, Naqs2 ringbuffers.
1686 2021-08-05 18:03 RSPO New Station Code Started at Ottawa Hub, Naqs1 ringbuffers.
1687 2021-10-13 19:56 SABG New Station Start Started recording from Sable Island at London Hub.
1688 2021-10-19 19:12 SABG Gain Check Gain 1.0, SPS 100.
1689 2021-10-19 19:13 SABG Comments Small 1.2 m antenna used.
1690 2021-10-19 19:13 SABG DC Filter Check Off.
1691 2021-10-19 19:13 SABG VSAT Check 27/0 dB.
1692 2021-10-19 19:14 SABG Seismo Check ESP, s/n TBD at 43.93124 -60.00843 0.
1693 2021-08-25 17:34 SABG Comments Switched from SES-AM to Telesat, Carina 116.
1694 2021-08-25 17:35 SABG VSAT Check Signal from this 1.2 m antenna is strong at ERIP increase of 0 dB.
1695 2021-08-24 15:42 SEDN New Station Code Sedna, Hudson Bay, on Ottawa Hub Naqs2, Telesat.
1696 2021-08-24 15:52 SEDN New Station Code Sedna, Hudson Bay, on Ottawa Hub Naqs1, Telesat.
No Date Time UT Location Activity Comments
1697 2021-08-24 17:08 SEDN Hardware List Cyg245, Trd423, TS190, ESP T3A90, SSPB 1645-0503, redeployed from LDGN.
1698 2021-08-24 17:09 SEDN Site Geology Bedrock at the end of runway strip.
1699 2021-08-24 17:10 SEDN DC Filter Check Off.
1700 2021-08-24 17:10 SEDN Gain Check 1.0 at 40 sps.
1701 2021-08-24 17:10 SEDN Require GPS Lock On.
1702 2021-08-24 17:11 SEDN Battery Check 16 new batteries, 12 solar panels.
1703 2021-08-24 17:11 SEDN VSAT Check 32 dB isolation.
1704 2021-07-04 19:09 SILB Closed End of Operation.
1705 2021-06-09 12:31 SILO Hardware List Cyg 378, TRD365, ESP T3E88. SPS 100, Gain 1.
1706 2021-06-09 12:31 SILO New Station Code Start recording in SILO ringbuffers.
1707 2021-06-09 12:31 SILO Site Geology Hard solid bedrock.
1708 2021-06-09 12:31 SILO VSAT Start X-pol 0, Co-pol 27 dB.
1709 2021-08-11 19:24 SILO DC Filter Check Unchanged from ON at 0.001 Hz.
1710 2021-08-11 19:24 SILO Gain Check Software gain unchanged from 1.0.
1711 2022-02-21 00:01 SILO VSAT Check Modulation increased from 64 to 112KB to improve bandwidth use.
1712 2022-05-05 13:00 SILO VSAT Check IFL loss faked from 2 to 5 dB. Eirp increase is at 9 dB in attempt to reduce ReTX.
No Date Time UT Location Activity Comments
1713 2022-04-12 18:40 SILO VSAT Check Changed modulation from 112 to 64KB to improve transmission.
1714 2021-08-23 17:30 SNO New Station Start First Taurus Internet station started at SNO Lab, Sudbury.
1715 2021-07-31 19:15 SNPN New Station Code Start recording in SNPN ringbuffers.
1716 2021-08-13 23:28 SNPN Firmware Upgrade Upgraded to Libra V5.20, TS V1.20, TR V1.50 EPLD V1.08
1717 2021-09-14 21:30 SNPN SSPB Check Corrected SSPB temperature profile from flat 45 dB to manufacturer specs.
1718 2021-09-20 00:49 SNPN Firmware Upgrade Upgraded to Libra Firmware version 5.31.
1719 2021-07-19 19:54 SNPN Before Re-Point Before any re-adjustment, co-pol was 28 and x-pol was 0 dB, indicating stable antenna base.
1720 2021-07-19 19:54 SNPN Feed Horn Box New Feed Horn protective box installed. New cables and conduits installed.
1721 2021-07-19 19:54 SNPN Firmware Upgrade Libra 5.6201 installed.
1722 2021-07-19 19:54 SNPN Re-Point After re-pointing, co-pol was 32 with x-pol of 1 dB.
1723 2021-07-19 19:54 SNPN Service Visit Site was working OK. Stopped for service. Solar panel rack was loose, fixed.
1724 2021-07-19 20:45 SNPN Restart Calibrated ok, mass-centred ok.
1725 2021-09-27 22:13 SNPN Bird-protection No bird damage observed but additional tie-wraps were installed on feed arms.
1726 2021-09-27 22:13 SNPN Re-Point X-pol 0 Co-pol 20 dB.
1727 2021-09-27 22:13 SNPN Service Visit Visit initiated due to poor VSAT performance.
1728 2021-09-27 22:13 SNPN VSAT Check Bad co-pol due probably to shifting of antenna. X-pol 0 Co-pol 7dB.
No Date Time UT Location Activity Comments
1729 2021-09-27 22:36 SNPN Cold-Boot Restarted from cold boot. Eirp increase adjusted to 5 dB at Hub.
1730 2021-08-21 16:12 SNPN Service Visit To add MT.
1731 2021-08-21 16:13 SNPN Hardware Swap Moved from London to Ottawa Hub.
1732 2021-08-21 16:16 SNPN Restart Recording at Ottawa Hub on Naqs1 and Naqs2.
1733 2021-08-21 16:26 SNPN Comments MT installation in progress.
1734 2021-08-21 20:50 SNPN Comments MT data started. Temperature option of SOH monitor disconnected.
1735 2021-08-21 20:59 SNPN Comments MT team are digging electric lines.
1736 2021-08-23 15:11 SNPN Comments MT station was found to be in GPS mode. Switched to normal mode by Zoc telenet login, and then ^C command.
1737 2021-07-19 20:59 SNPN Service Visit MT electordes changed.
1738 2021-07-19 20:59 SNPN VSAT Check Repoint from 15/0 to 30/1 dB.
1739 2021-07-19 21:00 SNPN Comments GPS antenna connector was loose. Fixed and antenna mast installed.
1740 2021-07-19 21:01 SNPN Comments Large blase planned at 6 local time at Snap Lake.
1741 2021-07-19 21:01 SNPN Comments Staion looked good, couple of loose bolts fixed.
1742 2021-09-20 04:00 SNQN New Station Code Internet, Ottawa Hub, at Sanikiliuq, Hudson Bay.
1743 2021-09-20 04:10 SNQN Hardware List Taurus 207, ESP T3P17.
1744 2021-09-20 04:20 SNQN 40 sps, Gain 1, No DC filter, Require GPS Lock Checked.
No Date Time UT Location Activity Comments
1745 2021-09-20 04:25 SNQN Internet Station On VPN.
1746 2021-08-21 21:09 SRLN New Station Start At Sarcpa Lake old Dewline station, recording on Naqs2, Ottawa.
1747 2021-08-21 21:12 SRLN New Station Start At Sarcpa Lake old Dewline station, recording on Naqs1, Ottawa.
1748 2021-08-21 21:14 SRLN Hardware List Cyg244, Trd136, Seismometer Esp T3B37(confirm with pics).
1749 2021-08-21 21:15 SRLN Gain Check Gain of 1.0, sps 40.
1750 2021-08-21 21:16 SRLN DC Filter Check DC filter is not on.
1751 2021-08-21 21:17 SRLN Require GPS Lock Require GPS Lock is on.
1752 2021-08-21 21:18 SRLN Site Geology Solid bedrock.
1753 2021-08-21 21:19 SRLN Comments Antenna and parts of solar panel frames bolted to rock.
1754 2021-12-04 17:30 SSIB DC Filter Check Starting with DC filter on at 0.001 Hz.
1755 2021-12-04 17:30 SSIB Gain Check Starting with software gain of 0.1.
1756 2021-12-04 17:30 SSIB New Station Code Start recording in SSIB ringbuffers.
1757 2021-06-16 17:11 SSIB Closed End of Operation.
1758 2021-11-09 17:15 SSNO Internet Station Started at shallow SNO, Ottawa Hub Naqs1,2.
1759 2021-11-09 17:16 SSNO Hardware List Taurus 476. Trillium 120P s/n to be reported.
1760 2021-11-09 17:17 SSNO Gain Check Gain 1.0 at 200 sps.
No Date Time UT Location Activity Comments
1761 2021-11-09 17:18 SSNO DC Filter Check Off.
1762 2021-11-09 17:19 SSNO Require GPS to Lock on.
1763 2021-11-09 STCO New Station Start SOSN/POLARIS. AC Powered.
1764 2021-06-26 STCO Hardware Swap Replaced Cygnus 155/HRD 185 to Cygnus 182/Trident 82.
1765 2021-08-09 21:30 STCO Firmware Upgrade Upgraded to Libra V5.20, TS V1.20, TR V1.50 EPLD V1.08
1766 2021-09-16 21:55 STCO DC Filter Change From off to on. At 0.001 mHz corner frequency.
1767 2021-09-16 21:55 STCO Gain Change From 1.048576 to 0.1.
1768 2021-09-24 21:45 STCO Firmware Upgrade Upgraded to Libra Firmware version 5.31.
1769 2022-04-19 00:30 STCO Require GPS Lock New Instrument installed.
1770 2021-06-17 23:45 STCO VSAT Check SSPB mounting screw is damaged, cound not adjust SSPB. X/Co 35/2, 2 dB improvement.
1771 2021-06-17 23:50 STCO Service Visit New SSPB A0304A, LNB 16201509.
1772 2021-08-18 01:43 STLN New Station Start Stellar Camp, Hudson Bay, Ottawa Naqs2, Telesat.
1773 2021-08-18 01:44 STLN Hardware List Cyg 242, Trd 183, TS 138, ESP T3A67, redeployed LGSN.
1774 2021-08-18 01:45 STLN New Station Start Stellar Camp, Hudson Bay, Ottawa Naqs1.
1775 2021-08-18 01:45 STLN VSAT Check 36 dB isolation!
1776 2021-08-18 01:49 STLN Comments Dataserver restarted on Naqs1.
No Date Time UT Location Activity Comments
1777 2021-08-18 12:33 STLN DC Filter Check Off.
1778 2021-08-18 12:33 STLN Gain Check Gain of 1 at 40 sps.
1779 2021-08-18 12:33 STLN Require GPS Lock On.
1780 2021-08-18 15:05 STLN Site Geology On Esker deposits, 1 Km from camp, south of airport.
1781 2021-08-18 15:06 STLN Battery Check 12 New battereies.
1782 2021-08-18 15:07 STLN Hardware List SSPB A0028A.
1783 2021-09-05 12:46 SULB New Station Code South of Vanderhoof, BC, Ottawa Hub, Telesat.
1784 2021-09-05 12:47 SULB Hardware List Cyg219, Trd080, TS090, SSPB A0156A, 3ESP T3D96.
1785 2021-09-05 12:49 SULB DC Filter Check Off.
1786 2021-09-05 12:49 SULB Gain Check 1.0 at 100 sps.
1787 2021-09-05 12:49 SULB Require GPS Lock On.
1788 2021-09-05 12:50 SULB Site Geology Gravel pit.
1789 2021-09-05 12:50 SULB VSAT Check 27.5 dB isolation on Telesat.
1790 2021-09-05 12:51 SULB Comments This is the 7th and last of the Nechako Basin.
1791 2021-09-05 14:55 SULB Comments Trident SN in Remote Cygnus Ringbuffer corrected to TRD080.
1792 2021-06-23 21:15 SUNO New Station Code Start recording in SUNO ringbuffers.
No Date Time UT Location Activity Comments
1793 2021-06-23 21:15 SUNO Hardware List CYG388 TRD360 CMG-40T T4776, SPS 100, Gain 1.
1794 2021-06-23 21:15 SUNO Site Geology Bedrock
1795 2021-06-23 21:15 SUNO VSAT Start co-pol 34 x-pol 2
1796 2021-08-11 19:50 SUNO DC Filter Check Unchanged from ON at 0.001 Hz.
1797 2021-08-11 19:50 SUNO Gain Check Software gain unchanged from 1.0.
1798 2021-10-31 13:04 SUNO Firmware Upgrade Trident 1.7210 in B. B is active.
1799 2021-10-31 15:06 SUNO Service Visit Station was working, some hardware needed tigthening.
1800 2021-10-31 15:08 SUNO VSAT Check Before 32/5 after 32/1 dB.
1801 2021-10-31 15:09 SUNO VSAT Start Cold boot and restarted VSAT.
1802 2021-11-03 17:33 SUNO Firmware Upgrade Trident 1.7210 in A. A is active.
1803 2021-11-03 17:34 SUNO Comments Libra firmware is out of date (5.50)
1804 2021-12-08 20:00 SUNO Service Visit Added 6 more batteries to power system.
1805 2021-12-08 20:30 SUNO Battery Check Started charging batteries with generator.
1806 2021-12-08 20:54 SUNO Battery Check Battery volt is now at 12.58VDC.
1807 2021-12-09 15:20 SUNO Battery Check Stopped generator. Batteries are at 12.94 VDC.
1808 2021-12-09 15:28 SUNO Seismo Check NS component has large DC offset, looks odd. Bad Seismo??
No Date Time UT Location Activity Comments
1809 2022-02-21 22:01 SUNO VSAT Check EIRP level reduced from 10 to 6 dB balance levels.
1810 2021-06-19 17:00 SUNO Service Visit To upgrade seismometer. VSAT kept on. Bus cable disconnected to swap seismo.
1811 2021-06-19 17:14 SUNO Seismo Swap CMG-40T s/n T4776 removed, ESP s/n T3J32 installed.
1812 2021-08-20 13:48 SUNO Service Visit To add four more solar panels.
1813 2021-08-20 14:03 SUNO Restart Station restarted adding solar panels. A total of 12 panels power the system now.
1814 2021-08-20 14:04 SUNO VSAT Check Co/X 33/1 dB, unchanged.
1815 2021-08-20 16:04 SUNO Cold-Boot Standard procedure.
1816 2021-08-20 16:06 SUNO Gain Check Unchaged at 1.0 (this is default for FedNor1).
1817 2021-08-20 16:07 SUNO DC Filter Check Unchaged at on and 1 mHz.
1818 2021-08-20 16:08 SUNO Comments Station in good condition.
1819 2021-08-28 00:54 TALB New Station Code Tatla Lake, BC, Ottawa Hub, Naqs1,2, Telesat.
1820 2021-08-28 00:55 TALB Hardware List Cyg285, Trd83, TS194, 3ESP T3E02, SSPB 1W A0124A.
1821 2021-08-28 00:56 TALB DC Filter Check Off.
1822 2021-08-28 00:56 TALB Gain Check 1.0 at 100 sps.
1823 2021-08-28 00:57 TALB Battery Check 6 new batteries, 8 solar panels.
1824 2021-08-28 00:57 TALB Require GPS Lock On.
No Date Time UT Location Activity Comments
1825 2021-08-28 00:57 TALB Site Geology Soft soil
1826 2021-08-28 00:57 TALB VSAT Check 28 dB isolation.
1827 2021-08-27 00:48 THMB New Station Code Started Thunder Mountain, BC, Ottawa Hub Naqs1,2 on Telesat.
1828 2021-08-27 00:49 THMB Hardware List Cyg271, Trd256, TS230, ESP T3D79, SSPB A0159A.
1829 2021-08-27 00:51 THMB Battery Check 6 new batteries, 8 solar panels.
1830 2021-08-27 00:51 THMB DC Filter Check Off.
1831 2021-08-27 00:51 THMB Gain Check Gain of 1.0 at 100 sps.
1832 2021-08-27 00:51 THMB Require GPS Lock On.
1833 2021-08-27 00:51 THMB VSAT Check 28 dB isolation.
1834 2021-08-27 00:52 THMB Comments Initial problem with GPS, long wait before fine lock.
1835 2021-08-27 00:52 THMB Site Geology Soft soil.
1836 2021-08-27 00:53 THMB Comments Bad seismometer cable, replaced with good one.
1837 2021-06-27 20:00 THMB Service Visit To improve VSAT.
1838 2021-06-27 20:01 THMB Seismo Check Re-levelled. Cable connector cleaned up.
1839 2021-06-27 20:01 THMB VSAT Check 36 dB separation.
1840 2021-06-27 20:03 THMB Comments Mass-centered and calibrated OK.
No Date Time UT Location Activity Comments
1841 2021-09-29 02:15 TIGG New Station Start Tignish, PEI, first of re-deployment.
1842 2021-10-02 23:31 TIGG Recording Hub London.
1843 2021-10-02 23:32 TIGG Gain Check Gain 1.0, sps 100.
1844 2021-10-02 23:33 TIGG DC Filter Check DC filter is off.
1845 2021-10-02 23:33 TIGG Require GPS Lock On.
1846 2021-10-02 23:34 TIGG Hardware List Cyg172, TRD78, sspb h0018a, esp s/n t3d95, IP 172.
1847 2021-10-02 23:35 TIGG VSAT Check 0/16 dB. Poor.
1848 2021-08-30 20:58 TIGG Carina Swap Switched from SES-AM to Telesat at Ottawa Hub.
1849 2021-08-30 20:59 TIGG SSPB Check Gain adjusted to 50 dB, EIRP adjusted to match.
1850 2021-08-30 20:59 TIGG VSAT Check 28 dB isolation while raining at the site.
1851 2021-07-25 19:26 TIMO New Station Code O62T is now started in TIMO at Ottawa Hub Naqs 2.
1852 2021-07-25 19:28 TIMO New Station Code Started at Ottawa Hub Naqs1 in name TIMO.
1853 2021-07-27 01:40 TIMO Cables Problem with RF out at SSPB end fixed. Station restarted.
1854 2021-07-29 15:43 TOBO DC Filter Check On at 0.001 Hz.
1855 2021-07-29 15:43 TOBO Gain Check Station Started at software gain of 0.5, sample rate 100 sps.
1856 2021-07-29 15:43 TOBO Hardware List CYG413 TRD363 CMG-40T 4401, SSPB A0308A
No Date Time UT Location Activity Comments
1857 2021-07-29 15:43 TOBO New Station Code Start recording in TOBO ringbuffers.
1858 2021-07-29 15:43 TOBO Site Geology Limestone bedrock.
1859 2021-07-29 15:43 TOBO VSAT Start Co-pol 28 X-pol 2
1860 2021-11-22 20:45 TOBO Temp Halt Station stopped. No AC power, Battery volts at 8.7 VDC.
1861 2022-03-26 22:45 TOBO Service Visit To replace seismo. Station found in good and secure condition.
1862 2022-03-26 22:55 TOBO Hardware Swap 40T removed, ESP (T3G89) installed.
1863 2022-03-26 23:10 TOBO Comments Laptop did not work so no re-pointing possible. No software changes. Serviced by KA, database entry IA.
1864 2021-11-04 12:17 TOBO Hub Swap Finished hub swap. TX disabled on carina 28 and 29.
1865 2022-05-04 19:00 TORO New Station Code Started recording at TORO ringbuffres.
1866 2022-05-04 19:10 TORO Hardware List Cyg173 TRD059 CMG-40T T4565 SSPB A0157A
1867 2022-05-04 19:11 TORO Gain Check Software gain 0.5 at 100 SPS.
1868 2022-05-04 19:12 TORO DC Filter Check Filter on at 1 mHZ (check 1 mHz again).
1869 2022-05-04 19:15 TORO Site Geology Man-made fill.
1870 2022-05-10 15:10 TORO DC Filter Change DC filter incorrectly set to 0.1 Hz.
1871 2022-05-10 15:10 TORO Hardware Swap Trd. 077 installed, 059 removed.
1872 2022-05-10 18:33 TORO DC Filter Change Corrected cutoff freq. to 0.001 HZ.
No Date Time UT Location Activity Comments
1873 2021-06-07 13:32 TORO Cables Seismo cable replaced to fix EW component. Some noise heard before cable swap, not after.
1874 2021-06-07 13:34 TORO Comments No system shut down. Cable swap took a couple of minutes. Waveform show good traces for now.
1875 2021-09-02 13:10 TORO Service Visit To improve bad VSAT power.
1876 2021-09-02 13:30 TORO Re-Point Antenna was off. Before 20 dB. After, 31/0 dB.
1877 2021-09-02 13:47 TORO Comments Non-POLARIS lock on battery box. BAD!!!
1878 2021-05-18 14:?? TSJB Hardware List Guralp ESP T3D84 Cygnus 266 Trident 76
1879 2021-05-18 14:?? TSJB New Station Code Start recording in TSJB ringbuffers.
1880 2021-06-18 23:50 TSJB Closed End of Operation.
1881 2021-11-27 00:11 TWBB New Station Code Start recording in TWBB ringbuffers.
1882 2021-11-27 00:12 TWBB Gain Check Starting with software gain of 0.1.
1883 2021-11-27 00:12 TWBB DC Filter Check Starting with DC filter on at 0.001 Hz.
1884 2021-08-05 18:00 TWBB Service Visit Station vandalized, battery box left open and was full of water. Battery positive terminal corroded. Positive lead disconnected. Box drained. Batteries cleaned and connected. Appear to be working.
1885 2021-08-05 19:40 TWBB VSAT Check Co/X 34/5 dB.
1886 2021-08-05 19:42 TWBB Cold-Boot Standard procedure.
1887 2021-08-05 19:49 TWBB Restart Station restarted, all OK.
1888 2021-08-26 18:30 TWBB Service Visit To replace broken Cygnus 177.
No Date Time UT Location Activity Comments
1889 2021-08-26 18:35 TWBB Hardware Swap Cyg177 removed, Cyg285 installed.
1890 2021-08-26 18:40 TWBB VSAT Check Co/x 18/0, low.
1891 2021-08-26 18:42 TWBB Comments Site is back to normal.
1892 2021-06-20 18:06 TWBB Closed End of Operation.
1893 2021-11-25 TWGB New Station Code Start recording in TWGB ringbuffers.
1894 2021-11-25 19:40 TWGB Gain Check Starting with software gain of 0.1.
1895 2021-11-25 19:40 TWGB DC Filter Check Starting with DC filter on at 0.001 Hz.
1896 2021-08-05 20:15 TWGB Service Visit On the way back from TWBB, power cable from junction box bitten but OK. Loose connector fixed.
1897 2022-05-10 19:11 TWGB Closed End of Operation.
1898 2021-11-20 23:30 TWKB Gain Check Starting with software gain of 0.1.
1899 2021-11-20 23:30 TWKB DC Filter Check Starting with DC filter on at 0.001 Hz.
1900 2021-11-20 23:30 TWKB New Station Code Start recording in TWKB ringbuffers.
1901 2021-08-25 17:15 TWKB Service Visit Station shut down to check seismometer orientation.
1902 2021-08-25 18:00 TWKB Orientation Change Seismometer orientation was 179 degrees worng. Corrected.
1903 2021-08-25 18:15 TWKB Restart Restarted.
1904 2022-02-12 19:20 TWKB Seismo Check HHN Component is bust. Mass-center did not help. Calibration shows odd signal on HHN but other two chans. ok. Seismo stuck?
No Date Time UT Location Activity Comments
1905 2022-03-15 21:40 TWKB Hardware Swap TRD071, CMG-3ESP T3D91 removed. TRD425, CMG-3ESP T3G52 installed.
1906 2022-03-15 21:41 TWKB DC Filter Check DC filter on at 0.0010 Hz
1907 2022-03-15 21:42 TWKB Gain Check Software gain = 0.5
1908 2022-05-14 10:26 TWKB Closed End of Operation.
1909 2021-07-27 TYNO New Station Start First POLARIS station. HRD digitizer and CMG-3ESP. SOSN/POLARIS. AC Powered.
1910 2021-10-31 22:00 TYNO Hardware Swap Replaced Cygnus 154/HRD 184 to Cygnus 282/Trident 259.
1911 2022-04-11 TYNO Site Visit Maintenance visit.
1912 2021-12-19 18:00 TYNO Service Visit Station down about 1/2 hour to swap batteries.
1913 2021-12-19 18:35 TYNO Battery Check AC-type batteries and LVD installed.
1914 2021-09-15 17:30 TYNO Require GPS Lock GPS receiver installed on Libra com port, recording at London Hub.
1915 2022-03-16 14:30 TYNO Service Visit To improve VSAT.
1916 2022-03-16 14:45 TYNO VSAT Check Co/X improved to 28/0 dB. Power level up by 5 dB.
1917 2022-03-16 15:00 TYNO Cold-Boot System restarted.
1918 2021-06-17 22:00 TYNO VSAT Check Co/x was good at 32/0 dB. Station need sspb temperature sensor.
1919 2021-07-16 UWO Hub Start First POLARIS Hub started at UWO.
1920 2021-08-14 22:14 UWO-Hub Hub Down UWO Hub Power Failure. Loss of at least 1/2 hour of data on the network.
No Date Time UT Location Activity Comments
1921 2022-02-26 UWOO New Station Start AC powered. Noisy site on UWO campus.
1922 2022-05-08 UWOO Closed Finished with test station.
1923 2021-06-11 13:00 VIMO Hardware List Cyg 364, Trd 370, ESP T3F09. SPS 100, Gain 1.
1924 2021-06-11 13:00 VIMO New Station Code Start recording in VIMO ringbuffers.
1925 2021-06-11 13:00 VIMO Site Geology Limestone.
1926 2021-06-11 13:00 VIMO VSAT Start X-pol 0, Co-pol 32 dB.
1927 2021-08-11 19:30 VIMO DC Filter Check Unchanged from ON at 0.001 Hz.
1928 2021-08-11 19:30 VIMO Gain Check Software gain unchanged from 1.0.
1929 2021-11-03 18:09 VIMO Comments Libra firmware is out of date (5.50)
1930 2021-11-03 18:35 VIMO Firmware Upgrade Trident 1.7210 in A and B. B is active.
1931 2022-02-21 00:03 VIMO VSAT Start Modulation reduced from 112 to 64 KB to improve chance of connecting to the station.
1932 2022-02-21 00:04 VIMO VSAT Check EIRP level increased from 6 to 8 dB to get the station up again.
1933 2022-02-21 22:00 VIMO VSAT Check Station back up after VSAT adjustments.
1934 2021-06-20 14:56 VSAT Comments New GE Test Frequencies: Tx 14031.0 MHz, Rx 11731.0 MHz reported by KA.
1935 2021-08-20 20:36 WAGN New Station Code Wager Bay, Hudson Bay, Naqs2, Ottawa Hub, Telesat.
1936 2021-08-20 20:47 WAGN Hardware List Cyg232, TS188, Trd179, ESP T3A89, SSPB H0017A, redeployed from YNEN.
No Date Time UT Location Activity Comments
1937 2021-08-20 20:48 WAGN DC Filter Check Off.
1938 2021-08-20 20:48 WAGN Gain Check 1.0 at 40 sps.
1939 2021-08-20 20:48 WAGN Require GPS Lock On.
1940 2021-08-20 20:48 WAGN VSAT Check 30 dB isolation on Telesat.
1941 2021-08-20 20:49 WAGN Battery Check 16 new batteries, 12 solar panels.
1942 2021-08-20 20:49 WAGN Site Geology On Esker near small stream 50 m off runway.
1943 2021-08-20 20:50 WAGN New Station Code Wager Bay, Hudson Bay, Naqs1, Ottawa Hub, Telesat.
1944 2021-08-20 20:54 WAGN Comments Dataserver started on Naqs1.
1945 2021-08-21 11:59 WAGN Comments Site location is at Sila Lodge air strip, Wager Bay.
1946 2021-06-23 16:25 WEMQ New Station Code Started on Naqs2 at Ottawa around 16:25 and on Naqs1 around 16:31.
1947 2021-06-23 16:30 WEMQ Hardware List Cyg601, trd 833 ts 737, ESP, S/N T3P32, Sspb a0017a.
1948 2021-06-23 16:31 WEMQ VSAT Check 32 dB isolation but very difficult to start. Had to increase Eirp to +10 dB, fake sspb gain by 3 dB and IFL Loss to 5 dB.
1949 2021-06-23 16:32 WEMQ Gain Check SPS 100, Gain 1.0, Vpp 40.
1950 2021-06-23 16:33 WEMQ DC Filter Check No DC filter.
1951 2021-06-23 16:35 WEMQ Site Geology Solid bedrock.
1952 2021-06-27 17:35 WEMQ Service Visit To check VSAT antenna. No problems noted.
No Date Time UT Location Activity Comments
1953 2022-01-29 WLVO New Station Start SOSN/POLARIS. AC Powered.
1954 2021-09-16 23:25 WLVO DC Filter Change From off to on. At 0.001 mHz corner frequency.
1955 2021-09-16 23:25 WLVO Gain Change From 1.048576 to 0.1.
1956 2021-09-16 23:35 WLVO Firmware Upgrade Upgraded to Libra V5.20, TS V1.20, TR V1.50 EPLD V1.08
1957 2021-09-23 17:39 WLVO Firmware Upgrade Upgraded to Libra Firmware version 5.31.
1958 2021-11-07 18:35 WLVO Hardware Swap Replaced Cygnus 171/Trident 63 to Cygnus 276/Trident 258.
1959 2022-04-10 17:10 WLVO Site Visit Maintenance visit.
1960 2022-04-10 17:10 WLVO VSAT Check Realigned. Co-pol 28, X-pol 0 dB.
1961 2021-12-16 16:00 WLVO Service Visit To install DC power system. AC power not available any more.
1962 2021-12-16 19:30 WLVO Reopened Station started with solar power. No change in configuration.
1963 2021-12-16 19:56 WLVO VSAT Check No change at 30/0 dB.
1964 2022-03-16 17:30 WLVO Service Visit To improve VSAT.
1965 2022-03-16 17:45 WLVO VSAT Check Co/X improved to 27/0 dB. Power level up by 6 dB.
1966 2022-03-16 18:00 WLVO Cold-Boot Station restarted.
1967 2021-08-27 21:12 YBKN New Station Code Baker Lake, Hudson Bay, Ottawa Hub, Naqs1,2, Telesat.
1968 2021-08-27 21:13 YBKN Hardware List Cyg283, Trd936, TS187, ESP T3B15, SSPB, 1W H0026A.
No Date Time UT Location Activity Comments
1969 2021-08-27 21:14 YBKN Gain Check 1.0 at 40 sps.
1970 2021-08-27 21:15 YBKN DC Filter Check Off.
1971 2021-08-27 21:15 YBKN Require GPS Lock On.
1972 2021-08-27 21:15 YBKN Site Geology Gravel, airport area.
1973 2021-08-27 21:15 YBKN VSAT Check 27 dB Isolation.
1974 2021-08-27 21:16 YBKN Battery Check 2 DC-type batteries used in this AC site.
1975 2021-08-27 21:17 YBKN Comments Dish in low-lying area, 20 m away from Env. Canada Bldg.
1976 2021-08-27 21:17 YBKN Power Check AC Power to be connected in a day or so.
1977 2021-09-07 YMBN New Station Start First station of NWT array. DC (Solar Panels).
1978 2021-07-21 YMBN Hardware Swap Cygnus 157/HRD 210 replaced with Cygnus 246/Trident 235.
1979 2021-07-21 YMBN Temp Stop Stopped recording in ringbuffers YMBN to swap electronics.
1980 2021-07-21 19:00 YMBN SR Change Sampling rate from to 20 sps to 40 sps.
1981 2021-07-21 23:00 YMBN Restart Started recording in new YMBN ringbuffers.
1982 2021-09-19 22:40 YMBN Firmware Upgrade Upgraded to Libra Firmware version 5.31.
1983 2021-09-19 22:40 YMBN Repair Wildlife damage repaired. Station re-started.
1984 2021-09-25 18:55 YMBN SSPB Check Corrected SSPB temperature profile from flat 45 dB to manufacturer specs.
No Date Time UT Location Activity Comments
1985 2021-07-09 20:42 YMBN Service Visit Site was working OK. Stopped for service.
1986 2021-07-09 21:48 YMBN Before Re-Point Before any re-adjustment, co-pol was 27 and x-pol was 0 dB, indicating stable antenna base.
1987 2021-07-09 21:48 YMBN Clean up Garbage and old battery box prepared for removal.
1988 2021-07-09 21:48 YMBN Feed Horn Box New Feed Horn protective box installed. New conduits installed.
1989 2021-07-09 21:48 YMBN Firmware Upgrade Libra 5.6201 installed. SSPB S/N H0013A.
1990 2021-07-09 21:48 YMBN Re-Point After re-pointing, co-pol was 28 with x-pol of 0 dB. Not much improved as expected.
1991 2021-07-09 21:48 YMBN Seismo Check Blue tarp was off the vault. Buried the vault. No mass-centering needed. Calibrated ok.
1992 2021-09-15 16:41 YMBN MT Station Libra ports not showing any read or write. MT data stopped.
1993 2021-09-17 13:51 YMBN MT Station Rebooting Cygnus did not restart the MT instrument . MT Station is down.
1994 2021-09-17 14:00 YMBN MT Station Rebooted MT instrument by sending Ctrl-C on telnet port. Station restarted.
1995 2021-07-10 14:48 YMBN Closed End of Operation.
1996 2021-07-24 18:00 YNEN LVD Check The Low Volt Disconnect set to 11.0 VDC instead of 11.5.
1997 2021-07-24 18:00 YNEN New Station Code Start recording in YNEN ringbuffers.
1998 2021-07-25 01:00 YNEN New channel prototype Channel prototype to Broad band, Trident and CMG-40T seismometer (BHZ-2 to BHZ-3, BHN-2 to BHN-3, BHE-2 to BHE-3).
1999 2021-08-21 19:45 YNEN Firmware Upgrade Upgraded to Libra V5.20, TS V1.20, TR V1.50 EPLD V1.08
2000 2021-09-22 16:51 YNEN Firmware Upgrade Upgraded to Libra Firmware version 5.31.
No Date Time UT Location Activity Comments
2001 2021-09-22 16:51 YNEN Repair Wildlife damage repaired. Station re-started.
2002 2021-09-25 18:10 YNEN SSPB Check Corrected SSPB temperature profile from flat 45 dB to manufacturer specs.
2003 2021-07-06 20:11 YNEN Feed Horn Box New Feed Horn protective box installed. New conduits installed.
2004 2021-07-06 20:11 YNEN Firmware Upgrade Libra 5.6201, TS 1.23, Trd 1.7201, EPLD 1.23
2005 2021-07-06 20:11 YNEN Hardware Swap New SSPB installed. S/N H0017A.
2006 2021-07-06 20:11 YNEN Hub Swap Served by Ottawa Hub.
2007 2021-07-06 20:11 YNEN Re-Point Co-pol 25, X-pol 0 dB.
2008 2021-07-06 20:11 YNEN SOH Monitor Working OK.
2009 2021-07-06 20:11 YNEN Service Visit Wildlife damage to SSPB.
2010 2021-08-12 12:18 YNEN DC Filter Check Unchanged from Off.
2011 2021-08-12 12:18 YNEN Gain Change Software gain changed from 1.048576 to 1.0.
2012 2021-11-04 16:44 YNEN Firmware Upgrade Trident EPLD 1.23 installed in B.
2013 2021-11-04 16:54 YNEN Firmware Upgrade Trident 1.7210 in A and B. B is active.
2014 2021-07-09 15:59 YNEN Closed End of Operation.
2015 2021-06-26 20:47 YOSQ New Station Code Started on Naqs2 at Ottawa around 20:47 and on Naqs1 around 20:52 UT.
2016 2021-06-26 20:50 YOSQ Hardware List Cyg-599, trd-832, ts 834, esp s/n t3n25, sspb a0018a 2 watt.
No Date Time UT Location Activity Comments
2017 2021-06-26 20:51 YOSQ Gain Check SPS 100, Gain 1.0, Vpp 40.
2018 2021-06-26 20:52 YOSQ DC Filter Check No DC filter.
2019 2021-06-26 20:53 YOSQ VSAT Check Antenna was not aligned correctly. Had to re-define sweet spot.
2020 2021-06-26 20:55 YOSQ Site Geology Seismometer vault in sand over solid bedrock
2021 2021-05-15 15:20 YOUB Hardware List Guralp ESP T3D81 Cygnus 187 Trident 81
2022 2021-05-15 15:20 YOUB New Station Code Start recording in YOUB ringbuffers.
2023 2021-09-06 22:31 YRTN New Station Code Started recording at Ottawa Naqs2 in YRTN ringbuffers.
2024 2021-09-06 22:33 YRTN New Station Code Started recording at Ottawa Naqs1 in YRTN ringbuffers.
2025 2021-09-06 22:35 YRTN Hardware List Cygnus 550, Trident 767, TimeServer 633, sspb A0400A. ESP .
2026 2021-09-06 22:36 YRTN Seismo Check Bottom of seismometer vault was wet.
2027 2021-09-06 22:37 YRTN Gain Check Started with gain of 1.0 at 40 sps.
2028 2021-09-06 22:38 YRTN DC Filter Check Started with DC filter off.
2029 2021-10-07 15:40 YRTN Service Visit To replace seismometer, repoint
2030 2021-10-07 15:41 YRTN Comments Seismo vault flooded. seimo sub-merged. Vault re-built with raised pier.
2031 2021-10-07 15:42 YRTN Hardware Swap Cyg 217, Trd 070, ESP T3L28 installed. New sspb a0498a .
2032 2021-10-07 15:43 YRTN VSAT Check Faild sspb replaced. Power is good. Eirp inc at +2 dB.
No Date Time UT Location Activity Comments
2033 2021-10-07 15:45 YRTN Gain Check Unchaged at 1 and 40 sps.
2034 2021-10-07 15:46 YRTN DC Filter Check Unchaged at no DC filter.